### v1

# Large Language Model Can Continue Evolving From Mistakes

Abstract
Large Language Models (LLMs) demonstrate impressive performance in various downstream tasks. However, they may still generate incorrect responses in certain scenarios due to the knowledge deficiencies and the flawed pre-training data. Continual Learning (CL) is a commonly used method to address this issue. Traditional CL is task-oriented, using novel or factually accurate data to retrain LLMs from scratch. However, this method requires more task-related training data and incurs expensive training costs. To address this challenge, we propose the Continue Evolving from Mistakes (CEM) method, inspired by the ’summarize mistakes’ learning skill, to achieve iterative refinement of LLMs. Specifically, the incorrect responses of LLMs indicate knowledge deficiencies related to the questions. Therefore, we collect corpora with these knowledge from multiple data sources and follow it up with iterative supplementary training for continuous, targeted knowledge updating and supplementation. Meanwhile, we developed two strategies to construct supplementary training sets to enhance the LLM’s understanding of the corpus and prevent catastrophic forgetting. We conducted extensive experiments to validate the effectiveness of this CL method. In the best case, our method resulted in a 17.00% improvement in the accuracy of the LLM.
Large Language Model Can Continue Evolving From Mistakes
Haokun Zhao, Haixia Han, Jie Shi, Chengyu Du, Jiaqing Liang, Yanghua Xiao Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University hkzhao23@m.fudan.edu.cn
1 Introduction
Figure 1: Continual evaluation can help us uncover knowledge deficiency of the LLM and thus evolve the model through continuous learning.
With the exponential expansion of both data and parameter size, large language models (LLMs) demonstrate outstanding performance across a series of downstream tasks Chang et al. (2023). However, LLMs have the potential to generate incorrect outputs, a phenomenon often referred to as ’hallucination’ Huang et al. (2023). In addition, it is important to note that satisfactory answers generated by LLMs cannot be achieved within a single training session Han et al. (2024). Continual refinement and optimization processes are necessary. This optimization process ensures that the LLM can undergo iterative refinement, continuously enhancing its performance.
There are two main reasons for the poor performance of LLM on specific questions. Firstly, some of these inquiries exceed the knowledge boundaries of LLM, indicating either a deficiency in domain expertise or outdated information within the LLM Ren et al. (2023); Zhang et al. (2023). Secondly, the LLM may incorporate erroneous knowledge derived from potentially unreliable sources in the training data, resulting in the generation of inaccuracies or misconceptions that contradict factual information Rae et al. (2021); Zhao et al. (2023).
To rectify the model’s knowledge deficits and ensure alignment with contemporary world knowledge, a common method is periodic training from scratch using the domain or latest corpus, such as the most recently archived Wikipedia. However, this method incurs significant training expenses Patterson et al. (2021). Another method involves pre-training on a corpus that includes accurate information. However, this method may result in a decline in other aspects of the model’s performance Li et al. (2022). This challenge has prompted research into mitigating catastrophic forgetting in continuous learning. Existing studies can be roughly grouped into three categories, e.g., experience replay methods Chaudhry et al. (2019); Wu et al. (2021), regularization-based methods Kirkpatrick et al. (2017a); Chen et al. (2022), and dynamic architecture methods Mallya et al. (2018). Recently, researchers have designed some hybrid methods that take advantage of the aforementioned techniques Chen et al. (2023); He et al. (2024).
In this paper, we focus on the process of uncovering knowledge that the model has not yet mastered, for supplementation. Throughout each iterative phase of model evolution, we identify the limitations of the model’s knowledge and address them in subsequent updates. This idea is inspired by the ’summarizing mistakes’ learning skill observed in human students. An evaluation is conducted initially to identify a student’s weaknesses in a subject. References related to these errors are then provided to assist the student in mastering the associated knowledge. For instance, if a student is struggling with chemistry, they can review questions answered incorrectly in chemistry tests to pinpoint areas needing improvement and extract relevant knowledge, such as the properties of benzene, from the corresponding chapter in the chemistry textbook. This process enhances his performance in chemistry by identifying and learning from weaknesses or areas of unfamiliarity.
In this paper, we propose the Continue Evolving from Mistakes (CEM) method. The aim is to systematically address the model’s knowledge gaps and rectify incorrect knowledge through continuous evolution. Our contributions are summarized as follows:
  • To endow LLMs with the capacity for continuous iterative evolution, we have drawn inspiration from the ’summarize mistakes’ learning skill observed in human students. We introduce CEM, which explores the knowledge boundaries from LLM mistakes, thereby continuously enhancing the model’s capabilities.
  • To assist LLMs in acquiring and comprehending factual knowledge from supplementary corpora, as well as mitigating knowledge forgetting, we propose two strategies for constructing supplementary corpora.
  • We conduct experiments on three open-source LLMs and two complex knowledge Question-Answering (QA) tasks. The results demonstrate the efficacy of the CEM.
2 Related Work
Instruction fine-tuning LMs often struggle to maintain alignment well with human intent after pre-training. However, this issue can be alleviated by employing techniques such as Instruction Fine-Tuning (IFT) Ouyang et al. (2022); Stiennon et al. (2022) and Reinforcement Learning from Human or AI Feedback (RLHF/RLAIF)Gao et al. (2022); Bai et al. (2022). IFT involves training an LM to generate responses based on provided prompts, which might optionally include task instructions. The data used for fine-tuning is usually less than the original pre-training corpus. These datasets used in IFT are often curated through crowd-sourcing or extracted from an LM that is already capable of generating instruction-following examples. Studies have shown that IFT can enhance zero-shot generalization to various unseen tasks. To optimize the fine-tuning process, various strategies have been proposed to make it more parameter-efficient. These strategies include using adapters Hu et al. (2021), prompt tuning Li and Liang (2021), and other techniques.
Continual learning with LLM Continual learning (CL) aims to enable a model to learn from a continuous data stream across time while reducing catastrophic forgetting of previously acquired knowledge Biesialska et al. (2020). With CL, a deployed LLM has the potential to adapt to the changing world without costly re-training from scratch Bubeck et al. (2023). Early works Gururangan et al. (2020); Röttger and Pierrehumbert (2021); Lazaridou et al. (2021); Dhingra et al. (2022) empirically analyze continuing LM pre-training on emerging domain or temporal data, showing the potential to update the base LM with new knowledge. Jang et al. (2022b) explicitly categorize world knowledge as timeinvariant, outdated, and new knowledge, which should be retained, acquired, and updated respectively by an LM when learning continually. Jin et al. (2022); Jang et al. (2022a) additionally implement traditional CL methods to alleviate catastrophic forgetting, a phenomenon in which previously learned knowledge or abilities are degraded due to overwritten parameters Kirkpatrick et al. (2017b).
3 Methods
The objective of this study is to enhance the knowledge base of LLM through continual learning, resulting in iterative improvements in model performance.
3.1 Task Formulation
Question-Answering (QA) tasks are commonly used to evaluate models’ world knowledge and reasoning skills. Our focus is on enhancing the models’ world knowledge rather than their reasoning abilities. These tasks usually involve a series of single-choice questions with multiple options. The ability of a LLM to correctly identify the correct answer within these questions indicates its possession of the relevant knowledge. Conversely, failure to do so suggests a deficiency in the model’s understanding of the knowledge pertinent to the question.
Given a question-answer example $(qi,ai)\in\mathcal{Q}$, a test set for question answering $\mathcal{Q}$, and a reasoning model $\mathcal{M}_{0}$.
First, the reasoning model $\mathcal{M}_{0}$ is utilized to answer all questions within the test set $\mathcal{Q}$. Subsequently, questions $q_{i}$ that receive incorrect responses are collected,
$\widetilde{r_{i}}=\mathcal{M}_{0}(q_{i}),\quad Ans(\widetilde{r_{i}})\neq a_{i}$
where $\widetilde{r_{i}}$ denotes the response that includes incorrect answers to the question $q_{i}$, and $Ans(\cdot)$ is utilized to extract the final answer from the response.
Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplementary corpus $C_{S_{i}}$,
$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\dots\}$
where $Search(\cdot)$ is used to search for relevant knowledge based on the question and retrieve supplementary corpus.
Subsequently, based on the supplementary corpus $C_{S}$, construct the supplementary training set $D$ to fine-tune model $\mathcal{M}_{0}$,
$D=Construct(C_{S}),\quad\mathcal{M}_{1}=Train(\mathcal{M}_{0},D)$
where $Construct(\cdot)$ is utilized to build an supplementary training set, $Train(\cdot)$ denotes fine-tuning. $\mathcal{M}_{1}$, represents the model that has undergone one iteration of supplementary training.
We can continue to employ $\mathcal{M}_{1}$ as the reasoning model on the test set $\mathcal{Q}$, for iterative supplementary training, thereby achieving continual optimization of the model’s performance.
3.2 CEM Method
Specifically, the CEM method proposed in this paper involves three main steps: (1) testing the LLM and collecting incorrectly answered questions, (2) collecting supplementary corpora from multiple data sources, and (3) constructing a supplementary training set and fine-tuning the LLM. Fig.2 shows the process.
Figure 2: The pipeline of CEM method.
Testing LLM and Collecting Incorrectly Answered Questions. We utilize the LLM to answer questions from a QA dataset. Subsequently, by utilizing pattern matching to extract answers from the LLM’s responses, we verify the accuracy of these answers and collect the questions that the LLM has answered incorrectly.
Collecting Supplementary Corpora from Multiple Data Sources. We argue that the questions answered incorrectly by the LLM reveal knowledge deficiencies. Based on the implicit information contained within questions incorrectly answered by the LLM, we collect supplementary corpora from multiple data sources, including online encyclopedias and search engines.
(1) Search Engine: We submit questions to search engines and use their retrieval techniques to crawl relevant content from the vast corpus of Chinese-language internet. Search engines are characterized by fast response time, high relevance matching, and diverse and rich content. We conduct experiments on two search engines, Baidu and Bing, and optimized our corpus crawling strategy.
Specifically, we search for questions that the LLM answers incorrectly. The search engine ranks results based on relevance. We select the top-k results with the highest relevance and access their detailed pages to crawl the main textual content as supplementary corpora. To prevent the LLM from simply memorizing answers, we filter out search results that may contain both the original question and its answer. In this experiment, we utilize Bing as the search engine and set $k=4$, indicating that the main textual content from the detailed pages of the top-4 search results with the highest relevance was retrieved. Fig.5 shows process of collecting corpora from search engine.
(2) Online Encyclopedia: Compared to using search engines for retrieval, online encyclopedias provide more credible and professional search results, despite having lower hit rates than search engines. In our experiment, we experiment two online encyclopedias, Baidubaike and Wikipedia.
Specifically, since the organization of encyclopedias typically relies on keyword entries, we conduct entity extraction from the questions and options of the collected incorrectly answered questions, using the extracted entity keywords for retrieval. The main text of the retrieved results serve as supplementary corpora.
In our experiment, we use Wikipedia as the online encyclopedia and instruct ChatGPT-3.5 using a few-shot prompt for entity extraction. Appendix Fig.6 shows process of collecting corpora from online encyclopedia.
Constructing Supplementary Training Set and Fine-Tuning LLM. The Supplementary Training Sets are constructed separately based on the Supplementary Corpus collected from Wikipedia and Bing. They are then employed to fine-tune the LLM. The construction method of the training set will be explained in detail in Section 3.3.
After completing one round of supplementary training, the model is retested on the QA dataset to gather a new set of incorrectly answered questions, which serve as the first step for the next round of supplementary training.
It should be noted that supplementary training should not be endless, as this may lead to severe catastrophic forgetting and the boundary effect of diminishing accuracy improvement.
3.3 Supplementary Training Data
Corpus Format Definition
The construction of the supplementary training set involves the following four types of corpus.
(1) Format Corpus: Table 1 shows an example of a Format Corpus. Prior to implementing the CEM method for supplementary training, we fine-tune the LLM using Format Corpora. This equips the LLM with the ability to respond to multiple-choice questions in a standard format, facilitating the extraction of answers from the LLM’s responses and allowing for the collection of incorrectly answered questions. To maintain this ability, we incorporate a certain proportion of Format Corpus during the supplementary training to prevent the LLM’s response skills from declining.
Example 1: A Case of Format Corpus
Question:
(task)Please select the correct answer for the following single choice questions:
(question)In the 2nd century, what became the Roman Empire’s internal sea was _____
A. Red Sea B. Caspian Sea C. Mediterranean Sea D. Black Sea
Answer:
(answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
Table 1: An example of a Format Corpus, designed to illustrate its corpus format.
(2) Supplement Corpus: Corpus obtained through multiple data sources.
(3) Extract Corpus: Table 2 shows an example of an Extract Corpus. Inspired by the human learning process: despite studying from the same textbooks, students who are better at extracting knowledge from these materials always achieve superior grades. We posit that merely providing the LLM with a supplementary corpus containing knowledge is inadequate. Consequently, we designed the Extract Corpus to train the LLM to effectively capture and comprehend knowledge from supplementary corpora, enhancing its utilization of these materials.
Example 2: A Case of Extract Corpus
Question:
(task)Please select the correct answer for the following single choice questions:
(question)In the 2nd century, what became the Roman Empire’s internal sea was _____
A. Red Sea B. Caspian Sea C. Mediterranean Sea D. Black Sea
(extract prompt)Please answer in the context of the following background material:
(supplement)The Roman Empire … During the reign of Tulajin (98-117), the Roman Empire reached its greatest extent with the Mediterranean Sea becoming the empire’s inland sea. In its heyday, it controlled about 5 million square kilometers of land, making it one of the largest monarchies in the ancient history of the world.
Answer:
(answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
Table 2: An example of a Extract Corpus, designed to illustrate its corpus format. The ellipsis indicates that some supplementary text has been omitted for the sake of clarity.
(4) General Corpus: Replaying previous training data can effectively mitigate catastrophic forgetting when training LLM on new dataHe et al. (2021). Therefore, we utilize the training data from ShareGPT dataset  [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_
unfiltered
as General Corpus to preserve the model’s general capabilities.
Supplementary Training Set Construction Methods. Through the combination of four types of corpora, we have developed three innovative methods for constructing supplementary training sets. These methods are detailed below.
(1) Normal Method: To mitigate the potential catastrophic forgetting resulting from supplementary training, we mix the General Corpus with the Supplement Corpus, ensuring an equal quantity of those two type of corpus across all construction methods,
$D_{N}=Construct(C_{G},C_{S})$
where $Construct(\cdot)$ is utilized for constructing supplementary training set based on corpora, $C_{G}$ represents the General Corpus, $C_{S}$ indicates the Supplement Corpus, and $D_{N}$ denotes the supplementary training set constructed via the Normal Method.
(2) Extract Method: To evaluate whether the Extract Corpus enhances the LLM’s ability to capture and comprehend knowledge from supplementary corpora as expected, and to associate questions with relevant knowledge when answering, we added the Extract Corpus into $D_{N}$ to generate the supplementary training set $D_{E}$.
$D_{E}=Construct(C_{G},C_{S},C_{E})$
We replaced the Extract Corpus with the General Corpus to obtain the control group $D_{E_{C}}$.
$D_{E_{C}}=Construct(C_{G}^{\prime},C_{S})$
(3) Review Method: Under higher temperature parameter, the responses of the LLM tend to be diverse. When conducting repeat QA tests on the LLM, it may change its answers to certain ’less confident’ questions due to the proximity of predicted probabilities. This means that after supplementary training, questions initially answered correctly may be answered incorrectly, resulting in a transition from right to wrong.
Therefore, we strengthen the LLM’s memory of the knowledge implied by correctly answered questions by reviewing the responses provided correctly by the LLM in the previous round.
Specifically, in this round of supplementary training, we sample the correct responses made by the LLM in the previous round and process the correctly answered questions into either ’Format Corpus’ or ’Extract Corpus’ formats. We define it as the Review Corpus, and based on the proportion of ’Format Corpus’ and ’Extract Corpus’ formats in the Review Corpus, we construct $R_{1}$, $R_{2}$, and $R_{3}$,
$C_{R_{1}}=Format(\alpha S_{QA})+Extract((1-\alpha)S_{QA})$
$C_{R_{2}}=Format(S_{QA})$
$C_{R_{3}}=Extract(S_{QA})$
where $\alpha\in(0,1)$, $S_{QA}$ denotes the QAs sampled from the correct responses in the previous round of LLM. $Format(\cdot)$ is employed to process the sampled QAs into Format Corpus format, while $Extract$ is used to process the sampled QAs into Extracted Corpus format.
Note that although Review Corpus looks like a mixture of Extract Corpus and Format Corpus in its compositional form, the corpus of the former comes from the correct responses made by the model and is a proper subset of the test set. In contrast, the corpus used by Extract Corpus and Format Corpus is completely independent of the test set.
We added the Review Corpus into the foundation of $D_{E}$ to generate the supplementary training set $D_{R_{i}}$,
$D_{R_{i}}=Construct(C_{G},C_{S},C_{E},C_{R_{i}}),\quad i=1,2,3$
And we replaced the Review Corpus with the General Corpus to obtain the control group $D_{R_{C}}$.
$D_{R_{C}}=Construct(C_{G}^{\prime\prime},C_{S},C_{E}),$
4 Experiments
4.1 Experimental Settings
Datasets. We conduct experiments on two question-answering datasets, including Xiezhi [^2] https://github.com/mikegu721/xiezhibenchmark and CMMLU [^3] https://github.com/haonan-li/CMMLU.
Xiezhi, a comprehensive evaluation suite designed to assess holistic domain knowledge. It comprises multiple-choice questions across 516 diverse disciplines, ranging from 13 different subjects. Gu et al. (2023).
CMMLU contains about 12,000 questions spanning elementary to advanced professional levels across 67 subjectsLi et al. (2024).
In our experiments, we select 27,265 multiple-choice questions from Xiezhi and 11,917 multiple-choice questions from CMMLU. Appendix Table 7 shows the details of the data division.
Base language models. The objective of our work is to evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even those with smaller parameter scales. We select open-source models with parameter ranges from 6 to 13 billion for our experiment. We use CuteGPT-13B [^4] https://huggingface.co/XuYipei/kw-cutegpt-13b-base, ChatGLM-6B [^5] https://huggingface.co/THUDM/chatglm3-6b, Qwen-7B [^6] https://huggingface.co/Qwen/Qwen-7B-Chat as our instruction fine-tuning base models for CEM. These models have been fine-tuned on the instruction data enabling them to effectively follow instructions. CuteGPT is based on the original Llama model structure, expands the Chinese vocabulary and performs pre-trainingLiang et al. (2023). ChatGLM-6B also is an open bilingual language model. It is trained for about 1T tokens of Chinese and English corpusDu et al. (2022). Qwen-7B-Chat is a Transformer-based large language model, which is pretrained on a large volume of data, including web texts, books, codes, etcBai et al. (2023).
In the instruction fine-tuning stage, CuteGPT-13B and Qwen-7B-chat employ full fine-tuning, and ChatGLM-6B employs Prompt-tuningLiu et al. (2022).
Supplementary training dataset. We construct a variety of supplementary training sets and their respective control groups for the experiments following the method described in Section 3.3. To distinguish between the model using different supplementary training sets, we establish the following naming rules:
(1) Baseline: it represents the model that has not undergone supplementary training and has only been pre-finetuned.
(2) N: N denotes Normal, indicating that the model is supplementary trained using the supplementary training set constructed by the Normal Method.
(3) E and EC: E denotes Extract, indicating that the model is supplementary trained using the supplementary training set constructed by the Extract Method. To obtain its respective control group, Extract Control (EC), we replace the Extract Corpus in the training set with the General Corpus.
(4) R1, R2, R3 and RC: R denotes Review, indicating that the model is supplementary trained using the supplementary training set constructed by the Review Method and based on the proportion of ’Format Corpus’ and ’Extract Corpus’ formats in the Review Corpus, we further name them as R1, R2, and R3. To obtain their respective control group, Review Control (RC), we replace the Review Corpus in the training set with the General Corpus.
In addition, we distinguish between the data sources from which the supplementary corpus is collected and establish the following naming rules:
(1) Wiki: it represents the Supplementary Corpus of the supplementary training set used by the model is from Wikipedia.
(2) Bing: it represents the Supplementary Corpus of the supplementary training set is from Bing.
(3) Mix: it represents the Supplementary Corpus of the supplementary training set is collected from both Bing and Wikipedia and concatenated.
Metrics. The effects of supplementary training will be evaluated using the following metrics:
(1) ACC: it represents the accuracy of the answer of LLM.
(2) W2R: it indicates the percentage of questions where the LLM changes incorrect answers to correct ones in the test set.
(3) R2W: it indicates the percentage of questions where the LLM changes correct answers to incorrect ones in the test set.
In Section 4.2, we combine the above two naming rules to describe the experimental results clearly. For instance, Wiki_E denotes supplementary training using the dataset constructed through the Extract Method. The Supplementary Corpus of this dataset is collected from Wikipedia.
4.2 Results and Analysis
Figure 3: This figure shows the accuracy of various models using the CEM method on the Xiezhi and CMMLU tasks. The supplementary training set is constructed using the Normal Method.
CEM method enhances the performance of various LLMs across diverse tasks. We conduct tests on LLMs that undergo supplementary training using the CEM method. The training sets are constructed via the Normal Method. The accuracy results are presented in Fig. 3. More complete experimental results are in Appendix Table 4 and 5. After implementing CEM, all LLMs showed significant improvement in accuracy for both QA tasks, demonstrating the effectiveness of CEM in addressing knowledge deficiencies and enhancing LLMs’ performance. For instance, after one round of of CEM optimization with Bing as its supplementary corpus source, the accuracy of Qwen-7B-Chat on the CMMLU task surged by 16.08%.
Smaller LLMs can better benefit from CEM. The effectiveness of LLM tuning with CEM depends on the innate abilities of the LLM and also varies for LLMs with different parameter scales. According to the experimental results presented in Fig. 3, among the open-source models mentioned, ChatGLM-6B, which has the fewest parameters, shows the most significant improvement, up to 13.17% on the Xiezhi task. In contrast, CuteGPT-13B, which has the highest initial answer accuracy, exhibits a smaller improvement of up to 3.63% compared to the smaller LLM.
Extract Corpora and Review Corpora can significantly boost accuracy. We conducted parallel experiments on the supplementary training set construction method described in section 3.3, using CuteGPT-13B on Xiezhi dataset. The results are presented in Fig. 4. CEM with Extract Corpus outperforms its control group in all cases with higher Accuracy and W2R metric, e.g., Mix_E shows a 3.11% increase in accuracy to its control group Mix_EC, and a 10.96% increase compared to the baseline.
Figure 4: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplementary training set constructed using various methods. ’Avg of Ri’ and ’Best of Ri’ represent the average and best results, respectively, of the three Review Methods.
Meanwhile, in alignment with expectations, CEM with the Review Corpus ($C_{R_{i=1,2,3}}$), consistently outperform their control group across all instances and significantly reduce the R2W metric. For example, Wiki_R1 shows a 5.74% improvement in accuracy and a 3.70% reduction in R2W compared to its control group, Wiki_RC, and a 12.25% improvement in accuracy and a 5.89% reduction in R2W compared to the baseline. The results indicate that the W2R is improved by the Extract Corpus. This suggests that the model’s capacity to identify and comprehend the corresponding knowledge in the corpus is enhanced, leading to better error correction. Review Corpus significantly reduced R2W, indicating that the model’s forgetting of correct knowledge was suppressed.
In addition, the three construction methods, $R_{1}$,$R_{2}$ and $R_{3}$, are similar in their ability to enhance LLMs’ performance, which means either form of Review Corpus can be effective. Therefore, if not specified, $R$ will refer to the $R_{1}$ Method in the following.
Supplementary corpora from multiple sources do not always improve the CEM approach more than a single source. The effectiveness of CEM depends on the recall rate when searching corpora and the quality of supplementary corpora, as this determines whether the corpus contains knowledge that the LLM lacks. Moreover, noise and errors within low-quality corpora can negatively impact the LLM. Results in Fig. 4 show that there is no significant difference in the effectiveness of CEM, whether using Wikipedia or Bing as a single data source, or when using multiple data source concatenated. When using the supplementary training set constructed by Review Method, Mix_R has lower accuracy than Wiki_R and Bing_R despite having more data volumn, This may be due to the information redundancy caused by the high overlap of the supplementary corpus from different sources. Additionally, an excess of supplementary corpus may dilute the proportion of generalization, excerpt, and comment corpus, thereby hindering their effectiveness in reducing model catastrophic forgetting and enhancing model knowledge extraction.
Metrics Wiki Bing
N E R N E R
Acc Baseline 42.93 42.93 42.93 42.93 42.93 42.93
Round1 46.56 52.02 55.18 45.91 48.46 53.86
Round2 51.69 55.20 59.93 46.37 48.74 58.33
W2R comp. to Baseline 24.46 26.03 28.29 22.07 23.32 27.01
comp. to Round1 20.00 17.56 16.43 18.49 18.33 16.61
R2W comp. to Baseline 15.70 13.76 11.29 18.63 17.51 11.61
comp. to Round1 14.87 14.39 11.68 18.03 18.06 12.14
Table 3: We employ the identical training data construction method as in the first round to conduct a second round of supplementary training on the LLM. The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task.
CEM can continuously evolve the LLM through multiple iterations. We conduct a second round of supplementary training on CuteGPT-13B on Xiezhi task using the same dataset construction method and data source as the first round. The results are presented in Table 3. After the second round of supplementary training, the accuracy of LLM’s answer on both two QA tasks has significantly improved compared to the previous round. Both the Extract Method and Review Method are effective in improving the accuracy of the model. For instance, Wiki_R shows a 4.75% improvement in accuracy compared to its previous round, and a 17.00% improvement in accuracy compared to baseline.
Furthermore, the Review Method consistently exhibits significantly lower R2W metric compared to other supplementary dataset construction methods, e.g., the R2W metric relative to Baseline of Wiki_R decreases by 7.02% compared to Wiki_N and 5.90% compared to Wiki_E. We argue that the Review Method performs best because it significantly reduces the probability of changing a previously correct response to an incorrect one. This enables the model to be evolved by two or more rounds of supplemental training.
5 Conclusion
In this paper, we propose and validate the Continue Evolving from Mistakes (CEM) method, which assesses LLM’s knowledge deficiencies through LLM’s mistakes and targets supplementary LLM’s knowledge. We first devise a general process for supplementary training: (1) testing LLM and collecting mistakes, (2) collecting supplementary corpora from multiple data sources, (3) constructing a supplementary training set and fine-tuning the LLM. Furthermore, we introduce multiple methods for constructing training data to enhance the LLM’s capability to extract and comprehend knowledge within the corpora, and to mitigate the forgetting of LLM. We conduct experiments on several opensource LLMs to validate the efficacy of CEM. The results consistently show that CEM supplements the LLM’s knowledge deficiencies and improves the accuracy of the answers. For instance, on the Xiezhi task, the accuracy of the CuteGPT-13B improves by up to 12.25% in the first round and up to 17.00% in the second round.
Limitations: This paper focuses on the model’s performance on question-answering task. It is yet to be determined if CEM can consistently enhance the model’s other abilities and its performance in non-QA tasks.
In future research, we will further explore three or more rounds of supplementary training to improve the LLM and to mitigate the catastrophic forgetting. In addition, we will experiment with other data sources for corpora and optimize supplementary corpus collection strategies to equip the LLM with more advanced continual iterative evolutionary capabilities.
Appendix A Process of Collecting Corpora from Multiple Sources
Fig.5 and Fig.6 show the process of obtaining supplementary data from Wikipedia and Bing respectively.
Figure 5: The process of collecting corpora from Bing.
Figure 6: The process of collecting corpora from Wikipedia.
Appendix B Experimental Results of Various LLMs on QA tasks
Table 4 and Table 5 show the complete experimental results of LLMs using the CEM method on Xiezhi and CMMLU tasks. We use the Normal method to construct the supplementary training dataset.
Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
Baseline Wiki Bing Baseline Wiki Bing Baseline Wiki Bing
Acc 42.93 45.91 46.56 41.62 48.71 47.87 40.44 51.95 53.61
W2R - 22.20 21.91 - 22.17 20.94 - 26.01 25.86
R2W - 18.57 18.93 - 15.08 14.70 - 14.50 12.69
Table 4: The table shows the experimental results of LLMs using the CEM method on the Xiezhi task.
Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
Baseline Wiki Bing Baseline Wiki Bing Baseline Wiki Bing
Acc 26.44 32.30 35.32 24.72 39.12 40.80 23.00 35.62 35.26
W2R - 23.84 25.14 - 25.24 26.94 - 25.96 26.32
R2W - 17.98 16.26 - 11.04 11.02 - 13.34 14.06
Table 5: The table shows the experimental results of LLMs using the CEM method on the CMMLU task.
Appendix C Experimental Results with Various Supplementary Training Dataset Construction Methods
Table 6 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplementary training set constructed using various methods.
Data Source Metrics Baseline N E EC R1 R2 R3 RC
Wiki Acc 42.93 46.56 52.03 50.69 55.18 54.67 53.40 49.44
W2R - 22.20 24.12 23.42 24.93 24.61 24.82 22.90
R2W - 18.57 15.02 15.65 12.68 12.87 14.35 16.38
Bing Acc 42.93 45.91 48.47 45.11 53.86 54.53 53.90 46.58
W2R - 21.91 23.15 21.37 24.25 25.18 24.99 22.28
R2W - 18.93 17.61 19.18 13.32 13.58 14.02 18.62
Mix Acc 42.93 46.61 53.89 50.78 53.42 54.43 53.37 53.07
W2R - 21.99 25.12 23.67 24.04 25.06 25.00 24.81
R2W - 18.31 14.16 15.81 13.55 13.46 14.56 14.67
Table 6: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplementary training set construction methods.
Appendix D Data Division
Table 7 shows the the data division of Xiezhi and CMMLU. The multiple-choice questions are divided into three parts. The Pre-Finetuning part is utilized to train the model to answer questions in a standard format before supplementary training. The Testing part is used to construct the test set, and the remaining part is used to construct the Format Corpus and Extract Corpus for supplementary training.
Dataset Total Pre-Finetuning Testing Supplementary Training
Xiezhi 27,265 10,000 10,000 7,265
CMMLU 11,917 4,000 5,000 2,917
Table 7: The data division of Xiezhi and CMMLU.

### v3

@@
 Abstract
-Large Language Models (LLMs) demonstrate impressive performance in various downstream tasks.
-However, they may still generate incorrect responses in certain scenarios due to the knowledge deficiencies and the flawed pre-training data.
-Continual Learning (CL) is a commonly used method to address this issue.
-Traditional CL is task-oriented, using novel or factually accurate data to retrain LLMs from scratch.
-However, this method requires more task-related training data and incurs expensive training costs.
-To address this challenge, we propose the Continue Evolving from Mistakes (CEM) method, inspired by the ’summarize mistakes’ learning skill, to achieve iterative refinement of LLMs.
-Specifically, the incorrect responses of LLMs indicate knowledge deficiencies related to the questions.
-Therefore, we collect corpora with these knowledge from multiple data sources and follow it up with iterative supplementary training for continuous, targeted knowledge updating and supplementation.
-Meanwhile, we developed two strategies to construct supplementary training sets to enhance the LLM’s understanding of the corpus and prevent catastrophic forgetting.
-We conducted extensive experiments to validate the effectiveness of this CL method.
-In the best case, our method resulted in a 17.00% improvement in the accuracy of the LLM.
+As world knowledge evolves and new task paradigms emerge, Large Language Models (LLMs) often fall short of meeting new demands due to knowledge deficiencies and outdated information.
+Continual Learning (CL) is crucial for keeping LLMs up-to-date and addressing these deficiencies.
+However, traditional CL approaches struggle to balance task-width generality with task-depth specificity and often lack efficient data collection strategies, leading to increased training costs without addressing the model’s most critical needs.
+Inspired by the ‘summarizing mistakes’ learning skill, we propose the Continue Evolving from Mistakes (CEM) method.
+This iterative approach continually evaluates LLMs to identify knowledge deficiencies based on their mistakes, collecting relevant data from multiple sources to supplement training in a targeted manner.
+To enhance the model’s utilization of supplemental knowledge and prevent forgetting, we developed three dataset construction strategies that integrate various types of continual pretraining (CPT) data and continual instruction tuning (CIT) data.
+Extensive experiments demonstrate the efficacy of the CEM method, achieving up to a 17% improvement in LLM accuracy in the best scenarios.
+Additionally, further experiments confirm the potential of combining CEM with other catastrophic forgetting mitigation strategies, enabling multi-round iterative optimization.
 Large Language Model Can Continue Evolving From Mistakes
-Haokun Zhao, Haixia Han, Jie Shi, Chengyu Du, Jiaqing Liang, Yanghua Xiao Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University hkzhao23@m.fudan.edu.cn
+[…]
 1 Introduction
-Figure 1: Continual evaluation can help us uncover knowledge deficiency of the LLM and thus evolve the model through continuous learning.
-With the exponential expansion of both data and parameter size, large language models (LLMs) demonstrate outstanding performance across a series of downstream tasks Chang et al.
+With the exponential growth of data and parameter sizes, coupled with the employment of instruction fine-tuning, Large Language Models (LLMs) have showcased superior performance across a wide array of downstream tasks Chang et al.
+(2023); Wei et al.
+(2022); Peng et al.
 (2023).
-However, LLMs have the potential to generate incorrect outputs, a phenomenon often referred to as ’hallucination’ Huang et al.
+However, in real-world applications, the ongoing emergence of new knowledge and task requirements necessitates continual task-specific adaptation and knowledge supplementation for LLMs.
+Otherwise, LLMs are likely to generate incorrect outputs based on outdated or lacking knowledge (i.e., ‘hallucinations’) Huang et al.
 (2023).
-In addition, it is important to note that satisfactory answers generated by LLMs cannot be achieved within a single training session Han et al.
-(2024).
-Continual refinement and optimization processes are necessary.
-This optimization process ensures that the LLM can undergo iterative refinement, continuously enhancing its performance.
-There are two main reasons for the poor performance of LLM on specific questions.
-Firstly, some of these inquiries exceed the knowledge boundaries of LLM, indicating either a deficiency in domain expertise or outdated information within the LLM Ren et al.
+Accordingly, Continual Learning(CL) is proposed to ensure LLMs’ alignment with the ever-changing world knowledge.
+Given the significant training costs, it is more popular to continually pre-train or fine-tune LLMs on a sequence of emerging domain-specific or task-specific data, rather than periodically training them from scratch on the latest corpora (e.g., the most recently archived Wikipedia) Gururangan et al.
+(2020b).
+However, common CL tends to be task-width oriented, whereby the model is trained using new task-relevant data to enhance its applicability across a wider range of tasksGururangan et al.
+(2020b); Cossu et al.
+(2022), as opposed to focusing on the continual improvement of the model’s performance within a specific domain or task, i.e., task-depth oriented.
+In addition, building CL datasets is often challenging.
+Blind data collection fails to ensure access to the high-quality data that the model most requires for task-solving, and may introduce significant training costs due to irrelevant data.
+To address these issues, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide a data-efficient method for continually enhancing the performance of LLMs on specific tasks through muiltiple evaluation and supplementation with mistake-relevant knowledge.
+Our method draws inspiration from the ‘summarizing mistakes’ learning skill of human students.
+For instance, when a student is struggling with math, they can leverage test mistakes to identify lacking knowledge and extract relevant information from the textbook.
+Through repeated math tests and summaries, improvements in their math scores have been noted.
+While recent studies have uncovered successful utilization of mistakes in constructing CoT dataAn et al.
+(2024), thereby enhancing the model’s performance in mathematical reasoning, there exists potential to extend the leveraging of mistakes to diverse tasks and integrate it within CL paradigms.
+Recent studies have shown that the poor performance of LLM on specific tasks may result from two factors: (1) issues beyond its knowledge boundaries, indicating a lack of domain expertise Ren et al.
 (2023); Zhang et al.
-(2023).
-Secondly, the LLM may incorporate erroneous knowledge derived from potentially unreliable sources in the training data, resulting in the generation of inaccuracies or misconceptions that contradict factual information Rae et al.
+(2023b), and (2) knowledge errors within LLM, including those stemming from unreliable pre-training data sources and outdated information Rae et al.
 (2021); Zhao et al.
 (2023).
-To rectify the model’s knowledge deficits and ensure alignment with contemporary world knowledge, a common method is periodic training from scratch using the domain or latest corpus, such as the most recently archived Wikipedia.
-However, this method incurs significant training expenses Patterson et al.
+We argue that mistakes effectively expose the LLM’s knowledge deficiencies and errors.
+Addressing these directly represents the most efficient approach to enhancing the model.
+Therefore, at the start of each CEM round, we test the model and gather cases of incorrect or poor performance.
+We deconstruct these ‘bad cases’ to identify the most lacking or erroneous knowledge points, then obtain supplemental data from search engines and online encyclopedias to address these gaps.
+To enhance the LLM’s knowledge acquisition from this supplemental data and prevent forgetting, we devised multiple dataset construction strategies utilizing both unsupervised Continual Pre-training (CPT) data and supervised Continual Instruction Tuning (CIT) dataWu et al.
 (2024).
-Another method involves pre-training on a corpus that includes accurate information.
-However, this method may result in a decline in other aspects of the model’s performance Li et al.
+By integrating CEM with strategies to mitigate catastrophic forgetting, we enable iterative, continual model evolution.
+Our contributions are summarized as follows: (1) We propose the CEM method.
+This method explores knowledge boundaries from LLM’s mistakes, endowing LLMs with the capacity for continual iterative evolution.
+(2) To assist LLMs in acquiring and comprehending factual knowledge from supplemental corpora, and to mitigate knowledge forgetting, we propose three strategies for constructing supplemental training sets.
+(3) We conduct experiments on multiple open-source LLMs for two complex knowledge Question-Answering (QA) tasks.
+Across all scenarios, the CEM method consistently exhibit significant improvements, achieving an optimal accuracy enhancement of 17%.
+(4) We verify that CEM can be used together with methods for mitigating catastrophic forgetting Li et al.
+(2022), such as replay-based methods Chaudhry et al.
+(2019); Wu et al.
 (2021).
-This challenge has prompted research into mitigating catastrophic forgetting in continuous learning.
-Existing studies can be roughly grouped into three categories, e.g., experience replay methods Chaudhry et al.
-(2019); Wu et al.
-(2021), regularization-based methods Kirkpatrick et al.
-(2017a); Chen et al.
-(2022), and dynamic architecture methods Mallya et al.
-(2018).
-Recently, researchers have designed some hybrid methods that take advantage of the aforementioned techniques Chen et al.
-(2023); He et al.
-(2024).
-In this paper, we focus on the process of uncovering knowledge that the model has not yet mastered, for supplementation.
-Throughout each iterative phase of model evolution, we identify the limitations of the model’s knowledge and address them in subsequent updates.
-This idea is inspired by the ’summarizing mistakes’ learning skill observed in human students.
-An evaluation is conducted initially to identify a student’s weaknesses in a subject.
-References related to these errors are then provided to assist the student in mastering the associated knowledge.
-For instance, if a student is struggling with chemistry, they can review questions answered incorrectly in chemistry tests to pinpoint areas needing improvement and extract relevant knowledge, such as the properties of benzene, from the corresponding chapter in the chemistry textbook.
-This process enhances his performance in chemistry by identifying and learning from weaknesses or areas of unfamiliarity.
-In this paper, we propose the Continue Evolving from Mistakes (CEM) method.
-The aim is to systematically address the model’s knowledge gaps and rectify incorrect knowledge through continuous evolution.
-Our contributions are summarized as follows:
-  • To endow LLMs with the capacity for continuous iterative evolution, we have drawn inspiration from the ’summarize mistakes’ learning skill observed in human students.
-We introduce CEM, which explores the knowledge boundaries from LLM mistakes, thereby continuously enhancing the model’s capabilities.
-  • To assist LLMs in acquiring and comprehending factual knowledge from supplementary corpora, as well as mitigating knowledge forgetting, we propose two strategies for constructing supplementary corpora.
-  • We conduct experiments on three open-source LLMs and two complex knowledge Question-Answering (QA) tasks. The results demonstrate the efficacy of the CEM.
+This ensures that catastrophic forgetting can be suppressed even with multiple iterations.
 2 Related Work
@@
 The data used for fine-tuning is usually less than the original pre-training corpus.
-These datasets used in IFT are often curated through crowd-sourcing or extracted from an LM that is already capable of generating instruction-following examples.
+These datasets used in IFT are often curated through crowd-sourcing or extracted from an LM that is already capable of generating instruction-following examplesHan et al.
+(2024).
 Studies have shown that IFT can enhance zero-shot generalization to various unseen tasks.
@@
 (2022), prompt tuning Li and Liang (2021), and other techniques.
-Continual learning with LLM Continual learning (CL) aims to enable a model to learn from a continuous data stream across time while reducing catastrophic forgetting of previously acquired knowledge Biesialska et al.
+Continual Pre-training (CPT)
+Continual pretraining involves regularly updating LLMs with the latest facts or adapting them to specialized domains.
+To continually pre-train LLMs to update facts, researchers utilize dynamic datasets to assimilate real-time data from various sources such as news feeds Sun et al.
+(2019), scholarly articles Cossu et al.
+(2022), and social media Cossu et al.
 (2022).
-With CL, a deployed LLM has the potential to adapt to the changing world without costly re-training from scratch Bubeck et al.
+Continual pre-training enhances domain knowledge through two approaches: (1) domain-incremental pre-training, which accumulates knowledge across multiple domains, and (2) domain-specific continual learning, which refines a general model into a domain expert by training on domain-specific datasets and tasks.
+In domain-incremental pre-training, Cossu et al.
+(2022) continually pre-trained LLMs on new multimodal data streams, preparing them for various downstream tasks.
+Ke et al.
+(2023) introduces a soft-masking mechanism to update LLMs with domain corpora, aiming to boost performance while preserving general knowledge.
+Notably, there has been research on continuous pretraining for specific domains such as finance Xie et al.
+(2023) and e-commerce Ma et al.
 (2023).
-Early works Gururangan et al.
-(2020); Röttger and Pierrehumbert (2021); Lazaridou et al.
-(2021); Dhingra et al.
-(2022) empirically analyze continuing LM pre-training on emerging domain or temporal data, showing the potential to update the base LM with new knowledge.
-Jang et al.
-(2022b) explicitly categorize world knowledge as timeinvariant, outdated, and new knowledge, which should be retained, acquired, and updated respectively by an LM when learning continually.
-Jin et al.
-(2022); Jang et al.
-(2022a) additionally implement traditional CL methods to alleviate catastrophic forgetting, a phenomenon in which previously learned knowledge or abilities are degraded due to overwritten parameters Kirkpatrick et al.
-(2017b).
+Continual Instruction Tuning (CIT)
+Continual Instruction Tuning involves continually fine-tuning LLMs to learn how to follow instructions and transfer knowledge for future tasks Zhang et al.
+(2023a).
+It can be broadly classified into two categories: 1) task-incremental CIT and 2) domain-incremental CIT.
+Task-incremental CIT aims to continually fine-tune LLMs on a series of task-specific instructions to enable them to solve new tasks.
+A straightforward solution is to continuously generate instruction-tuning data for new tasks and directly fine-tune LLMs on it Wang et al.
+(2023).
+However, studies have shown this process would cause a catastrophic forgetting of the learned knowledge and problem-solving skills in previous tasks Kotha et al.
+(2024).
+Domain-incremental CIT aims to continually fine-tuning of LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains.
+TAPT Gururangan et al.
+(2020a) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain.
+AdaptLLM Cheng et al.
+(2024) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
 3 Methods
-The objective of this study is to enhance the knowledge base of LLM through continual learning, resulting in iterative improvements in model performance.
+The objective of this work is to enhance the knowledge of LLM through continual learning, resulting in iterative improvements in model performance.
 3.1 Task Formulation
@@
 First, the reasoning model $\mathcal{M}_{0}$ is utilized to answer all questions within the test set $\mathcal{Q}$. Subsequently, questions $q_{i}$ that receive incorrect responses are collected,
-$\widetilde{r_{i}}=\mathcal{M}_{0}(q_{i}),\quad Ans(\widetilde{r_{i}})\neq a_{i}$
+$\widetilde{r_{i}}=\mathcal{M}_{0}(q_{i}),\quad Ans(\widetilde{r_{i}})\neq a_{i},$
 where $\widetilde{r_{i}}$ denotes the response that includes incorrect answers to the question $q_{i}$, and $Ans(\cdot)$ is utilized to extract the final answer from the response.
-Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplementary corpus $C_{S_{i}}$,
-$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\dots\}$
-where $Search(\cdot)$ is used to search for relevant knowledge based on the question and retrieve supplementary corpus.
-Subsequently, based on the supplementary corpus $C_{S}$, construct the supplementary training set $D$ to fine-tune model $\mathcal{M}_{0}$,
-$D=Construct(C_{S}),\quad\mathcal{M}_{1}=Train(\mathcal{M}_{0},D)$
-where $Construct(\cdot)$ is utilized to build an supplementary training set, $Train(\cdot)$ denotes fine-tuning.
-$\mathcal{M}_{1}$, represents the model that has undergone one iteration of supplementary training.
-We can continue to employ $\mathcal{M}_{1}$ as the reasoning model on the test set $\mathcal{Q}$, for iterative supplementary training, thereby achieving continual optimization of the model’s performance.
+Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplemental corpus $C_{S_{i}}$,
+$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\dots\},$
+where $Search(\cdot)$ is used to search for relevant knowledge based on the question and obtain supplemental corpus.
+Subsequently, based on the supplemental corpus $C_{S}$, construct the supplemental training set $D$ to fine-tune model $\mathcal{M}_{0}$,
+$D=Construct(C_{S}),\quad\mathcal{M}_{1}=Train(\mathcal{M}_{0},D),$
+where $Construct(\cdot)$ is utilized to build an supplemental training set, $Train(\cdot)$ denotes fine-tuning.
+$\mathcal{M}_{1}$, represents the model that has undergone one iteration of supplemental training.
+We can continue to employ $\mathcal{M}_{1}$ as the reasoning model on the test set $\mathcal{Q}$, for iterative supplemental training, thereby achieving continual optimization of the model’s performance.
 3.2 CEM Method
-Specifically, the CEM method proposed in this paper involves three main steps: (1) testing the LLM and collecting incorrectly answered questions, (2) collecting supplementary corpora from multiple data sources, and (3) constructing a supplementary training set and fine-tuning the LLM.
-Fig.2 shows the process.
+We propose the CEM method, involving three main steps: (1) testing the LLM and collecting incorrectly answered questions, (2) collecting supplemental corpora from multiple data sources, and (3) constructing a supplemental training set and fine-tuning the LLM.
+Figure 1 shows the process.
 Figure 1: The pipeline of CEM method.
@@
 Subsequently, by utilizing pattern matching to extract answers from the LLM’s responses, we verify the accuracy of these answers and collect the questions that the LLM has answered incorrectly.
-Collecting Supplementary Corpora from Multiple Data Sources.
+Collecting Supplemental Corpora from Multiple Data Sources.
 We argue that the questions answered incorrectly by the LLM reveal knowledge deficiencies.
-Based on the implicit information contained within questions incorrectly answered by the LLM, we collect supplementary corpora from multiple data sources, including online encyclopedias and search engines.
-(1) Search Engine: We submit questions to search engines and use their retrieval techniques to crawl relevant content from the vast corpus of Chinese-language internet.
-Search engines are characterized by fast response time, high relevance matching, and diverse and rich content.
-We conduct experiments on two search engines, Baidu and Bing, and optimized our corpus crawling strategy.
-Specifically, we search for questions that the LLM answers incorrectly.
+Based on the implicit information contained within questions incorrectly answered by the LLM, we collect supplemental corpora from multiple data sources, including search engines and online encyclopedias.
+(1) Search Engine: Search engines are characterized by fast response time, high relevance matching, and diverse and rich content.
+We search for questions that the LLM answers incorrectly.
 The search engine ranks results based on relevance.
-We select the top-k results with the highest relevance and access their detailed pages to crawl the main textual content as supplementary corpora.
+We select the top-k results with the highest relevance and access their detailed pages to crawl the main textual content as supplemental corpora.
 To prevent the LLM from simply memorizing answers, we filter out search results that may contain both the original question and its answer.
-In this experiment, we utilize Bing as the search engine and set $k=4$, indicating that the main textual content from the detailed pages of the top-4 search results with the highest relevance was retrieved.
-Fig.5 shows process of collecting corpora from search engine.
+Additionally, we optimize the strategy for crawling the corpus to filter out irrelevant information and spam, such as advertisements.
+In subsequent experiments, we utilize Bing as the search engine and set $k=4$. Figure 4 in Appendix shows a case of corpus collection from Bing.
 (2) Online Encyclopedia: Compared to using search engines for retrieval, online encyclopedias provide more credible and professional search results, despite having lower hit rates than search engines.
-In our experiment, we experiment two online encyclopedias, Baidubaike and Wikipedia.
-Specifically, since the organization of encyclopedias typically relies on keyword entries, we conduct entity extraction from the questions and options of the collected incorrectly answered questions, using the extracted entity keywords for retrieval.
-The main text of the retrieved results serve as supplementary corpora.
-In our experiment, we use Wikipedia as the online encyclopedia and instruct ChatGPT-3.5 using a few-shot prompt for entity extraction.
-Appendix Fig.6 shows process of collecting corpora from online encyclopedia.
-Constructing Supplementary Training Set and Fine-Tuning LLM.
-The Supplementary Training Sets are constructed separately based on the Supplementary Corpus collected from Wikipedia and Bing.
+Since the organization of encyclopedias typically relies on keyword entries, we conduct entity extraction from the questions and options of the collected incorrectly answered questions, using the extracted entity keywords for retrieval.
+The main text of the retrieved results serve as supplemental corpora.
+In subsequent experiment, we use Wikipedia as the online encyclopedia and instruct GPT-3.5 using a few-shot prompt for entity extraction.
+Figure 5 in Appendix shows a case of corpus collection from Wikipedia.
+Constructing Supplemental Training Set and Fine-Tuning LLM.
+The supplemental training sets are constructed separately based on the supplemental corpus collected from Wikipedia and Bing.
 They are then employed to fine-tune the LLM.
 The construction method of the training set will be explained in detail in Section 3.3.
-After completing one round of supplementary training, the model is retested on the QA dataset to gather a new set of incorrectly answered questions, which serve as the first step for the next round of supplementary training.
-It should be noted that supplementary training should not be endless, as this may lead to severe catastrophic forgetting and the boundary effect of diminishing accuracy improvement.
-3.3 Supplementary Training Data
-Corpus Format Definition
-The construction of the supplementary training set involves the following four types of corpus.
-(1) Format Corpus: Table 1 shows an example of a Format Corpus.
-Prior to implementing the CEM method for supplementary training, we fine-tune the LLM using Format Corpora.
-This equips the LLM with the ability to respond to multiple-choice questions in a standard format, facilitating the extraction of answers from the LLM’s responses and allowing for the collection of incorrectly answered questions.
-To maintain this ability, we incorporate a certain proportion of Format Corpus during the supplementary training to prevent the LLM’s response skills from declining.
-Example 1: A Case of Format Corpus
+After completing one round of supplemental training, the model is retested on the QA dataset to evaluate its performance and gather a new set of incorrectly answered questions.
+This can serve as the first step for the next round of supplemental training.
+It is important to note that supplemental training should not be never-ending, as this may lead to severe catastrophic forgetting and the boundary effect of diminishing accuracy improvement.
+In the following experiments, we will examine how the number of supplemental training rounds affects model improvement.
+3.3 Supplemental Training Data
+Definition of Corpus Type. The construction of the supplemental training set involves the following five types of corpus.
+(1) Supplemental Corpus($C_{\underline{S}}$): Supplemental Corpus obtained from multiple data sources contains background knowledge that the LLM lacks and is used similarly to continuous pre-training Gupta et al.
+(2023).
+Example 1: A Case of Normative Corpus
 Question:
@@
 (answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
-Table 1: An example of a Format Corpus, designed to illustrate its corpus format.
-(2) Supplement Corpus: Corpus obtained through multiple data sources.
-(3) Extract Corpus: Table 2 shows an example of an Extract Corpus.
-Inspired by the human learning process: despite studying from the same textbooks, students who are better at extracting knowledge from these materials always achieve superior grades.
-We posit that merely providing the LLM with a supplementary corpus containing knowledge is inadequate.
-Consequently, we designed the Extract Corpus to train the LLM to effectively capture and comprehend knowledge from supplementary corpora, enhancing its utilization of these materials.
-Example 2: A Case of Extract Corpus
+Table 1: A case of a Normative Corpus, designed to illustrate its corpus format.
+Example 2: A Case of Extractive Corpus
 Question:
@@
 (answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
-Table 2: An example of a Extract Corpus, designed to illustrate its corpus format. The ellipsis indicates that some supplementary text has been omitted for the sake of clarity.
-(4) General Corpus: Replaying previous training data can effectively mitigate catastrophic forgetting when training LLM on new dataHe et al.
+Table 2: A case of a Extractive Corpus, designed to illustrate its corpus format. The ellipsis indicates that some supplemental text has been omitted for clarity.
+(2) Normative Corpus($C_{\underline{N}}$): To equip the LLM with the ability to respond to multiple-choice questions in a normative format, we sample multiple-choice questions that are independent of the test set and process them into the format shown in Table 1 to build Normative Corpus.
+Prior to supplemental training, we finetune the LLM using Normative Corpus(i.e.
+pre-finetuning), facilitating the extraction of answers from the LLM’s responses and allowing for the collection of incorrectly answered questions.
+To maintain this ability, we incorporate a small proportion of Normative Corpus during the supplemental training to prevent the LLM’s response skills from declining.
+(3) Extractive Corpus($C_{\underline{E}}$): Inspired by the human learning process: despite studying from the same textbooks, students who are better at extracting knowledge from these materials always achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
+To improve the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample multiple-choice questions that are independent of the test set and the Normative Corpus.
+These questions are processed into the format shown in Table 2 to build the Extractive Corpus.
+(4) Review Corpus($C_{\underline{R}}$): The responses of the LLM tend to be diverse under high temperature Caccia et al.
 (2018).
-Therefore, we utilize the training data from ShareGPT dataset  [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_
+During repeated QA testing, LLM may change its answers to certain ‘less confident’ questions due to the proximity of predicted probabilities.
+Consequently, after supplemental training, questions initially answered correctly may be answered incorrectly, thus affecting the accuracy.
+Therefore, to strengthen the LLM’s memory of the knowledge implied by correctly answered questions, we sample the responses provided correctly of the LLM in the previous round and process some of them in Normative Corpus format and the rest in Extractive Corpus format to build the Review Corpus.
+The composition of the Review Corpus can be described as:
+$C_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
+where $\alpha\in[0,1]$, denotes the proportion of responses processed into the Normative Corpus format.
+$r$ denotes the correct responses in the previous round of LLM.
+$Normative(\cdot)$ converts responses into the Normative Corpus format, and $Extractive(\cdot)$ converts responses into the Extractive Corpus format.
+(5) General Corpus($C_{\underline{G}}$): Replaying previous training data can effectively mitigate catastrophic forgetting when training LLM on new dataHe et al.
+(2021).
+Therefore, to preserve the model’s general capabilities on other tasks, we use the data from ShareGPT dataset  [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_
 unfiltered
-as General Corpus to preserve the model’s general capabilities.
-Supplementary Training Set Construction Methods.
-Through the combination of four types of corpora, we have developed three innovative methods for constructing supplementary training sets.
-These methods are detailed below.
-(1) Normal Method: To mitigate the potential catastrophic forgetting resulting from supplementary training, we mix the General Corpus with the Supplement Corpus, ensuring an equal quantity of those two type of corpus across all construction methods,
-$D_{N}=Construct(C_{G},C_{S})$
-where $Construct(\cdot)$ is utilized for constructing supplementary training set based on corpora, $C_{G}$ represents the General Corpus, $C_{S}$ indicates the Supplement Corpus, and $D_{N}$ denotes the supplementary training set constructed via the Normal Method.
-(2) Extract Method: To evaluate whether the Extract Corpus enhances the LLM’s ability to capture and comprehend knowledge from supplementary corpora as expected, and to associate questions with relevant knowledge when answering, we added the Extract Corpus into $D_{N}$ to generate the supplementary training set $D_{E}$.
-$D_{E}=Construct(C_{G},C_{S},C_{E})$
-We replaced the Extract Corpus with the General Corpus to obtain the control group $D_{E_{C}}$.
-$D_{E_{C}}=Construct(C_{G}^{\prime},C_{S})$
-(3) Review Method: Under higher temperature parameter, the responses of the LLM tend to be diverse.
-When conducting repeat QA tests on the LLM, it may change its answers to certain ’less confident’ questions due to the proximity of predicted probabilities.
-This means that after supplementary training, questions initially answered correctly may be answered incorrectly, resulting in a transition from right to wrong.
-Therefore, we strengthen the LLM’s memory of the knowledge implied by correctly answered questions by reviewing the responses provided correctly by the LLM in the previous round.
-Specifically, in this round of supplementary training, we sample the correct responses made by the LLM in the previous round and process the correctly answered questions into either ’Format Corpus’ or ’Extract Corpus’ formats.
-We define it as the Review Corpus, and based on the proportion of ’Format Corpus’ and ’Extract Corpus’ formats in the Review Corpus, we construct $R_{1}$, $R_{2}$, and $R_{3}$,
-$C_{R_{1}}=Format(\alpha S_{QA})+Extract((1-\alpha)S_{QA})$
-$C_{R_{2}}=Format(S_{QA})$
-$C_{R_{3}}=Extract(S_{QA})$
-where $\alpha\in(0,1)$, $S_{QA}$ denotes the QAs sampled from the correct responses in the previous round of LLM.
-$Format(\cdot)$ is employed to process the sampled QAs into Format Corpus format, while $Extract$ is used to process the sampled QAs into Extracted Corpus format.
-Note that although Review Corpus looks like a mixture of Extract Corpus and Format Corpus in its compositional form, the corpus of the former comes from the correct responses made by the model and is a proper subset of the test set.
-In contrast, the corpus used by Extract Corpus and Format Corpus is completely independent of the test set.
-We added the Review Corpus into the foundation of $D_{E}$ to generate the supplementary training set $D_{R_{i}}$,
-$D_{R_{i}}=Construct(C_{G},C_{S},C_{E},C_{R_{i}}),\quad i=1,2,3$
-And we replaced the Review Corpus with the General Corpus to obtain the control group $D_{R_{C}}$.
-$D_{R_{C}}=Construct(C_{G}^{\prime\prime},C_{S},C_{E}),$
-4 Experiments
-4.1 Experimental Settings
-Datasets. We conduct experiments on two question-answering datasets, including Xiezhi [^2] https://github.com/mikegu721/xiezhibenchmark and CMMLU [^3] https://github.com/haonan-li/CMMLU.
+as General Corpus.
+Supplemental Training Set Construction Strategy.
+We have developed three innovative strategies for constructing supplemental training sets by combining five types of corpus.
+The strategies are detailed below.
+(1) Plain Strategy: To mitigate the potential catastrophic forgetting and to maintain the LLM’s ability to provide normative responses, we mix three types of corpus to construct the supplemental training set:
+$D_{P}=Concat(C_{S},C_{N},C_{G}),$
+where $C_{S}$ denotes the Supplemental Corpus, $C_{N}$ denotes the Normative Corpus, $C_{G}$ denotes the General Corpus, $Concat{(\cdot)}$ merges various types of corpus together and $D_{P}$ denotes the supplemental training set constructed via the Plain Strategy.
+(2) Extractive Strategy: To enhance the LLM’s ability to capture and comprehend knowledge from supplemental material, we integrate the Extractive Corpus:
+$D_{E}=Concat(C_{S},C_{N},C_{G},C_{E}),$
+where $C_{E}$ denotes the Extractive Corpus and $D_{E}$ denotes the supplemental training set constructed via the Extractive Strategy.
+(3) Review Strategy: To enhance the model’s retention of acquired knowledge and prevent correct responses from being altered to incorrect ones over multiple iterations, we integrate the Review Corpus:
+$D_{R}=Concat(C_{S},C_{N},C_{G},C_{E},C_{R}),$
+where $C_{R}$ denotes the Review Corpus and $D_{R}$ denotes the supplemental training set constructed via the Review Strategy.
+4 Experimental Setup
+4.1 Datasets
+We conduct experiments on two question-answering datasets, including Xiezhi [^2] https://github.com/mikegu721/xiezhibenchmark and CMMLU [^3] https://github.com/haonan-li/CMMLU.
 Xiezhi, a comprehensive evaluation suite designed to assess holistic domain knowledge.
@@
 In our experiments, we select 27,265 multiple-choice questions from Xiezhi and 11,917 multiple-choice questions from CMMLU. Appendix Table 7 shows the details of the data division.
-Base language models.
-The objective of our work is to evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even those with smaller parameter scales.
-We select open-source models with parameter ranges from 6 to 13 billion for our experiment.
-We use CuteGPT-13B [^4] https://huggingface.co/XuYipei/kw-cutegpt-13b-base, ChatGLM-6B [^5] https://huggingface.co/THUDM/chatglm3-6b, Qwen-7B [^6] https://huggingface.co/Qwen/Qwen-7B-Chat as our instruction fine-tuning base models for CEM.
-These models have been fine-tuned on the instruction data enabling them to effectively follow instructions.
-CuteGPT is based on the original Llama model structure, expands the Chinese vocabulary and performs pre-trainingLiang et al.
-(2023).
-ChatGLM-6B also is an open bilingual language model.
-It is trained for about 1T tokens of Chinese and English corpusDu et al.
-(2022).
-Qwen-7B-Chat is a Transformer-based large language model, which is pretrained on a large volume of data, including web texts, books, codes, etcBai et al.
-(2023).
-In the instruction fine-tuning stage, CuteGPT-13B and Qwen-7B-chat employ full fine-tuning, and ChatGLM-6B employs Prompt-tuningLiu et al. (2022).
-Supplementary training dataset.
-We construct a variety of supplementary training sets and their respective control groups for the experiments following the method described in Section 3.3.
-To distinguish between the model using different supplementary training sets, we establish the following naming rules:
-(1) Baseline: it represents the model that has not undergone supplementary training and has only been pre-finetuned.
-(2) N: N denotes Normal, indicating that the model is supplementary trained using the supplementary training set constructed by the Normal Method.
-(3) E and EC: E denotes Extract, indicating that the model is supplementary trained using the supplementary training set constructed by the Extract Method.
-To obtain its respective control group, Extract Control (EC), we replace the Extract Corpus in the training set with the General Corpus.
-(4) R1, R2, R3 and RC: R denotes Review, indicating that the model is supplementary trained using the supplementary training set constructed by the Review Method and based on the proportion of ’Format Corpus’ and ’Extract Corpus’ formats in the Review Corpus, we further name them as R1, R2, and R3.
-To obtain their respective control group, Review Control (RC), we replace the Review Corpus in the training set with the General Corpus.
-In addition, we distinguish between the data sources from which the supplementary corpus is collected and establish the following naming rules:
-(1) Wiki: it represents the Supplementary Corpus of the supplementary training set used by the model is from Wikipedia.
-(2) Bing: it represents the Supplementary Corpus of the supplementary training set is from Bing.
-(3) Mix: it represents the Supplementary Corpus of the supplementary training set is collected from both Bing and Wikipedia and concatenated.
-Metrics. The effects of supplementary training will be evaluated using the following metrics:
-(1) ACC: it represents the accuracy of the answer of LLM.
-(2) W2R: it indicates the percentage of questions where the LLM changes incorrect answers to correct ones in the test set.
-(3) R2W: it indicates the percentage of questions where the LLM changes correct answers to incorrect ones in the test set.
-In Section 4.2, we combine the above two naming rules to describe the experimental results clearly.
-For instance, Wiki_E denotes supplementary training using the dataset constructed through the Extract Method.
-The Supplementary Corpus of this dataset is collected from Wikipedia.
-4.2 Results and Analysis
-Figure 3: This figure shows the accuracy of various models using the CEM method on the Xiezhi and CMMLU tasks. The supplementary training set is constructed using the Normal Method.
-CEM method enhances the performance of various LLMs across diverse tasks.
-We conduct tests on LLMs that undergo supplementary training using the CEM method.
-The training sets are constructed via the Normal Method.
-The accuracy results are presented in Fig.
-3.
-More complete experimental results are in Appendix Table 4 and 5.
-After implementing CEM, all LLMs showed significant improvement in accuracy for both QA tasks, demonstrating the effectiveness of CEM in addressing knowledge deficiencies and enhancing LLMs’ performance.
-For instance, after one round of of CEM optimization with Bing as its supplementary corpus source, the accuracy of Qwen-7B-Chat on the CMMLU task surged by 16.08%.
-Smaller LLMs can better benefit from CEM.
-The effectiveness of LLM tuning with CEM depends on the innate abilities of the LLM and also varies for LLMs with different parameter scales.
-According to the experimental results presented in Fig.
-3, among the open-source models mentioned, ChatGLM-6B, which has the fewest parameters, shows the most significant improvement, up to 13.17% on the Xiezhi task.
-In contrast, CuteGPT-13B, which has the highest initial answer accuracy, exhibits a smaller improvement of up to 3.63% compared to the smaller LLM.
-Extract Corpora and Review Corpora can significantly boost accuracy.
-We conducted parallel experiments on the supplementary training set construction method described in section 3.3, using CuteGPT-13B on Xiezhi dataset.
-The results are presented in Fig.
-4.
-CEM with Extract Corpus outperforms its control group in all cases with higher Accuracy and W2R metric, e.g., Mix_E shows a 3.11% increase in accuracy to its control group Mix_EC, and a 10.96% increase compared to the baseline.
-Figure 4: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplementary training set constructed using various methods.
-’Avg of Ri’ and ’Best of Ri’ represent the average and best results, respectively, of the three Review Methods.
-Meanwhile, in alignment with expectations, CEM with the Review Corpus ($C_{R_{i=1,2,3}}$), consistently outperform their control group across all instances and significantly reduce the R2W metric.
-For example, Wiki_R1 shows a 5.74% improvement in accuracy and a 3.70% reduction in R2W compared to its control group, Wiki_RC, and a 12.25% improvement in accuracy and a 5.89% reduction in R2W compared to the baseline.
-The results indicate that the W2R is improved by the Extract Corpus.
-This suggests that the model’s capacity to identify and comprehend the corresponding knowledge in the corpus is enhanced, leading to better error correction.
-Review Corpus significantly reduced R2W, indicating that the model’s forgetting of correct knowledge was suppressed.
-In addition, the three construction methods, $R_{1}$,$R_{2}$ and $R_{3}$, are similar in their ability to enhance LLMs’ performance, which means either form of Review Corpus can be effective.
-Therefore, if not specified, $R$ will refer to the $R_{1}$ Method in the following.
-Supplementary corpora from multiple sources do not always improve the CEM approach more than a single source.
-The effectiveness of CEM depends on the recall rate when searching corpora and the quality of supplementary corpora, as this determines whether the corpus contains knowledge that the LLM lacks.
+4.2 Base Language Models
+To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs with parameter ranges from 6 to 13 billion.
+Our instruction fine-tuning base models include CuteGPT-13B [^4] https://huggingface.co/XuYipei/kw-cutegpt-13b-base, ChatGLM-6B [^5] https://huggingface.co/THUDM/chatglm3-6b, and Qwen-7B-Chat [^6] https://huggingface.co/Qwen/Qwen-7B-Chat.
+These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
+CuteGPT-13B is based on the original Llama model structure, expands the Chinese vocabulary and performs pre-trainingLiang et al. (2023).
+ChatGLM-6B is an open bilingual language model. It is trained for about 1T tokens of Chinese and English corpusDu et al. (2022).
+Qwen-7B-Chat is a Transformer-based large language model, which is pretrained on a large volume of data, including web texts, books, codes, etcBai et al. (2023).
+4.3 Metrics
+The effectiveness of supplemental training will be assessed using the following metrics:
+ACC: it represents the accuracy of the answer of LLM.
+W2R: it indicates the percentage of questions where the LLM changes incorrect answers to correct ones on the test set.
+R2W: it indicates the percentage of questions where the LLM changes correct answers to incorrect ones on the test set.
+5 Experiments
+We construct various supplemental training sets and their corresponding control groups for the experiments, following the method described in Section 3.3.
+To differentiate between models trained on different supplemental training sets, we establish the following naming conventions:
+INIT: it represents the model, which has undergone only pre-finetuning using the Normative Corpus, without any supplemental training.
+CEM-N: N denotes Normative, representing the model that has undergone supplemental training using the training set constructed by the Plain Strategy.
+[…]
+[…]
+We append suffixes to the aforementioned naming conventions to differentiate the data sources of the supplemental corpus.
+These suffixes include: ‘Wiki’, ‘Bing’ and ‘Mix’ where ‘Mix’ indicates that the Supplemental Corpus is collected from both Bing and Wikipedia and concatenated.
+For instance, ‘E_Wiki’ denotes the model supplemented trained using the dataset constructed through the Extractive Strategy.
+The Supplemental Corpus of this dataset is collected from Wikipedia.
+5.1 Main Results
+We conduct experiments to analyze the accuracy after applying CEM on the test data. The results are shown in Figure 2. Our observations are summarized as follows:
+Figure 2: This figure shows the accuracy of various models using the CEM method on the Xiezhi and CMMLU tasks. The supplemental training set is constructed using the Plain Strategy.
+Firstly, following the supplemental training, a noticeable enhancement in accuracy is observed across all models for the two given tasks.
+This implies that despite the high availability of supplemental data and the small size of the training set, the CEM method effectively complements the model’s knowledge deficiencies.
+For instance, the ChatGLM-6B model shows an improvement of up to 16.08% on the CMMLU task.
+Secondly, the effectiveness of tuning LLMs with CEM depends on their innate abilities and also varies across LLMs with different parameter scales.
+Among the mentioned open-source models, ChatGLM-6B, which has the fewest parameters, exhibits the most significant improvement, up to 13.17% on the Xiezhi task.
+In contrast, CuteGPT-13B, with the highest initial answer accuracy and the most parameters, shows a smaller improvement of up to 3.63% compared to the smaller LLM.
+Construction Strategy Analysis.
+To investigate the effectiveness of each component in supplemental training set, we conduct experiments on LLMs trained using various dataset construction strategies, the results are shown in Figure 3.
+Our observations are summarized as follows:
+Figure 3: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplemental training set constructed using various strategies.
+’Avg of $R_{\alpha}$’ and ’Best of $R_{\alpha}$’ represent the average and best results, respectively, of the Review Strategy with different $\alpha$.
+Firstly, Extractive Corpus and Review Corpus can significantly boost LLMs’ performance.
+CEM with Extractive Corpus ($C_{E}$) outperforms its control group in all cases with higher Accuracy and W2R metric, e.g., E_Mix shows a 3.11% increase in accuracy to its control group E’_Mix, and a 10.96% increase compared to INIT.
+Meanwhile, in alignment with expectations, CEM with the Review Corpus ($C_{R_{\alpha=0,0.5,1}}$), consistently outperform their control group across all instances and significantly reduce the R2W metric.
+For example, R0.5_Wiki shows a 5.74% improvement in accuracy and a 3.70% reduction in R2W compared to its control group, R’_Wiki, and a 12.25% improvement in accuracy and a 5.89% reduction in R2W compared to the INIT.
+The results indicate an improvement in W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the corpus.
+This enhancement leads to more effective error correction.
+Additionally, the Review Corpus significantly reduced R2W, indicating suppression of the model’s tendency to forget correct knowledge.
+Secondly, Supplemental Corpus from multiple sources do not always improve the CEM method more than a single source.
+The effectiveness of CEM depends on the recall rate when searching corpus and the quality of supplemental corpus, as this determines whether the corpus contains knowledge that the LLM lacks.
 Moreover, noise and errors within low-quality corpora can negatively impact the LLM.
-Results in Fig.
-4 show that there is no significant difference in the effectiveness of CEM, whether using Wikipedia or Bing as a single data source, or when using multiple data source concatenated.
-When using the supplementary training set constructed by Review Method, Mix_R has lower accuracy than Wiki_R and Bing_R despite having more data volumn, This may be due to the information redundancy caused by the high overlap of the supplementary corpus from different sources.
-Additionally, an excess of supplementary corpus may dilute the proportion of generalization, excerpt, and comment corpus, thereby hindering their effectiveness in reducing model catastrophic forgetting and enhancing model knowledge extraction.
+Results in Figure 3 show that there is no significant difference in the effectiveness of CEM, whether using Wikipedia or Bing as a single data source, or when using multiple data source concatenated.
+When using the supplemental training set constructed by Review Strategy, R_Mix has lower accuracy than R_Wiki and R_Bing despite having more data volumn, This may be due to the information redundancy caused by the high overlap of the supplemental corpus from different sources.
+Additionally, an excess of supplemental corpus may dilute the proportion of General, Extractive, and Normative Corpus, thereby hindering their effectiveness in reducing model catastrophic forgetting and enhancing model knowledge extraction.
+Thirdly, proportion factor $\alpha$ has limited impact on the effectiveness of the Review Corpus.
+According to Table 6 in Appendix, R0, R0.5 and R1 are similar in their ability to enhance LLMs’ performance.
+This may be because the models themselves have already mastered the knowledge, and replaying it in either format serves to consolidate that knowledge.
+Therefore, if not specified, R will refer to the R0.5 Strategy in the following.
+5.2 Multiple Iterations Analysis
 Metrics Wiki Bing
-N E R N E R
-Acc Baseline 42.93 42.93 42.93 42.93 42.93 42.93
+CEM-N CEM-E CEM-R CEM-N CEM-E CEM-R
+Acc INIT 42.93 42.93 42.93 42.93 42.93 42.93
 Round1 46.56 52.02 55.18 45.91 48.46 53.86
 Round2 51.69 55.20 59.93 46.37 48.74 58.33
-W2R comp. to Baseline 24.46 26.03 28.29 22.07 23.32 27.01
-comp. to Round1 20.00 17.56 16.43 18.49 18.33 16.61
-R2W comp. to Baseline 15.70 13.76 11.29 18.63 17.51 11.61
-comp. to Round1 14.87 14.39 11.68 18.03 18.06 12.14
-Table 3: We employ the identical training data construction method as in the first round to conduct a second round of supplementary training on the LLM.
+W2R vs. INIT 24.46 26.03 28.29 22.07 23.32 27.01
+vs. Round1 20.00 17.56 16.43 18.49 18.33 16.61
+R2W vs. INIT 15.70 13.76 11.29 18.63 17.51 11.61
+vs. Round1 14.87 14.39 11.68 18.03 18.06 12.14
+Table 3: We employ the identical training data construction strategy as in the first round to conduct a second round of supplementary training on the LLM.
 The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task.
-CEM can continuously evolve the LLM through multiple iterations.
-We conduct a second round of supplementary training on CuteGPT-13B on Xiezhi task using the same dataset construction method and data source as the first round.
+We conduct a second round of supplemental training on CuteGPT-13B on Xiezhi task using the same dataset construction strategy and data source as the first round.
 The results are presented in Table 3.
 After the second round of supplementary training, the accuracy of LLM’s answer on both two QA tasks has significantly improved compared to the previous round.
-Both the Extract Method and Review Method are effective in improving the accuracy of the model.
-For instance, Wiki_R shows a 4.75% improvement in accuracy compared to its previous round, and a 17.00% improvement in accuracy compared to baseline.
-Furthermore, the Review Method consistently exhibits significantly lower R2W metric compared to other supplementary dataset construction methods, e.g., the R2W metric relative to Baseline of Wiki_R decreases by 7.02% compared to Wiki_N and 5.90% compared to Wiki_E.
-We argue that the Review Method performs best because it significantly reduces the probability of changing a previously correct response to an incorrect one.
+Both the Extractive Strategy and Review Strategy are effective in improving the accuracy of the model.
+For instance, R_Wiki shows a 4.75% improvement in accuracy compared to its previous round, and a 17.00% improvement in accuracy compared to INIT.
+Furthermore, the Review Strategy consistently exhibits significantly lower R2W metric compared to other supplementary dataset construction strategies, e.g., the R2W metric relative to INIT of Wiki_R decreases by 7.02% compared to Wiki_N and 5.90% compared to Wiki_E.
+We argue that the Review Strategy performs best because it significantly reduces the probability of changing a previously correct response to an incorrect one.
 This enables the model to be evolved by two or more rounds of supplemental training.
+6 Discussion
+Why can the Extractive Corpus be effective?
+We observe that models employing the Extractive Corpus exhibit higher accuracy and W2R metric.
+Unlike the Review Corpus, where improving model performance is intuitive, the precise mechanism by which the Extractive Corpus augments the model’s capability to capture pivotal knowledge in lengthy texts remains worthy of discussion.
+Extractive Corpus provides the model with both questions and supplemental material containing relevant knowledge.
+Questions and options typically present straightforwardly and can be condensed into several key points, whereas the supplemental material tends to intricately interweave numerous factual statements Jiang et al.
+(2024).
+Thus, we suggest that questions may be a structured induction of supplemental material, where the Extractive Corpus explicitly links complex information within the corpus to the resulting induction (i.e., the question and options).
+Questions and options can act as cues and guides for the model to gather relevant information from the corpus, considering how this knowledge is accessed through questions.
+This, in turn, can enhance the model’s ability to extract useful information from the corpus while eliminating redundant and irrelevant information.
 7 Conclusion
 In this paper, we propose and validate the Continue Evolving from Mistakes (CEM) method, which assesses LLM’s knowledge deficiencies through LLM’s mistakes and targets supplementary LLM’s knowledge.
-We first devise a general process for supplementary training: (1) testing LLM and collecting mistakes, (2) collecting supplementary corpora from multiple data sources, (3) constructing a supplementary training set and fine-tuning the LLM.
-Furthermore, we introduce multiple methods for constructing training data to enhance the LLM’s capability to extract and comprehend knowledge within the corpora, and to mitigate the forgetting of LLM.
-We conduct experiments on several opensource LLMs to validate the efficacy of CEM.
+We first devise a general process for supplementary training: (1) testing LLM and collecting mistakes, (2) collecting supplemental corpora from multiple data sources, (3) constructing a supplementl training set and fine-tuning the LLM.
+Furthermore, we introduce multiple strategies for constructing training data to enhance the LLM’s capability to extract and comprehend knowledge within the corpora, and to mitigate the forgetting of LLM.
+We conduct experiments on several open-source LLMs to validate the efficacy of CEM.
 The results consistently show that CEM supplements the LLM’s knowledge deficiencies and improves the accuracy of the answers.
 For instance, on the Xiezhi task, the accuracy of the CuteGPT-13B improves by up to 12.25% in the first round and up to 17.00% in the second round.
-Limitations: This paper focuses on the model’s performance on question-answering task.
-It is yet to be determined if CEM can consistently enhance the model’s other abilities and its performance in non-QA tasks.
-In future research, we will further explore three or more rounds of supplementary training to improve the LLM and to mitigate the catastrophic forgetting.
+8 Limitations
+This paper focuses on the model’s performance on question-answering task. It is yet to be determined if CEM can consistently enhance the model’s other abilities and its performance in non-QA tasks.
+In future research, we will further explore three or more rounds of supplemental training to improve the LLM and to mitigate the catastrophic forgetting.
 In addition, we will experiment with other data sources for corpora and optimize supplementary corpus collection strategies to equip the LLM with more advanced continual iterative evolutionary capabilities.
 Appendix A Process of Collecting Corpora from Multiple Sources
-Fig.5 and Fig.6 show the process of obtaining supplementary data from Wikipedia and Bing respectively.
+Fig.4 and Fig.5 show the process of obtaining supplemental data from Wikipedia and Bing respectively.
 Figure 4: The process of collecting corpora from Bing.
@@
 Appendix B Experimental Results of Various LLMs on QA tasks
-Table 4 and Table 5 show the complete experimental results of LLMs using the CEM method on Xiezhi and CMMLU tasks. We use the Normal method to construct the supplementary training dataset.
+Table 4 and Table 5 show the complete experimental results of LLMs using the CEM method on Xiezhi and CMMLU tasks. We use the Plain Strategy to construct the supplemental training dataset.
 Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
-Baseline Wiki Bing Baseline Wiki Bing Baseline Wiki Bing
+INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
 Acc 42.93 45.91 46.56 41.62 48.71 47.87 40.44 51.95 53.61
@@
 Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
-Baseline Wiki Bing Baseline Wiki Bing Baseline Wiki Bing
+INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
 Acc 26.44 32.30 35.32 24.72 39.12 40.80 23.00 35.62 35.26
@@
 Table 5: The table shows the experimental results of LLMs using the CEM method on the CMMLU task.
-Appendix C Experimental Results with Various Supplementary Training Dataset Construction Methods
-Table 6 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplementary training set constructed using various methods.
-Data Source Metrics Baseline N E EC R1 R2 R3 RC
+Appendix C Experimental Results with Various supplemental Training Dataset Construction Strategies
+Table 6 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplemental training set constructed using various strategies.
+Data Metrics INIT CEM-N CEM-E CEM-E’ CEM-R0.5 CEM-R1 CEM-R0 CEM-R’
 Wiki Acc 42.93 46.56 52.03 50.69 55.18 54.67 53.40 49.44
@@
 R2W - 18.31 14.16 15.81 13.55 13.46 14.56 14.67
-Table 6: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplementary training set construction methods.
+Table 6: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplemental training set construction methods.
 Appendix D Data Division
@@
 The multiple-choice questions are divided into three parts.
-The Pre-Finetuning part is utilized to train the model to answer questions in a standard format before supplementary training.
-The Testing part is used to construct the test set, and the remaining part is used to construct the Format Corpus and Extract Corpus for supplementary training.
-Dataset Total Pre-Finetuning Testing Supplementary Training
+The Pre-Finetuning part is utilized to train the model to answer questions in a standard format before supplemental training.
+The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
+Dataset Total Pre-Finetuning Testing supplemental Training
 Xiezhi 27,265 10,000 10,000 7,265

### v4

@@
 Abstract
-As world knowledge evolves and new task paradigms emerge, Large Language Models (LLMs) often fall short of meeting new demands due to knowledge deficiencies and outdated information.
-Continual Learning (CL) is crucial for keeping LLMs up-to-date and addressing these deficiencies.
-However, traditional CL approaches struggle to balance task-width generality with task-depth specificity and often lack efficient data collection strategies, leading to increased training costs without addressing the model’s most critical needs.
-Inspired by the ‘summarizing mistakes’ learning skill, we propose the Continue Evolving from Mistakes (CEM) method.
-This iterative approach continually evaluates LLMs to identify knowledge deficiencies based on their mistakes, collecting relevant data from multiple sources to supplement training in a targeted manner.
-To enhance the model’s utilization of supplemental knowledge and prevent forgetting, we developed three dataset construction strategies that integrate various types of continual pretraining (CPT) data and continual instruction tuning (CIT) data.
-Extensive experiments demonstrate the efficacy of the CEM method, achieving up to a 17% improvement in LLM accuracy in the best scenarios.
-Additionally, further experiments confirm the potential of combining CEM with other catastrophic forgetting mitigation strategies, enabling multi-round iterative optimization.
+As world knowledge evolves and new task paradigms emerge, Continual Learning (CL) is crucial for keeping Large Language Models (LLMs) up-to-date and addressing their shortcomings.
+In practical applications, LLMs often require both continual instruction tuning (CIT) and continual pre-training (CPT) to adapt to new task paradigms and acquire necessary knowledge for task-solving.
+However, it remains challenging to collect CPT data that addresses the knowledge deficiencies in models while maintaining adequate volume, and improving the efficiency of utilizing this data also presents significant difficulties.
+Inspired by the ’summarizing mistakes’ learning skill, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide a data-efficient approach for collecting CPT data and continually improving LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge.
+To efficiently utilize these CPT data and mitigate forgetting, we design a novel CL training set construction paradigm that integrates parallel CIT and CPT data.
+Extensive experiments demonstrate the efficacy of the CEM method, achieving up to a 17% improvement in accuracy in the best case.
+Furthermore, additional experiments confirm the potential of combining CEM with catastrophic forgetting mitigation methods, enabling iterative and continual model evolution.
 Large Language Model Can Continue Evolving From Mistakes
-[…]
+Haokun Zhao1, Haixia Han2, Jie Shi1, Chengyu Du 1, Jiaqing Liang3 thanks: Corresponding Author, Yanghua Xiao 1,2 Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University1 Shanghai Institute of Artificial Intelligence for Education, East China Normal University2 Shanghai Key Laboratory of Data Science, School of Data Science, Fudan University3 {hkzhao23,jshi22}@m.fudan.edu.cn, {haixiahan03,cydu2024}@gmail.com, {liangjiaqing,shawyh}@fudan.edu.cn
 1 Introduction
-With the exponential growth of data and parameter sizes, coupled with the employment of instruction fine-tuning, Large Language Models (LLMs) have showcased superior performance across a wide array of downstream tasks Chang et al.
+With the exponential growth of data and parameter sizes, Large Language Models (LLMs) have demonstrated superior performance across numerous downstream tasks Chang et al.
 (2023); Wei et al.
@@
 (2023).
-However, in real-world applications, the ongoing emergence of new knowledge and task requirements necessitates continual task-specific adaptation and knowledge supplementation for LLMs.
-Otherwise, LLMs are likely to generate incorrect outputs based on outdated or lacking knowledge (i.e., ‘hallucinations’) Huang et al.
+However, in real-world applications, the continual emergence of new knowledge and task requirements necessitates ongoing task adaptation and knowledge supplementation for LLMs.
+Otherwise, LLMs may fall into hallucinations and generate incorrect responses Huang et al.
 (2023).
-Accordingly, Continual Learning(CL) is proposed to ensure LLMs’ alignment with the ever-changing world knowledge.
-Given the significant training costs, it is more popular to continually pre-train or fine-tune LLMs on a sequence of emerging domain-specific or task-specific data, rather than periodically training them from scratch on the latest corpora (e.g., the most recently archived Wikipedia) Gururangan et al.
+Therefore, Continual Learning (CL) is proposed to ensure LLMs’ alignment with the evolving world knowledge and demands.
+Figure 1: The figure illustrates two potential triggers for poor model performance: (1) Task Paradigm Unfamiliarity, and (2) Lack of task-relevant knowledge.
+Unfamiliarity with the task paradigm can cause deviations from expected interaction styles, while insufficient task knowledge may lead to hallucinations.
+Instruction tuning addresses only the former.
+Wu et al.
+(2024) categorize Continual Learning into Continual Pre-training (CPT) and Continual Instruction Tuning (CIT).
+CPT requires a larger volume of training data to enrich LLMs’ knowledge for new domains, while CIT continually fine-tunes LLMs on emerging task-specific instruction data to adapt the models to evolving task paradigms with much smaller training costs and data requirements Razdaibiedina et al.
+(2023); Gururangan et al.
+(2020b); Cossu et al.
+(2022).
+However, recent work has highlighted the limitations of instruction tuning, suggesting it is merely a superficial alignment process focused on learning interaction styles and LLMs primarily acquire knowledge during the pre-training stage Zhou et al.
+(2023).
+This will trigger the bad response shown in Figure 1, where LLMs’ poor performance may result from a lack of requisite knowledge to solve the task rather than unfamiliarity with the task paradigm Ren et al.
+(2023); Zhang et al.
 (2023b).
-However, common CL tends to be task-width oriented, whereby the model is trained using new task-relevant data to enhance its applicability across a wider range of tasksGururangan et al.
-(2020b); Cossu et al.
-(2022), as opposed to focusing on the continual improvement of the model’s performance within a specific domain or task, i.e., task-depth oriented.
-In addition, building CL datasets is often challenging.
-Blind data collection fails to ensure access to the high-quality data that the model most requires for task-solving, and may introduce significant training costs due to irrelevant data.
-To address these issues, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide a data-efficient method for continually enhancing the performance of LLMs on specific tasks through muiltiple evaluation and supplementation with mistake-relevant knowledge.
-Our method draws inspiration from the ‘summarizing mistakes’ learning skill of human students.
-For instance, when a student is struggling with math, they can leverage test mistakes to identify lacking knowledge and extract relevant information from the textbook.
-Through repeated math tests and summaries, improvements in their math scores have been noted.
-While recent studies have uncovered successful utilization of mistakes in constructing CoT dataAn et al.
-(2024), thereby enhancing the model’s performance in mathematical reasoning, there exists potential to extend the leveraging of mistakes to diverse tasks and integrate it within CL paradigms.
-Recent studies have shown that the poor performance of LLM on specific tasks may result from two factors: (1) issues beyond its knowledge boundaries, indicating a lack of domain expertise Ren et al.
-(2023); Zhang et al.
-(2023b), and (2) knowledge errors within LLM, including those stemming from unreliable pre-training data sources and outdated information Rae et al.
-(2021); Zhao et al.
+CPT is more effective than CIT in addressing this issue.
+However, using CPT to supplement the model with the lacking knowledge required for task-solving presents several issues: First, collecting CPT data is challenging, as it is difficult to a priori identify the lacking knowledge while ensuring a sufficiently large data scale.
+Blind data collection and insufficient data volume may introduce irrelevant information, weaken generalization, and pose overfitting risks Zhang et al.
+(2024); Gekhman et al.
 (2024).
-We argue that mistakes effectively expose the LLM’s knowledge deficiencies and errors.
-Addressing these directly represents the most efficient approach to enhancing the model.
-Therefore, at the start of each CEM round, we test the model and gather cases of incorrect or poor performance.
-We deconstruct these ‘bad cases’ to identify the most lacking or erroneous knowledge points, then obtain supplemental data from search engines and online encyclopedias to address these gaps.
-To enhance the LLM’s knowledge acquisition from this supplemental data and prevent forgetting, we devised multiple dataset construction strategies utilizing both unsupervised Continual Pre-training (CPT) data and supervised Continual Instruction Tuning (CIT) dataWu et al.
+Second, due to the weak constraints of self-supervised CPT data during parameters updates and the scale gap compared to initial pre-training data, efficiently utilizing CPT data and avoiding catastrophic forgetting remain difficult.
+To address these issues, recent work has explored converting CPT data into CIT data, such as transforming it into reading comprehension questions to avoid the problem of insufficient data volume.
+Cheng et al.
+(2024) or using highly relevant QA pairs for CIT before CPT to enhance the data utilization efficiency Jiang et al.
 (2024).
-By integrating CEM with strategies to mitigate catastrophic forgetting, we enable iterative, continual model evolution.
-Our contributions are summarized as follows: (1) We propose the CEM method.
-This method explores knowledge boundaries from LLM’s mistakes, endowing LLMs with the capacity for continual iterative evolution.
-(2) To assist LLMs in acquiring and comprehending factual knowledge from supplemental corpora, and to mitigate knowledge forgetting, we propose three strategies for constructing supplemental training sets.
-(3) We conduct experiments on multiple open-source LLMs for two complex knowledge Question-Answering (QA) tasks.
-Across all scenarios, the CEM method consistently exhibit significant improvements, achieving an optimal accuracy enhancement of 17%.
-(4) We verify that CEM can be used together with methods for mitigating catastrophic forgetting Li et al.
-(2022), such as replay-based methods Chaudhry et al.
+However, these methods demand high-quality data and the assistance of oracle LLM, and they overlook the crucial challenge of CPT mentioned above.
+Inspired by the ‘summarizing mistakes’ learning skill of human students, we argue that a model’s mistakes in ’exams’ reflect its inherent knowledge deficiencies.
+CPT data that directly addresses these gaps is the most effective for supplementing the model’s knowledge.
+Currently, the successful utilization of mistakes in constructing CoT data An et al.
+(2024); Tong et al.
+(2024) has enhanced the model’s performance in mathematics and reasoning, indicating the potential to leverage mistakes in continual learning.
+In this paper, we propose the Continue Evolving from Mistakes(CEM) method, aiming to provide a data-efficient method for collecting CPT data and continually enhancing the performance of LLMs through multiple evaluation and supplementation with mistake-relevant knowledge.
+Specifically, at the start of each CEM round, we test the model, identify cases of poor performance, and deconstruct them to pinpoint knowledge deficiencies.
+We then gather abundant supplemental data based on mistake analysis from search engines and online encyclopedias as CPT data to address these gaps.
+To efficiently utilize these CPT data and mitigate forgetting, we designed a novel CL training set construction paradigm that integrates parallel CIT and CPT data.
+By integrating CEM with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al.
+(2022), we enable iterative and continual model evolution.
+Our contributions are summarized as follows: (1) We propose the CEM method, which identifies knowledge deficiencies from LLMs’ mistakes and collects abundant relevant CPT data, enabling continual iterative evolution.
+(2) To assist LLMs in acquiring and comprehending factual knowledge from CPT data and mitigating knowledge forgetting, we propose three strategies for constructing supplemental training set using a parallel paradigm with CIT and CPT data.
+(3) We conduct experiments on multiple open-source LLMs for two complex knowledge QA tasks.
+The CEM method consistently demonstrates significant improvements across all scenarios, , achieving up to a 17.00% accuracy enhancement.
+(4) We verify that CEM can be combined with methods to mitigate catastrophic forgetting, such as replay-based methods Chaudhry et al.
 (2019); Wu et al.
+(2021), ensuring suppression of catastrophic forgetting over multiple iterations.
+2 Related Work
+Continual Pre-training (CPT)
+Continual pre-training involves regularly updating LLMs with the latest facts or adapting them to specialized domains Wu et al.
 (2024).
-This ensures that catastrophic forgetting can be suppressed even with multiple iterations.
-2 Related Work
-Instruction fine-tuning LMs often struggle to maintain alignment well with human intent after pre-training.
-However, this issue can be alleviated by employing techniques such as Instruction Fine-Tuning (IFT) Ouyang et al.
-(2022); Stiennon et al.
-(2022) and Reinforcement Learning from Human or AI Feedback (RLHF/RLAIF)Gao et al.
-(2022); Bai et al.
-(2022).
-IFT involves training an LM to generate responses based on provided prompts, which might optionally include task instructions.
-The data used for fine-tuning is usually less than the original pre-training corpus.
-These datasets used in IFT are often curated through crowd-sourcing or extracted from an LM that is already capable of generating instruction-following examplesHan et al.
-(2024).
-Studies have shown that IFT can enhance zero-shot generalization to various unseen tasks.
-To optimize the fine-tuning process, various strategies have been proposed to make it more parameter-efficient.
-These strategies include using adapters Hu et al.
-(2022), prompt tuning Li and Liang (2021), and other techniques.
-Continual Pre-training (CPT)
-Continual pretraining involves regularly updating LLMs with the latest facts or adapting them to specialized domains.
-To continually pre-train LLMs to update facts, researchers utilize dynamic datasets to assimilate real-time data from various sources such as news feeds Sun et al.
+To continually pre-train LLMs and update facts, researchers utilize dynamic datasets to assimilate real-time data from various sources such as news feeds Sun et al.
 (2019), scholarly articles Cossu et al.
@@
 Ke et al.
-(2023) introduces a soft-masking mechanism to update LLMs with domain corpora, aiming to boost performance while preserving general knowledge.
-Notably, there has been research on continuous pretraining for specific domains such as finance Xie et al.
+(2023) introduced a soft-masking mechanism to update LLMs with domain corpora, aiming to boost performance while preserving general knowledge.
+Notably, there has been research on continual pre-training for specific domains such as finance Xie et al.
 (2023) and e-commerce Ma et al.
@@
 (2023a).
-It can be broadly classified into two categories: 1) task-incremental CIT and 2) domain-incremental CIT.
+It can be broadly classified into two categories: (1) task-incremental CIT and (2) domain-incremental CIT Wu et al.
+(2024).
 Task-incremental CIT aims to continually fine-tune LLMs on a series of task-specific instructions to enable them to solve new tasks.
@@
 (2023).
-However, studies have shown this process would cause a catastrophic forgetting of the learned knowledge and problem-solving skills in previous tasks Kotha et al.
+However, studies have shown this process causes catastrophic forgetting of the learned knowledge and problem-solving skills from previous tasks Kotha et al.
 (2024).
-Domain-incremental CIT aims to continually fine-tuning of LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains.
+Domain-incremental CIT aims to continually fine-tune LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains.
 TAPT Gururangan et al.
@@
 3 Methods
-The objective of this work is to enhance the knowledge of LLM through continual learning, resulting in iterative improvements in model performance.
 3.1 Task Formulation
-Question-Answering (QA) tasks are commonly used to evaluate models’ world knowledge and reasoning skills.
-Our focus is on enhancing the models’ world knowledge rather than their reasoning abilities.
-These tasks usually involve a series of single-choice questions with multiple options.
-The ability of a LLM to correctly identify the correct answer within these questions indicates its possession of the relevant knowledge.
-Conversely, failure to do so suggests a deficiency in the model’s understanding of the knowledge pertinent to the question.
+Question Answering (QA) tasks are commonly used to evaluate LLM’s domain knowledge.
+These tasks typically involve a series of single-choice questions with multiple options.
+If an LLM correctly answers these questions, it indicates a strong grasp of the relevant knowledge.
+Conversely, incorrect answers reveal deficiencies in understanding.
 Given a question-answer example $(qi,ai)\in\mathcal{Q}$, a test set for question answering $\mathcal{Q}$, and a reasoning model $\mathcal{M}_{0}$.
@@
 where $\widetilde{r_{i}}$ denotes the response that includes incorrect answers to the question $q_{i}$, and $Ans(\cdot)$ is utilized to extract the final answer from the response.
-Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplemental corpus $C_{S_{i}}$,
-$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\dots\},$
+Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplemental corpus $C_{S_{i}}$ as CPT data,
+$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\ldots\},$
 where $Search(\cdot)$ is used to search for relevant knowledge based on the question and obtain supplemental corpus.
@@
 3.2 CEM Method
-We propose the CEM method, involving three main steps: (1) testing the LLM and collecting incorrectly answered questions, (2) collecting supplemental corpora from multiple data sources, and (3) constructing a supplemental training set and fine-tuning the LLM.
-Figure 1 shows the process.
+We propose the CEM method, as illustrated in Figure 2. The main steps are as follows:
 Figure 2: The pipeline of CEM method.
-Testing LLM and Collecting Incorrectly Answered Questions.
-We utilize the LLM to answer questions from a QA dataset.
-Subsequently, by utilizing pattern matching to extract answers from the LLM’s responses, we verify the accuracy of these answers and collect the questions that the LLM has answered incorrectly.
-Collecting Supplemental Corpora from Multiple Data Sources.
-We argue that the questions answered incorrectly by the LLM reveal knowledge deficiencies.
-Based on the implicit information contained within questions incorrectly answered by the LLM, we collect supplemental corpora from multiple data sources, including search engines and online encyclopedias.
-(1) Search Engine: Search engines are characterized by fast response time, high relevance matching, and diverse and rich content.
-We search for questions that the LLM answers incorrectly.
-The search engine ranks results based on relevance.
-We select the top-k results with the highest relevance and access their detailed pages to crawl the main textual content as supplemental corpora.
-To prevent the LLM from simply memorizing answers, we filter out search results that may contain both the original question and its answer.
-Additionally, we optimize the strategy for crawling the corpus to filter out irrelevant information and spam, such as advertisements.
-In subsequent experiments, we utilize Bing as the search engine and set $k=4$. Figure 4 in Appendix shows a case of corpus collection from Bing.
-(2) Online Encyclopedia: Compared to using search engines for retrieval, online encyclopedias provide more credible and professional search results, despite having lower hit rates than search engines.
-Since the organization of encyclopedias typically relies on keyword entries, we conduct entity extraction from the questions and options of the collected incorrectly answered questions, using the extracted entity keywords for retrieval.
-The main text of the retrieved results serve as supplemental corpora.
+Step 1&2: Testing LLM and Collecting Incorrectly Answered Questions.
+We use the LLM to answer questions from a QA dataset.
+By employing pattern matching to extract answers from the LLM’s responses, we calculate the accuracy and collect the questions, along with their options, that the LLM answered incorrectly.
+Step 3: Obtaining Supplemental Corpus from Multiple Data Sources.
+Based on the implicit information of knowledge deficiencies contained within questions incorrectly answered by the LLM, we collect supplemental corpus from multiple data sources, including search engines and online encyclopedias.
+(1) Search Engine: Search engines are characterized by high relevance matching and diverse content.
+We search for questions the LLM answers incorrectly.
+The search engine ranks results by relevance, and we select the top-k results to access their detailed pages, extracting the main textual content as supplemental corpus.
+To prevent the LLM from merely memorizing answers, we filter out results that contain both the original question and its answer.
+Additionally, we optimize the corpus crawling strategy to exclude irrelevant information and spam, such as advertisements.
+In subsequent experiments, we utilize Bing as the search engine and set $k=4$. Figure 6 in Appendix shows the process of corpus collection from Bing.
+(2) Online Encyclopedia: Compared to search engines, online encyclopedias provide more credible and professional results, despite lower hit rates.
+Since encyclopedias are typically organized by keyword entries, we conduct entity extraction from the questions and options, using the extracted entity keywords for retrieval.
+The main text of the retrieved results serves as the supplemental corpus.
 In subsequent experiment, we use Wikipedia as the online encyclopedia and instruct GPT-3.5 using a few-shot prompt for entity extraction.
-Figure 5 in Appendix shows a case of corpus collection from Wikipedia.
-Constructing Supplemental Training Set and Fine-Tuning LLM.
+Figure 7 in Appendix shows the process of corpus collection from Wikipedia.
+Step 4&5: Constructing Supplemental Training Set and Fine-Tuning LLM.
 The supplemental training sets are constructed separately based on the supplemental corpus collected from Wikipedia and Bing.
 They are then employed to fine-tune the LLM.
-The construction method of the training set will be explained in detail in Section 3.3.
+The construction strategies of the training set will be explained in detail in Section 3.3.
 After completing one round of supplemental training, the model is retested on the QA dataset to evaluate its performance and gather a new set of incorrectly answered questions.
-This can serve as the first step for the next round of supplemental training.
-It is important to note that supplemental training should not be never-ending, as this may lead to severe catastrophic forgetting and the boundary effect of diminishing accuracy improvement.
-In the following experiments, we will examine how the number of supplemental training rounds affects model improvement.
+This can serve as the first step for the next round of CEM.
+It should be noted that supplemental training should not be endless, as it may lead to catastrophic forgetting and diminishing returns in accuracy improvement.
+Subsequent experiments explore the impact of the number of training rounds on model performance.
 3.3 Supplemental Training Data
 Definition of Corpus Type. The construction of the supplemental training set involves the following five types of corpus.
-(1) Supplemental Corpus($C_{\underline{S}}$): Supplemental Corpus obtained from multiple data sources contains background knowledge that the LLM lacks and is used similarly to continuous pre-training Gupta et al.
+(1) Supplemental Corpus($C_{\underline{S}}$): The supplemental corpus, obtained from multiple data sources, contains knowledge that the LLM lacks and is used as CPT data.
+Below are the three types of CIT data:
+(2) Normative Corpus($C_{\underline{N}}$): To enable the LLM to answer multiple-choice questions in a normative format, we sample questions independent of the test set and process them into the format shown in Table 1 to build the Normative Corpus.
+Prior to supplemental training, we finetune the LLM using the Normative Corpus (i.e., pre-finetuning), which facilitates extracting answers from the LLM’s responses to calculate initial performance and allows for the collection of incorrectly answered questions.
+To maintain this ability, we incorporate a small proportion of the Normative Corpus during supplemental training.
+(3) Extractive Corpus($C_{\underline{E}}$): Inspired by the human learning process, where despite studying from the same textbooks, students who excel at extracting knowledge achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
+To enhance the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample questions independent of the test set and the Normative Corpus.
+These questions are processed into the format shown in Table 2 to build the Extractive Corpus.
+(4) Review Corpus($C_{\underline{R}}$):
+The responses of the LLM tend to be diverse under high temperature settings Caccia et al.
 (2018).
-Example 1: A Case of Normative Corpus
+During repeated QA testing, the LLM may change its answers to certain ’less confident’ questions due to the proximity of predicted probabilities.
+Consequently, after supplemental training, questions initially answered correctly may be answered incorrectly, thus affecting accuracy.
+To reinforce the LLM’s memory of the knowledge implied by correctly answered questions, we sample the responses provided correctly by the LLM in the previous round.
+We process some of these responses in the Normative Corpus format and the rest in the Extractive Corpus format to build the Review Corpus.
+The composition of the Review Corpus is described as follows:
+$C_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
+where $\alpha\in[0,1]$, denotes the proportion of responses processed into the Normative Corpus format.
+$r$ denotes the correct responses in the previous round.
+$Normative(\cdot)$ converts responses into the Normative Corpus format, and $Extractive(\cdot)$ converts responses into the Extractive Corpus format.
+(5) General Corpus($C_{\underline{G}}$): We replay previous IFT data to mitigate catastrophic forgetting when training LLMs on new data He et al.
+(2021).
+When previous IFT data is unclear, we use the IFT data from the ShareGPT dataset [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_unfiltered as the General Corpus to preserve the model’s general capabilities on other tasks.
+Supplemental Training Set Construction Strategy.
+We propose a novel CL dataset construction paradigm that parallelizes CIT and CPT data.
+Specifically, we construct three types of supplemental training sets by combining five types of corpora.
+The following sections detail these construction strategies.
+(1) Plain Strategy:
+$D_{\underline{P}}=Concat(C_{S},C_{N},C_{G}),$
+where $C_{S}$ denotes the Supplemental Corpus, $C_{N}$ denotes the Normative Corpus, $C_{G}$ denotes the General Corpus, $Concat{(\cdot)}$ merges these corpora, and $D_{P}$ represents the supplemental training set constructed using the Plain Strategy.
+(2) Extractive Strategy:
+$D_{\underline{E}}=Concat(C_{S},C_{N},C_{G},{\color[rgb]{1,0,0}C_{E}}),$
+where $C_{E}$ denotes the Extractive Corpus and $D_{E}$ represents the supplemental training set constructed using the Extractive Strategy.
+(3) Review Strategy:
+$D_{\underline{R}}=Concat(C_{S},C_{N},C_{G},C_{E},{\color[rgb]{1,0,0}C_{R}}),$
+where $C_{R}$ denotes the Review Corpus and $D_{R}$ represents the supplemental training set constructed using the Review Strategy.
+4 Experimental Setup
+4.1 Datasets
+We conduct experiments on two question-answering datasets, including Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024).
+To investigate the impact of catastrophic forgetting, we also use the HotpotQA Yang et al.
+(2018) and GSM8K Cobbe et al.
+(2021) datasets to test the model’s decline in reasoning and mathematical abilities.
+Appendix Table 6 shows the details of the data division and usage.
+4.2 Base Language Models
+To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs with parameter ranges from 6 to 13 billion.
+Our instruction fine-tuning base models include CuteGPT-13B [^2] https://huggingface.co/XuYipei/kw-cutegpt-13b-base Liang et al.
+(2023), ChatGLM-6B [^3] https://huggingface.co/THUDM/chatglm3-6b Du et al.
+(2022), and Qwen-7B-Chat [^4] https://huggingface.co/Qwen/Qwen-7B-Chat Bai et al.
+(2023).
+These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
+4.3 Metrics
+The effectiveness of supplemental training will be assessed using the following metrics in the main experiments:
+ACC: it represents the accuracy of the answer of LLM.
+W2R: it indicates the percentage of questions where the LLM changes incorrect answers to correct ones on the test set.
+R2W: it indicates the percentage of questions where the LLM changes correct answers to incorrect ones on the test set.
+To analyze the potential performance degradation caused by catastrophic forgetting, we also use the following metrics:
+ER: The Enhancement Rate measures the relative improvement brought by CEM on a specific task compared to the initial performance.
+It is calculated by $ER=\frac{R_{k}^{*}-R_{k}^{0}}{R_{i}^{0}},$ where $k$ denotes the task using CEM, $R^{0}_{k}$ is the initial accuracy, and $R^{*}_{k}$ is the accuracy after CEM.
+AFR: The Average Forgetting Rate Wang et al.
+(2024) measures the average relative degradation brought by CEM on tasks, excluding the task using CEM, compared to their initial performance.
+It is calculated by $AFR=\frac{1}{N-1}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\left(\frac{R_{i}^{0}-R_{i}^{*}}{R_{i}^{0}}\right)$, where $R^{0}_{i}$ is the initial accuracy of task $i$ after the initial fine-tuning, and $R_{i}^{*}$ is the accuracy of taks $i$ after applying CEM on task $k$.
+5 Experiments
+We construct supplemental training sets and control groups for the experiments, following the strategies described in Section 3.3.
+To distinguish between models trained on different supplemental training sets, we use the following naming conventions:
+INIT represents the model that has undergone only pre-finetuning using the Normative Corpus, without any supplemental training.
+CEM-P: P denotes Plain, representing the model trained with the supplemental training set constructed by the Plain Strategy.
+[…]
+[…]
+We append suffixes to indicate the data sources of the Supplemental Corpus: ‘Wiki’, ‘Bing’ and ‘Mix’.
+‘Mix’ indicates that the Supplemental Corpus is collected from both Bing and Wikipedia and concatenated.
+For example, ‘E-Wiki’ denotes the model trained using the Extractive Strategy with a Supplemental Corpus from Wikipedia.
+5.1 Main Results
+We used the Plain Strategy to construct supplementary training set, using approximately 4,000 CPT data (based on the initial number of mistakes) and 4,000 CIT data for single-round CEM.
+The main results are shown in Figure 3.
+Our observations are summarized as follows:
+Figure 3: This figure shows the accuracy of different models using the CEM-P method on the Xiezhi and CMMLU tasks.
+Firstly, following the supplemental training, a noticeable enhancement in accuracy is observed across all models for the two tasks.
+This indicates that the CEM method effectively addresses the model’s knowledge deficiencies, even with the small training set and readily available supplemental data.
+For instance, the ChatGLM-6B improves by up to 16.08% on the CMMLU task.
+Secondly, the effectiveness of CEM varies with the innate abilities and parameter scales of LLMs.
+For example, on the Xiezhi task, ChatGLM-6B, with the fewest parameters, improves by up to 13.17%.
+In contrast, CuteGPT-13B, with the highest initial accuracy and most parameters, shows a smaller improvement of up to 3.63%.
+Construction Strategy Analysis.
+To investigate the effectiveness of each component in the supplemental training set, we conducted experiments on CuteGPT-13B with various dataset construction strategies.
+We set the volume of both Extractive Corpus and Review Corpus to 1,000.
+The results are shown in Figure 4.
+Below is a summary of our observations:
+[…]
+Firstly, Extractive and Review Corpus significantly enhance LLM performance.
+CEM-E consistently outperforms its control group, showing higher accuracy and W2R metrics.
+Similarly, CEM-R outperforms its control group and significantly reduces the R2W metric.
+The results indicate an improvement in the W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the CPT data.
+This improvement leads to more effective error correction.
+Additionally, the Review Corpus significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
+Secondly, Supplemental Corpus from multiple sources do not always improve the CEM method more than a single source.
+Results in Figure 4 show no significant difference in CEM effectiveness whether using Wikipedia, Bing, or a combination of both.
+Indeed, R-Mix has lower accuracy than R-Wiki and R-Bing despite having more data.
+This may be due to redundancy from overlapping sources.
+Additionally, an excess of supplemental corpus can dilute the proportion of CIT data, reducing their effectiveness in mitigating catastrophic forgetting and enhancing knowledge extraction, and increasing the risk of overfitting.
+Thirdly, proportion factor $\alpha$ has limited impact on the effectiveness of Review Corpus.
+Table 5 in Appendix shows that R0, R0.5, and R1 have similar effects on LLM performance.
+This suggests that reviewing the knowledge in any format consolidates it.
+Therefore, unless specified, R will refer to the R0.5 in the following.
+5.2 Multiple Iterations Analysis
+We conduct a second round of CEM on CuteGPT-13B for the Xiezhi task, using the same training set construction strategy and data source as the first round.
+We select CIT data independent of the first round but maintain the same data volume.
+The experimental results presented in Appendix Table 8 show that after the second round of CEM, the accuracy of the LLM’s answers significantly improved across all dataset construction strategies.
+In the best case, R-Wiki shows a 4.75% improvement in accuracy compared to its previous round and a 17.00% improvement compared to INIT.
+Further analysis shows that the Review Strategy consistently has significantly lower R2W metrics compared to other dataset construction strategies.
+For example, the R2W metric of R-Wiki decreases by 7.02% compared to P-Wiki and by 5.90% compared to E-Wiki.
+We argue that the Review Strategy significantly reduces the probability of changing a previously correct response to an incorrect one, thereby helping the model avoid severe forgetting during multiple rounds of supplementary training.
+Figure 5: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after multiple rounds of CEM combined with the random replay method.
+Here, ER indicates the model’s enhancement rate on the Xiezhi task, while AFR represents the average forgetting rate on the HotpotQA and GSM8K tasks.
+Catastrophic Forgetting Analysis
+To analyze catastrophic forgetting that may arise from iterations, we use CuteGPT-13B and test it on Xiezhi, HotpotQA, and GSM8K tasks to evaluate the model’s domain knowledge, reasoning, and mathematical abilities.
+We employ the CEM method to enhance performance on Xiezhi and assess any performance declines on the other two tasks.
+Initially, we sample 2,000 instances from each of the three datasets for pre-finetuning, then test to establish baseline.
+We use a method of random replay to mitigate catastrophic forgetting.
+Specifically, in each training round, we randomly replay 500 instances from both HotpotQA and GSM8k to replace the original General Corpus.
+Figure 5 shows that even after three rounds of CEM, performance on Xiezhi still exhibits significant improvement and maintains an upward trend.
+Additionally, the combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 70% in most cases.
+However, we observe that CEM-R, with its larger data capacity, achieves significant performance improvements but also shows the greatest decline on other tasks.
+This may be attributed to the increased volume of CPT data diluting the effect of the replay data.
+Combining CEM with other advanced replay-based CL methods warrants further investigation.
+6 Discussion
+Why can the Extractive Corpus be effective?
+LLMs using the Extractive Corpus show higher accuracy and W2R metrics.
+Unlike the Review Corpus, where the role in performance improvement is intuitive, the precise mechanism by which the Extractive Corpus augments the model’s capability to capture pivotal knowledge in lengthy texts remains worthy of discussion.
+Extractive Corpus provides the model with both questions and supplemental material containing relevant knowledge.
+Questions and options typically present straightforwardly and can be condensed into several key points, while the supplemental material tends to intricately interweave numerous factual statements Jiang et al.
+(2024).
+Thus, we suggest that questions act as a structured induction of the supplemental material, where the Extractive Corpus explicitly links complex information to the resulting induction (i.e., the question and options).
+Questions and options can serve as cues and guides for the model to gather relevant information from the corpus, helping it consider how this knowledge is accessed through questions.
+This process enhances the model’s ability to extract useful information from the corpus while filtering out redundant and irrelevant details.
+7 Conclusion
+In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes.
+We believe that the difficulty of collecting CPT data and the limitations of instruction tuning’s superficial alignment have become a bottleneck for further performance improvement.
+To address this, we design a pipeline to collect corpus from mistakes and a novel parallel dataset construction paradigm using both CIT and CPT data to achieve continual and efficient internal knowledge supplementation during fine-tuning stage.
+We conduct experiments on several open-source LLMs to validate the efficacy of CEM.
+The results consistently showing significant improvements in answer accuracy and potential for continual evolution through multiple iterations.
+In the best case, accuracy enhancement reaches up to 17.00%.
+Appendix A Format of CIT Data
+Table 1 and Table 2 show the IFT data formats of the Normative Corpus and the Extractive Corpus.
+An Example of Normative Corpus
 Question:
@@
 (answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
-Table 1: A case of a Normative Corpus, designed to illustrate its corpus format.
-Example 2: A Case of Extractive Corpus
+Table 1: An example illustrating the format of the Normative Corpus.
+An Example of Extractive Corpus
 Question:
@@
 (answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
-Table 2: A case of a Extractive Corpus, designed to illustrate its corpus format. The ellipsis indicates that some supplemental text has been omitted for clarity.
-(2) Normative Corpus($C_{\underline{N}}$): To equip the LLM with the ability to respond to multiple-choice questions in a normative format, we sample multiple-choice questions that are independent of the test set and process them into the format shown in Table 1 to build Normative Corpus.
-Prior to supplemental training, we finetune the LLM using Normative Corpus(i.e.
-pre-finetuning), facilitating the extraction of answers from the LLM’s responses and allowing for the collection of incorrectly answered questions.
-To maintain this ability, we incorporate a small proportion of Normative Corpus during the supplemental training to prevent the LLM’s response skills from declining.
-(3) Extractive Corpus($C_{\underline{E}}$): Inspired by the human learning process: despite studying from the same textbooks, students who are better at extracting knowledge from these materials always achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
-To improve the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample multiple-choice questions that are independent of the test set and the Normative Corpus.
-These questions are processed into the format shown in Table 2 to build the Extractive Corpus.
-(4) Review Corpus($C_{\underline{R}}$): The responses of the LLM tend to be diverse under high temperature Caccia et al.
-(2018).
-During repeated QA testing, LLM may change its answers to certain ‘less confident’ questions due to the proximity of predicted probabilities.
-Consequently, after supplemental training, questions initially answered correctly may be answered incorrectly, thus affecting the accuracy.
-Therefore, to strengthen the LLM’s memory of the knowledge implied by correctly answered questions, we sample the responses provided correctly of the LLM in the previous round and process some of them in Normative Corpus format and the rest in Extractive Corpus format to build the Review Corpus.
-The composition of the Review Corpus can be described as:
-$C_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
-where $\alpha\in[0,1]$, denotes the proportion of responses processed into the Normative Corpus format.
-$r$ denotes the correct responses in the previous round of LLM.
-$Normative(\cdot)$ converts responses into the Normative Corpus format, and $Extractive(\cdot)$ converts responses into the Extractive Corpus format.
-(5) General Corpus($C_{\underline{G}}$): Replaying previous training data can effectively mitigate catastrophic forgetting when training LLM on new dataHe et al.
-(2021).
-Therefore, to preserve the model’s general capabilities on other tasks, we use the data from ShareGPT dataset  [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_
-unfiltered
-as General Corpus.
-Supplemental Training Set Construction Strategy.
-We have developed three innovative strategies for constructing supplemental training sets by combining five types of corpus.
-The strategies are detailed below.
-(1) Plain Strategy: To mitigate the potential catastrophic forgetting and to maintain the LLM’s ability to provide normative responses, we mix three types of corpus to construct the supplemental training set:
-$D_{P}=Concat(C_{S},C_{N},C_{G}),$
-where $C_{S}$ denotes the Supplemental Corpus, $C_{N}$ denotes the Normative Corpus, $C_{G}$ denotes the General Corpus, $Concat{(\cdot)}$ merges various types of corpus together and $D_{P}$ denotes the supplemental training set constructed via the Plain Strategy.
-(2) Extractive Strategy: To enhance the LLM’s ability to capture and comprehend knowledge from supplemental material, we integrate the Extractive Corpus:
-$D_{E}=Concat(C_{S},C_{N},C_{G},C_{E}),$
-where $C_{E}$ denotes the Extractive Corpus and $D_{E}$ denotes the supplemental training set constructed via the Extractive Strategy.
-(3) Review Strategy: To enhance the model’s retention of acquired knowledge and prevent correct responses from being altered to incorrect ones over multiple iterations, we integrate the Review Corpus:
-$D_{R}=Concat(C_{S},C_{N},C_{G},C_{E},C_{R}),$
-where $C_{R}$ denotes the Review Corpus and $D_{R}$ denotes the supplemental training set constructed via the Review Strategy.
-4 Experimental Setup
-4.1 Datasets
-We conduct experiments on two question-answering datasets, including Xiezhi [^2] https://github.com/mikegu721/xiezhibenchmark and CMMLU [^3] https://github.com/haonan-li/CMMLU.
-Xiezhi, a comprehensive evaluation suite designed to assess holistic domain knowledge.
-It comprises multiple-choice questions across 516 diverse disciplines, ranging from 13 different subjects.
-Gu et al.
-(2023).
-CMMLU contains about 12,000 questions spanning elementary to advanced professional levels across 67 subjectsLi et al. (2024).
-In our experiments, we select 27,265 multiple-choice questions from Xiezhi and 11,917 multiple-choice questions from CMMLU. Appendix Table 7 shows the details of the data division.
-4.2 Base Language Models
-To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs with parameter ranges from 6 to 13 billion.
-Our instruction fine-tuning base models include CuteGPT-13B [^4] https://huggingface.co/XuYipei/kw-cutegpt-13b-base, ChatGLM-6B [^5] https://huggingface.co/THUDM/chatglm3-6b, and Qwen-7B-Chat [^6] https://huggingface.co/Qwen/Qwen-7B-Chat.
-These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
-CuteGPT-13B is based on the original Llama model structure, expands the Chinese vocabulary and performs pre-trainingLiang et al. (2023).
-ChatGLM-6B is an open bilingual language model. It is trained for about 1T tokens of Chinese and English corpusDu et al. (2022).
-Qwen-7B-Chat is a Transformer-based large language model, which is pretrained on a large volume of data, including web texts, books, codes, etcBai et al. (2023).
-4.3 Metrics
-The effectiveness of supplemental training will be assessed using the following metrics:
-ACC: it represents the accuracy of the answer of LLM.
-W2R: it indicates the percentage of questions where the LLM changes incorrect answers to correct ones on the test set.
-R2W: it indicates the percentage of questions where the LLM changes correct answers to incorrect ones on the test set.
-5 Experiments
-We construct various supplemental training sets and their corresponding control groups for the experiments, following the method described in Section 3.3.
-To differentiate between models trained on different supplemental training sets, we establish the following naming conventions:
-INIT: it represents the model, which has undergone only pre-finetuning using the Normative Corpus, without any supplemental training.
-CEM-N: N denotes Normative, representing the model that has undergone supplemental training using the training set constructed by the Plain Strategy.
-[…]
-[…]
-We append suffixes to the aforementioned naming conventions to differentiate the data sources of the supplemental corpus.
-These suffixes include: ‘Wiki’, ‘Bing’ and ‘Mix’ where ‘Mix’ indicates that the Supplemental Corpus is collected from both Bing and Wikipedia and concatenated.
-For instance, ‘E_Wiki’ denotes the model supplemented trained using the dataset constructed through the Extractive Strategy.
-The Supplemental Corpus of this dataset is collected from Wikipedia.
-5.1 Main Results
-We conduct experiments to analyze the accuracy after applying CEM on the test data. The results are shown in Figure 2. Our observations are summarized as follows:
-Figure 2: This figure shows the accuracy of various models using the CEM method on the Xiezhi and CMMLU tasks. The supplemental training set is constructed using the Plain Strategy.
-Firstly, following the supplemental training, a noticeable enhancement in accuracy is observed across all models for the two given tasks.
-This implies that despite the high availability of supplemental data and the small size of the training set, the CEM method effectively complements the model’s knowledge deficiencies.
-For instance, the ChatGLM-6B model shows an improvement of up to 16.08% on the CMMLU task.
-Secondly, the effectiveness of tuning LLMs with CEM depends on their innate abilities and also varies across LLMs with different parameter scales.
-Among the mentioned open-source models, ChatGLM-6B, which has the fewest parameters, exhibits the most significant improvement, up to 13.17% on the Xiezhi task.
-In contrast, CuteGPT-13B, with the highest initial answer accuracy and the most parameters, shows a smaller improvement of up to 3.63% compared to the smaller LLM.
-Construction Strategy Analysis.
-To investigate the effectiveness of each component in supplemental training set, we conduct experiments on LLMs trained using various dataset construction strategies, the results are shown in Figure 3.
-Our observations are summarized as follows:
-Figure 3: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplemental training set constructed using various strategies.
-’Avg of $R_{\alpha}$’ and ’Best of $R_{\alpha}$’ represent the average and best results, respectively, of the Review Strategy with different $\alpha$.
-Firstly, Extractive Corpus and Review Corpus can significantly boost LLMs’ performance.
-CEM with Extractive Corpus ($C_{E}$) outperforms its control group in all cases with higher Accuracy and W2R metric, e.g., E_Mix shows a 3.11% increase in accuracy to its control group E’_Mix, and a 10.96% increase compared to INIT.
-Meanwhile, in alignment with expectations, CEM with the Review Corpus ($C_{R_{\alpha=0,0.5,1}}$), consistently outperform their control group across all instances and significantly reduce the R2W metric.
-For example, R0.5_Wiki shows a 5.74% improvement in accuracy and a 3.70% reduction in R2W compared to its control group, R’_Wiki, and a 12.25% improvement in accuracy and a 5.89% reduction in R2W compared to the INIT.
-The results indicate an improvement in W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the corpus.
-This enhancement leads to more effective error correction.
-Additionally, the Review Corpus significantly reduced R2W, indicating suppression of the model’s tendency to forget correct knowledge.
-Secondly, Supplemental Corpus from multiple sources do not always improve the CEM method more than a single source.
-The effectiveness of CEM depends on the recall rate when searching corpus and the quality of supplemental corpus, as this determines whether the corpus contains knowledge that the LLM lacks.
-Moreover, noise and errors within low-quality corpora can negatively impact the LLM.
-Results in Figure 3 show that there is no significant difference in the effectiveness of CEM, whether using Wikipedia or Bing as a single data source, or when using multiple data source concatenated.
-When using the supplemental training set constructed by Review Strategy, R_Mix has lower accuracy than R_Wiki and R_Bing despite having more data volumn, This may be due to the information redundancy caused by the high overlap of the supplemental corpus from different sources.
-Additionally, an excess of supplemental corpus may dilute the proportion of General, Extractive, and Normative Corpus, thereby hindering their effectiveness in reducing model catastrophic forgetting and enhancing model knowledge extraction.
-Thirdly, proportion factor $\alpha$ has limited impact on the effectiveness of the Review Corpus.
-According to Table 6 in Appendix, R0, R0.5 and R1 are similar in their ability to enhance LLMs’ performance.
-This may be because the models themselves have already mastered the knowledge, and replaying it in either format serves to consolidate that knowledge.
-Therefore, if not specified, R will refer to the R0.5 Strategy in the following.
-5.2 Multiple Iterations Analysis
-Metrics Wiki Bing
-CEM-N CEM-E CEM-R CEM-N CEM-E CEM-R
-Acc INIT 42.93 42.93 42.93 42.93 42.93 42.93
-Round1 46.56 52.02 55.18 45.91 48.46 53.86
-Round2 51.69 55.20 59.93 46.37 48.74 58.33
-W2R vs. INIT 24.46 26.03 28.29 22.07 23.32 27.01
-vs. Round1 20.00 17.56 16.43 18.49 18.33 16.61
-R2W vs. INIT 15.70 13.76 11.29 18.63 17.51 11.61
-vs. Round1 14.87 14.39 11.68 18.03 18.06 12.14
-Table 3: We employ the identical training data construction strategy as in the first round to conduct a second round of supplementary training on the LLM.
-The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task.
-We conduct a second round of supplemental training on CuteGPT-13B on Xiezhi task using the same dataset construction strategy and data source as the first round.
-The results are presented in Table 3.
-After the second round of supplementary training, the accuracy of LLM’s answer on both two QA tasks has significantly improved compared to the previous round.
-Both the Extractive Strategy and Review Strategy are effective in improving the accuracy of the model.
-For instance, R_Wiki shows a 4.75% improvement in accuracy compared to its previous round, and a 17.00% improvement in accuracy compared to INIT.
-Furthermore, the Review Strategy consistently exhibits significantly lower R2W metric compared to other supplementary dataset construction strategies, e.g., the R2W metric relative to INIT of Wiki_R decreases by 7.02% compared to Wiki_N and 5.90% compared to Wiki_E.
-We argue that the Review Strategy performs best because it significantly reduces the probability of changing a previously correct response to an incorrect one.
-This enables the model to be evolved by two or more rounds of supplemental training.
-6 Discussion
-Why can the Extractive Corpus be effective?
-We observe that models employing the Extractive Corpus exhibit higher accuracy and W2R metric.
-Unlike the Review Corpus, where improving model performance is intuitive, the precise mechanism by which the Extractive Corpus augments the model’s capability to capture pivotal knowledge in lengthy texts remains worthy of discussion.
-Extractive Corpus provides the model with both questions and supplemental material containing relevant knowledge.
-Questions and options typically present straightforwardly and can be condensed into several key points, whereas the supplemental material tends to intricately interweave numerous factual statements Jiang et al.
-(2024).
-Thus, we suggest that questions may be a structured induction of supplemental material, where the Extractive Corpus explicitly links complex information within the corpus to the resulting induction (i.e., the question and options).
-Questions and options can act as cues and guides for the model to gather relevant information from the corpus, considering how this knowledge is accessed through questions.
-This, in turn, can enhance the model’s ability to extract useful information from the corpus while eliminating redundant and irrelevant information.
-7 Conclusion
-In this paper, we propose and validate the Continue Evolving from Mistakes (CEM) method, which assesses LLM’s knowledge deficiencies through LLM’s mistakes and targets supplementary LLM’s knowledge.
-We first devise a general process for supplementary training: (1) testing LLM and collecting mistakes, (2) collecting supplemental corpora from multiple data sources, (3) constructing a supplementl training set and fine-tuning the LLM.
-Furthermore, we introduce multiple strategies for constructing training data to enhance the LLM’s capability to extract and comprehend knowledge within the corpora, and to mitigate the forgetting of LLM.
-We conduct experiments on several open-source LLMs to validate the efficacy of CEM.
-The results consistently show that CEM supplements the LLM’s knowledge deficiencies and improves the accuracy of the answers.
-For instance, on the Xiezhi task, the accuracy of the CuteGPT-13B improves by up to 12.25% in the first round and up to 17.00% in the second round.
-8 Limitations
-This paper focuses on the model’s performance on question-answering task. It is yet to be determined if CEM can consistently enhance the model’s other abilities and its performance in non-QA tasks.
-In future research, we will further explore three or more rounds of supplemental training to improve the LLM and to mitigate the catastrophic forgetting.
-In addition, we will experiment with other data sources for corpora and optimize supplementary corpus collection strategies to equip the LLM with more advanced continual iterative evolutionary capabilities.
-Appendix A Process of Collecting Corpora from Multiple Sources
+Table 2: An example illustrating the format of the Extractive Corpus. The ellipsis indicates that some supplemental text has been omitted for clarity.
+Appendix B Process of Collecting Corpora from Multiple Sources
 Fig.6 and Fig.7 show the process of obtaining supplemental data from Wikipedia and Bing respectively.
-Figure 4: The process of collecting corpora from Bing.
-Figure 5: The process of collecting corpora from Wikipedia.
-Appendix B Experimental Results of Various LLMs on QA tasks
+Figure 6: The process of collecting supplementary corpus from Bing.
+Figure 7: The process of collecting supplementary corpus from Wikipedia.
+Appendix C Experimental Results of LLMs on QA tasks
 Table 3 and Table 4 show the complete experimental results of LLMs using the CEM method on Xiezhi and CMMLU tasks. We use the Plain Strategy to construct the supplemental training dataset.
@@
 Table 4: The table shows the experimental results of LLMs using the CEM method on the CMMLU task.
-Appendix C Experimental Results with Various supplemental Training Dataset Construction Strategies
-Table 6 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplemental training set constructed using various strategies.
+Appendix D Experimental Results with Various supplemental Training Dataset Construction Strategies
+Table 5 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with supplemental training sets constructed using the Plain, Extractive, and Review strategies, as well as various data sources.
 Data Metrics INIT CEM-N CEM-E CEM-E’ CEM-R0.5 CEM-R1 CEM-R0 CEM-R’
@@
 Table 5: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplemental training set construction methods.
-Appendix D Data Division
-Table 7 shows the the data division of Xiezhi and CMMLU.
+Appendix E Data Division
+Table 6 and Table 7 shows the the data division of Xiezhi and CMMLU.
 The multiple-choice questions are divided into three parts.
@@
 CMMLU 11,917 4,000 5,000 2,917
-Table 7: The data division of Xiezhi and CMMLU.
+Table 6: The data division and usage of Xiezhi and CMMLU.
+Dataset Total Pre-Finetuning Testing Random Replay
+HotpotQA 7,000 2,000 5,000 500
+GSM8K 7,000 2,000 5,000 500
+Table 7: The data division and usage of HotpotQA and GSM8K, where the data used for Random Replay is included in pre-finetuning.
+Appendix F Experimental Results of Multiple Rounds of CEM
+We employ the identical training data construction strategy as in the first round to conduct a second round of CEM on the LLM.
+Table 8 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after two rounds of CEM.
+Metrics Wiki Bing
+CEM-N CEM-E CEM-R CEM-N CEM-E CEM-R
+Acc INIT 42.93 42.93 42.93 42.93 42.93 42.93
+Round1 46.56 52.02 55.18 45.91 48.46 53.86
+Round2 51.69 55.20 59.93 46.37 48.74 58.33
+W2R vs. INIT 24.46 26.03 28.29 22.07 23.32 27.01
+vs. Round1 20.00 17.56 16.43 18.49 18.33 16.61
+R2W vs. INIT 15.70 13.76 11.29 18.63 17.51 11.61
+vs. Round1 14.87 14.39 11.68 18.03 18.06 12.14
+Table 8: The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task, where "v.s.
+INIT" indicates W2R and R2W calculated based on the initial evaluation results, and "v.s.
+Round1" indicates W2R and R2W calculated based on the evaluation results after one round of CEM.
+Appendix G Experimental Results of Catastrophic Forgetting Analysis
+Table 9 shows the complete experimental results of CuteGPT-13B on the Xiezhi, GSM8K, and HotpotQA tasks after one, two, and three rounds of CEM, respectively.
+Round Strategy Task Accuracy Relative Rate
+Xiezhi GSM8K HotpotQA ER AFR
+INIT 40.84 13.20 80.60 1.00 1.00
+Round 1 CEM-P 42.02 6.64 51.60 1.03 0.730
+CEM-E 42.76 8.36 55.52 1.05 0.681
+CEM-R 51.49 7.54 55.62 1.26 0.673
+Round 2 CEM-P 48.68 9.94 57.84 1.19 0.723
+CEM-E 49.47 9.76 60.24 1.21 0.746
+CEM-R 54.23 6.88 56.68 1.33 0.703
+Round 3 CEM-P 50.58 10.02 60.96 1.24 0.757
+CEM-E 52.69 9.46 62.98 1.29 0.772
+CEM-R 58.91 7.48 59.86 1.44 0.678
+Table 9: The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi, GSM8K, and HotpotQA tasks.

### v5

@@
 Abstract
-As world knowledge evolves and new task paradigms emerge, Continual Learning (CL) is crucial for keeping Large Language Models (LLMs) up-to-date and addressing their shortcomings.
-In practical applications, LLMs often require both continual instruction tuning (CIT) and continual pre-training (CPT) to adapt to new task paradigms and acquire necessary knowledge for task-solving.
-However, it remains challenging to collect CPT data that addresses the knowledge deficiencies in models while maintaining adequate volume, and improving the efficiency of utilizing this data also presents significant difficulties.
-Inspired by the ’summarizing mistakes’ learning skill, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide a data-efficient approach for collecting CPT data and continually improving LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge.
-To efficiently utilize these CPT data and mitigate forgetting, we design a novel CL training set construction paradigm that integrates parallel CIT and CPT data.
-Extensive experiments demonstrate the efficacy of the CEM method, achieving up to a 17% improvement in accuracy in the best case.
-Furthermore, additional experiments confirm the potential of combining CEM with catastrophic forgetting mitigation methods, enabling iterative and continual model evolution.
+As world knowledge evolves and new task schemas emerge, Continual Learning (CL) is crucial for keeping Large Language Models (LLMs) up-to-date and addressing their shortcomings.
+LLMs typically require continual instruction tuning (CIT) and continual pre-training (CPT) to adapt to new tasks and acquire essential knowledge.
+However, collecting sufficient CPT data while addressing knowledge gaps remains challenging, as does optimizing the efficiency of utilizing this data.
+Inspired by the ‘summarizing mistakes’ strategy, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient approach aiming to collect CPT data and continually improve LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge.
+To enhance data utilization and mitigate forgetting, we introduce a novel training paradigm that combines CIT and CPT data.
+Experiments demonstrate that CEM significantly enhances model performance and continual evolution.
+The code and dataset are available in the GitHub  [^1] https://anonymous.4open.science/r/cem-BB25.
 Large Language Model Can Continue Evolving From Mistakes
@@
 1 Introduction
-With the exponential growth of data and parameter sizes, Large Language Models (LLMs) have demonstrated superior performance across numerous downstream tasks Chang et al.
+With the exponential growth of data and model sizes, Large Language Models (LLMs) have demonstrated superior performance across numerous downstream tasks Chang et al.
 (2023); Wei et al.
@@
 (2023).
-However, in real-world applications, the continual emergence of new knowledge and task requirements necessitates ongoing task adaptation and knowledge supplementation for LLMs.
-Otherwise, LLMs may fall into hallucinations and generate incorrect responses Huang et al.
+However, in real-world applications, the continual emergence of new knowledge and evolving task requirements necessitate ongoing updates and task-specific adaptations for LLMs.
+Without these, models risk inducing hallucinations due to knowledge boundaries and may cause task misalignment Huang et al.
 (2023).
-Therefore, Continual Learning (CL) is proposed to ensure LLMs’ alignment with the evolving world knowledge and demands.
-Figure 1: The figure illustrates two potential triggers for poor model performance: (1) Task Paradigm Unfamiliarity, and (2) Lack of task-relevant knowledge.
-Unfamiliarity with the task paradigm can cause deviations from expected interaction styles, while insufficient task knowledge may lead to hallucinations.
-Instruction tuning addresses only the former.
-Wu et al.
-(2024) categorize Continual Learning into Continual Pre-training (CPT) and Continual Instruction Tuning (CIT).
-CPT requires a larger volume of training data to enrich LLMs’ knowledge for new domains, while CIT continually fine-tunes LLMs on emerging task-specific instruction data to adapt the models to evolving task paradigms with much smaller training costs and data requirements Razdaibiedina et al.
-(2023); Gururangan et al.
-(2020b); Cossu et al.
+Additionally, addressing and correcting the shortcomings and errors exposed during practical use is crucial.
+Therefore, Continual Instruction Tuning (CIT) and Continual Pre-training (CPT) are proposed as the primary methods of Continual Learning (CL) to align LLMs with evolving knowledge and tasks, and to improve their shortcomings Wu et al.
 (2024).
-However, recent work has highlighted the limitations of instruction tuning, suggesting it is merely a superficial alignment process focused on learning interaction styles and LLMs primarily acquire knowledge during the pre-training stage Zhou et al.
+Figure 1: Two potential triggers for poor model performance: (1) Task Schema Unfamiliarity, and (2) Lack of Task-relevant Knowledge.
+Unfamiliarity with the task schema can cause deviations from expected interaction styles, while insufficient task knowledge may lead to hallucinations.
+Instruction tuning has been shown to be effective for addressing the former, but poor for the latter Gekhman et al.
+(2024a); Zhou et al.
 (2023).
-This will trigger the bad response shown in Figure 1, where LLMs’ poor performance may result from a lack of requisite knowledge to solve the task rather than unfamiliarity with the task paradigm Ren et al.
+Despite CIT’s smaller data size requirement and lower risk of overfitting compared to CPT, it is limited in new knowledge injection, as Zhou et al.
+(2023) highlight that instruction tuning is a superficial alignment focused on interaction styles (i.e., task schema) and does not effectively help the model acquire new knowledge.
+Thus, it cannot address all the issues shown in Figure 1, where LLMs’ poor performance may stem from a lack of requisite knowledge rather than unfamiliarity with the task schema Ren et al.
 (2023); Zhang et al.
 (2023b).
-CPT is more effective than CIT in addressing this issue.
-However, using CPT to supplement the model with the lacking knowledge required for task-solving presents several issues: First, collecting CPT data is challenging, as it is difficult to a priori identify the lacking knowledge while ensuring a sufficiently large data scale.
-Blind data collection and insufficient data volume may introduce irrelevant information, weaken generalization, and pose overfitting risks Zhang et al.
+In contrast, CPT is more effective in addressing the lack of knowledge.
+However, using CPT to supplement the model with lacking knowledge still presents several challenges.
+First, collecting CPT data is challenging as identifying the lacking knowledge a priori while ensuring a sufficiently large data size is difficult.
+Blind data collection and insufficient data volume may introduce irrelevant or redundant information, weaken generalization, and increase overfitting risks Zhang et al.
 (2024); Gekhman et al.
+(2024b).
+Second, efficiently utilizing CPT data and mitigating degradation are difficult due to weak constraints on parameter updates, leading to larger parameter shifts and a data size gap compared to the initial pre-training.
+To address these issues, the most common data collection methods involve using the latest Wikipedia snapshots or new academic papers in specialized fields Cossu et al.
+(2022); Jin et al.
 (2022).
-Second, due to the weak constraints of self-supervised CPT data during parameters updates and the scale gap compared to initial pre-training data, efficiently utilizing CPT data and avoiding catastrophic forgetting remain difficult.
-To address these issues, recent work has explored converting CPT data into CIT data, such as transforming it into reading comprehension questions to avoid the problem of insufficient data volume.
-Cheng et al.
-(2024) or using highly relevant QA pairs for CIT before CPT to enhance the data utilization efficiency Jiang et al.
-(2024).
-However, these methods demand high-quality data and the assistance of oracle LLM, and they overlook the crucial challenge of CPT mentioned above.
-Inspired by the ‘summarizing mistakes’ learning skill of human students, we argue that a model’s mistakes in ’exams’ reflect its inherent knowledge deficiencies.
-CPT data that directly addresses these gaps is the most effective for supplementing the model’s knowledge.
-Currently, the successful utilization of mistakes in constructing CoT data An et al.
-(2024); Tong et al.
-(2024) has enhanced the model’s performance in mathematics and reasoning, indicating the potential to leverage mistakes in continual learning.
-In this paper, we propose the Continue Evolving from Mistakes(CEM) method, aiming to provide a data-efficient method for collecting CPT data and continually enhancing the performance of LLMs through multiple evaluation and supplementation with mistake-relevant knowledge.
-Specifically, at the start of each CEM round, we test the model, identify cases of poor performance, and deconstruct them to pinpoint knowledge deficiencies.
-We then gather abundant supplemental data based on mistake analysis from search engines and online encyclopedias as CPT data to address these gaps.
-To efficiently utilize these CPT data and mitigate forgetting, we designed a novel CL training set construction paradigm that integrates parallel CIT and CPT data.
-By integrating CEM with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al.
+However, these methods lack focus and efficiency and fail to dynamically update model deficiencies.
+Additionally, recent work has explored using highly relevant QA pairs for CIT before CPT Jiang et al.
+(2024) or converting raw CPT corpus into reading comprehension instructions (i.e., CIT data) Cheng et al.
+(2024) to mitigate degradation and enhance the efficiency of knowledge absorption from CPT data.
+Nevertheless, reliance on oracle LLMs and the complexity of mining patterns limit their practicality.
+In this paper, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide (1) a practical pipeline for collecting CPT data based on model mistakes and (2) a novel training set construction paradigm to efficiently utilize these data, enabling continual model improvement while mitigating degradation.
+To enhance the targetedness and efficiency of data collection, we draw inspiration from the human learning skill of ‘summarizing mistakes’.
+We argue that a model’s mistakes in ‘exams’ reflect its inherent knowledge deficiencies.
+Therefore, we collect corpora from the internet that contain background knowledge related to these mistakes for directly address these deficiencies.
+To efficiently utilize these CPT data and mitigate forgetting, we design a novel training set construction paradigm that integrates parallel CIT and CPT data.
+In addition, by integrating CEM with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al.
 (2022), we enable iterative and continual model evolution.
-Our contributions are summarized as follows: (1) We propose the CEM method, which identifies knowledge deficiencies from LLMs’ mistakes and collects abundant relevant CPT data, enabling continual iterative evolution.
-(2) To assist LLMs in acquiring and comprehending factual knowledge from CPT data and mitigating knowledge forgetting, we propose three strategies for constructing supplemental training set using a parallel paradigm with CIT and CPT data.
-(3) We conduct experiments on multiple open-source LLMs for two complex knowledge QA tasks.
-The CEM method consistently demonstrates significant improvements across all scenarios, , achieving up to a 17.00% accuracy enhancement.
-(4) We verify that CEM can be combined with methods to mitigate catastrophic forgetting, such as replay-based methods Chaudhry et al.
-(2019); Wu et al.
-(2021), ensuring suppression of catastrophic forgetting over multiple iterations.
+The CEM method consistently demonstrates significant improvements across multiple models in extensive experiments, achieving up to a 17.00% accuracy enhancement.
 2 Related Work
@@
 (2024).
-Task-incremental CIT aims to continually fine-tune LLMs on a series of task-specific instructions to enable them to solve new tasks.
-A straightforward solution is to continuously generate instruction-tuning data for new tasks and directly fine-tune LLMs on it Wang et al.
-(2023).
-However, studies have shown this process causes catastrophic forgetting of the learned knowledge and problem-solving skills from previous tasks Kotha et al.
+Task-incremental Continual Instruction Tuning (CIT) aims to continually fine-tune LLMs on new task-specific instructions.
+While a straightforward approach involves generating new instruction-tuning data and fine-tuning LLMs directly Wang et al.
+(2023), this method can lead to catastrophic forgetting of previously learned knowledge and skills Kotha et al.
 (2024).
@@
 TAPT Gururangan et al.
-(2020a) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain.
+(2020) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain.
 AdaptLLM Cheng et al.
@@
 Step 3: Obtaining Supplemental Corpus from Multiple Data Sources.
-Based on the implicit information of knowledge deficiencies contained within questions incorrectly answered by the LLM, we collect supplemental corpus from multiple data sources, including search engines and online encyclopedias.
-(1) Search Engine: Search engines are characterized by high relevance matching and diverse content.
-We search for questions the LLM answers incorrectly.
-The search engine ranks results by relevance, and we select the top-k results to access their detailed pages, extracting the main textual content as supplemental corpus.
-To prevent the LLM from merely memorizing answers, we filter out results that contain both the original question and its answer.
-Additionally, we optimize the corpus crawling strategy to exclude irrelevant information and spam, such as advertisements.
-In subsequent experiments, we utilize Bing as the search engine and set $k=4$. Figure 6 in Appendix shows the process of corpus collection from Bing.
-(2) Online Encyclopedia: Compared to search engines, online encyclopedias provide more credible and professional results, despite lower hit rates.
-Since encyclopedias are typically organized by keyword entries, we conduct entity extraction from the questions and options, using the extracted entity keywords for retrieval.
-The main text of the retrieved results serves as the supplemental corpus.
-In subsequent experiment, we use Wikipedia as the online encyclopedia and instruct GPT-3.5 using a few-shot prompt for entity extraction.
-Figure 7 in Appendix shows the process of corpus collection from Wikipedia.
+We first employ the lightweight LLM, Qwen2-7B-instruct [^2] https://huggingface.co/Qwen/Qwen2-7B-Instruct Yang et al.
+(2024), to identify knowledge points involved in questions erroneously answered by the tested LLM.
+For further details, see Appendix A.
+The extracted knowledge points are then utilized as keywords to aggregate supplemental corpora from diverse sources, including search engines and online encyclopedias.
+(1) Search Engine: Search engines are known for high relevance matching and diverse content.
+We search for extracted knowledge points using Bing and select the top-k results based on relevance, scraping the main textual content as a supplemental corpus.
+To prevent LLMs from simply memorizing answers, we filter out results that contain both the original question and its answer.
+Moreover, we use newspaper3k [^3] https://github.com/codelucas/newspaper and readability [^4] https://github.com/buriy/python-readability to exclude irrelevant and spam content.
+(2) Online Encyclopedia: Compared to search engines, online encyclopedias offer more credible and professional results, despite having lower hit rates.
+Since encyclopedias are typically organized by keyword entries, we employ the state-of-the-art massive text embedding model, Conan-embedding-v1555https://huggingface.co/TencentBAC/Conan-embedding-v1 Li et al.
+(2024b), to generate vector embeddings for Wikipedia entry titles.
+FAISS [^6] https://github.com/facebookresearch/faiss Douze et al.
+(2024) is then used to build an efficient vector index, enabling searches based on inner product similarity.
+By calculating the similarity between input keyword embeddings and the FAISS-indexed entry embeddings, we retrieve the top-k results that exceed a predefined similarity threshold, serving as the supplemental corpus.
+Appendix C presents the experimental parameters and processes.
 Step 4&5: Constructing Supplemental Training Set and Fine-Tuning LLM.
@@
 (1) Supplemental Corpus($C_{\underline{S}}$): The supplemental corpus, obtained from multiple data sources, contains knowledge that the LLM lacks and is used as CPT data.
+We employ Qwen2-72B-Instruct [^7] https://huggingface.co/Qwen/Qwen2-72B-Instruct Yang et al.
+(2024) for entailment checking to verify whether the Supplemental Corpus contains information that helps the model answer correctly.
+Details are provided in the Appendix B.
 Below are the three types of CIT data:
-(2) Normative Corpus($C_{\underline{N}}$): To enable the LLM to answer multiple-choice questions in a normative format, we sample questions independent of the test set and process them into the format shown in Table 1 to build the Normative Corpus.
-Prior to supplemental training, we finetune the LLM using the Normative Corpus (i.e., pre-finetuning), which facilitates extracting answers from the LLM’s responses to calculate initial performance and allows for the collection of incorrectly answered questions.
-To maintain this ability, we incorporate a small proportion of the Normative Corpus during supplemental training.
+(2) Normative Corpus($C_{\underline{N}}$):
+To enhance the LLM’s understanding of task schemas and formats, task-relevant instruction fine-tuning is commonly applied.
+For QA tasks, we sample questions from the training set and format them as shown in Table 3 to build the Normative Corpus, enabling the LLM to answer multiple-choice questions in a standardized manner.
+Before supplemental training, we fine-tune the LLM using the Normative Corpus (i.e., pre-finetuning).
+A small portion of the Normative Corpus is incorporated during supplemental training to maintain the model’s ability to provide normative answers.
 (3) Extractive Corpus($C_{\underline{E}}$): Inspired by the human learning process, where despite studying from the same textbooks, students who excel at extracting knowledge achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
-To enhance the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample questions independent of the test set and the Normative Corpus.
-These questions are processed into the format shown in Table 2 to build the Extractive Corpus.
+To enhance the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample questions from the training set and format them as shown in Table 4 to build the Extractive Corpus.
+We employ Qwen2-72B-Instruct for entailment checking to filter the corpus.
+A minimal amount of Extractive Corpus data yields effective results, ensuring manageable inference costs.
+Details are provided in the Appendix E.
 (4) Review Corpus($C_{\underline{R}}$):
@@
 (2018).
-During repeated QA testing, the LLM may change its answers to certain ’less confident’ questions due to the proximity of predicted probabilities.
+During repeated QA testing, the LLM may change its answers to certain ‘less confident’ questions due to the proximity of predicted probabilities.
 Consequently, after supplemental training, questions initially answered correctly may be answered incorrectly, thus affecting accuracy.
@@
 (2021).
-When previous IFT data is unclear, we use the IFT data from the ShareGPT dataset [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_unfiltered as the General Corpus to preserve the model’s general capabilities on other tasks.
+When previous IFT data is unclear, we use the IFT data from the ShareGPT dataset [^8] https://huggingface.co/datasets/Aeala/ShareGPT_
+Vicuna_unfiltered
+as the General Corpus to preserve the model’s general capabilities on other tasks and instruction-following ability.
 Supplemental Training Set Construction Strategy.
@@
 4 Experimental Setup
+4.1 Training Setup
+We train the models on 2 A800 GPUs using ZeRO (Rajbhandari et al., 2020) stage 1 from DeepSpeed (Rasley et al., 2020).
+We adopt AdamW (Kingma and Ba, 2015) as the optimizer and set the batch size to 16 with a maximum sequence length of 2,048 and 1 training epoch.
+We explore the impact of full fine-tuning, LoRA fine-tuning Hu et al. (2022), and single-stage versus multi-stage training methods. For details, see the Appendix F.
 4.2 Datasets
-We conduct experiments on two question-answering datasets, including Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024).
+We conduct experiments on two question-answering datasets, including Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024a).
 To investigate the impact of catastrophic forgetting, we also use the HotpotQA Yang et al.
@@
 (2021) datasets to test the model’s decline in reasoning and mathematical abilities.
-Appendix Table 6 shows the details of the data division and usage.
+Appendix G shows the details of the data division and usage.
 4.3 Base Language Models
-To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs with parameter ranges from 6 to 13 billion.
-Our instruction fine-tuning base models include CuteGPT-13B [^2] https://huggingface.co/XuYipei/kw-cutegpt-13b-base Liang et al.
+To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs include CuteGPT-13B [^9] https://huggingface.co/XuYipei/kw-cutegpt-13b-base Liang et al.
 (2023), ChatGLM-6B [^10] https://huggingface.co/THUDM/chatglm3-6b Du et al.
@@
 ER: The Enhancement Rate measures the relative improvement brought by CEM on a specific task compared to the initial performance.
-It is calculated by $ER=\frac{R_{k}^{*}-R_{k}^{0}}{R_{i}^{0}},$ where $k$ denotes the task using CEM, $R^{0}_{k}$ is the initial accuracy, and $R^{*}_{k}$ is the accuracy after CEM.
+It is calculated by $ER=\frac{R_{k}^{*}-R_{k}^{0}}{R_{k}^{0}},$ where $k$ denotes the task using CEM, $R^{0}_{k}$ is the initial accuracy, and $R^{*}_{k}$ is the accuracy after CEM.
 AFR: The Average Forgetting Rate Wang et al.
@@
 5 Experiments
-We construct supplemental training sets and control groups for the experiments, following the strategies described in Section 3.3.
-To distinguish between models trained on different supplemental training sets, we use the following naming conventions:
-INIT represents the model that has undergone only pre-finetuning using the Normative Corpus, without any supplemental training.
-CEM-P: P denotes Plain, representing the model trained with the supplemental training set constructed by the Plain Strategy.
+We construct supplemental training sets and control groups for experiments as outlined in Section 3.3.
+To differentiate models trained on various supplemental sets, we employ the following naming conventions:
+INIT represents the model pre-finetuned solely with the Normative Corpus.
+CEM-P: ‘P’ for Plain, indicates the model trained using the supplemental training set constructed by the Plain Strategy.
 […]
 […]
-We append suffixes to indicate the data sources of the Supplemental Corpus: ‘Wiki’, ‘Bing’ and ‘Mix’.
-‘Mix’ indicates that the Supplemental Corpus is collected from both Bing and Wikipedia and concatenated.
-For example, ‘E-Wiki’ denotes the model trained using the Extractive Strategy with a Supplemental Corpus from Wikipedia.
+Suffixes ‘Wiki’, ‘Bing’, and ‘Mix’ indicate the data sources of the Supplemental Corpus, with ‘Mix’ signifying a combined set from Bing and Wikipedia.
+For instance, ‘EWiki’ denotes the model trained using the Extractive Strategy with a Supplemental Corpus from Wikipedia.
 5.1 Main Results
-We used the Plain Strategy to construct supplementary training set, using approximately 4,000 CPT data (based on the initial number of mistakes) and 4,000 CIT data for single-round CEM.
+We use the Plain Strategy to construct supplemental training set, using approximately 4,000 CPT data (based on the initial number of mistakes) and 4,000 CIT data for single-round CEM.
 The main results are shown in Figure 3.
@@
 Figure 3: This figure shows the accuracy of different models using the CEM-P method on the Xiezhi and CMMLU tasks.
-Firstly, following the supplemental training, a noticeable enhancement in accuracy is observed across all models for the two tasks.
-This indicates that the CEM method effectively addresses the model’s knowledge deficiencies, even with the small training set and readily available supplemental data.
-For instance, the ChatGLM-6B improves by up to 16.08% on the CMMLU task.
+Firstly, after supplemental training, a marked improvement in accuracy is noted across all models for both tasks.
+This suggests that the CEM method effectively addresses the model’s knowledge deficiencies, despite the limited training set and readily accessible supplemental data.
+For example, the ChatGLM-6B exhibits an increase of up to 16.08% in accuracy on the CMMLU task.
 Secondly, the effectiveness of CEM varies with the innate abilities and parameter scales of LLMs.
@@
 Additionally, the Review Corpus significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
-Secondly, Supplemental Corpus from multiple sources do not always improve the CEM method more than a single source.
-Results in Figure 4 show no significant difference in CEM effectiveness whether using Wikipedia, Bing, or a combination of both.
-Indeed, R-Mix has lower accuracy than R-Wiki and R-Bing despite having more data.
-This may be due to redundancy from overlapping sources.
-Additionally, an excess of supplemental corpus can dilute the proportion of CIT data, reducing their effectiveness in mitigating catastrophic forgetting and enhancing knowledge extraction, and increasing the risk of overfitting.
-Thirdly, proportion factor $\alpha$ has limited impact on the effectiveness of Review Corpus.
-Table 5 in Appendix shows that R0, R0.5, and R1 have similar effects on LLM performance.
+Secondly, a Supplemental Corpus from multiple sources does not invariably enhance the CEM method more than a single source.
+Results in Figure 4 indicate no substantial difference in CEM effectiveness, whether using Wikipedia, Bing, or their combination.
+Notably, RMix records lower accuracy compared to RWiki and RBing, despite a larger dataset.
+This may arise from data redundancies.
+Moreover, an excessive supplemental corpus can reduce the proportion of CIT data, undermining its role in mitigating catastrophic forgetting, improving knowledge extraction, and increasing the risk of overfitting.
+Thirdly, proportion factor $\boldsymbol{\alpha}$ has limited impact on the effectiveness of Review Corpus.
+Table 9 shows that R0, R0.5, and R1 have similar effects on LLM performance.
 This suggests that reviewing the knowledge in any format consolidates it.
@@
 5.2 Multiple Iterations Analysis
-We conduct a second round of CEM on CuteGPT-13B for the Xiezhi task, using the same training set construction strategy and data source as the first round.
-We select CIT data independent of the first round but maintain the same data volume.
-The experimental results presented in Appendix Table 8 show that after the second round of CEM, the accuracy of the LLM’s answers significantly improved across all dataset construction strategies.
-In the best case, R-Wiki shows a 4.75% improvement in accuracy compared to its previous round and a 17.00% improvement compared to INIT.
+The CEM method consistently yields significant improvements across multiple rounds, underscoring its iterative potential.
+We conduct a second round of CEM on CuteGPT-13B for the Xiezhi task, employing the same training strategies and data sources as in the first round.
+Independent CIT data of equivalent volume are used.
+Results in Table 10 indicate that this round significantly enhances the LLM’s accuracy across all dataset construction strategies.
+Notably, RWiki shows a 4.75% increase in accuracy over the previous round and a 17.00% improvement over INIT.
 Further analysis shows that the Review Strategy consistently has significantly lower R2W metrics compared to other dataset construction strategies.
-For example, the R2W metric of R-Wiki decreases by 7.02% compared to P-Wiki and by 5.90% compared to E-Wiki.
-We argue that the Review Strategy significantly reduces the probability of changing a previously correct response to an incorrect one, thereby helping the model avoid severe forgetting during multiple rounds of supplementary training.
-Figure 5: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after multiple rounds of CEM combined with the random replay method.
-Here, ER indicates the model’s enhancement rate on the Xiezhi task, while AFR represents the average forgetting rate on the HotpotQA and GSM8K tasks.
+For example, the R2W metric of RWiki decreases by 7.02% compared to PWiki and by 5.90% compared to EWiki.
+We argue that the Review Strategy significantly reduces the probability of changing a previously correct response to an incorrect one, thereby helping the model avoid severe forgetting during multiple rounds of supplemental training.
+Figure 5: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after multiple rounds of CEM.
+The left part represents results without replay, while the right part shows results with a random replay strategy.
+ER indicates the enhancement rate on the Xiezhi task, and AFR represents the average forgetting rate on the HotpotQA and GSM8K tasks.
 Catastrophic Forgetting Analysis
-To analyze catastrophic forgetting that may arise from iterations, we use CuteGPT-13B and test it on Xiezhi, HotpotQA, and GSM8K tasks to evaluate the model’s domain knowledge, reasoning, and mathematical abilities.
-We employ the CEM method to enhance performance on Xiezhi and assess any performance declines on the other two tasks.
-Initially, we sample 2,000 instances from each of the three datasets for pre-finetuning, then test to establish baseline.
-We use a method of random replay to mitigate catastrophic forgetting.
-Specifically, in each training round, we randomly replay 500 instances from both HotpotQA and GSM8k to replace the original General Corpus.
-Figure 5 shows that even after three rounds of CEM, performance on Xiezhi still exhibits significant improvement and maintains an upward trend.
-Additionally, the combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 70% in most cases.
-However, we observe that CEM-R, with its larger data capacity, achieves significant performance improvements but also shows the greatest decline on other tasks.
-This may be attributed to the increased volume of CPT data diluting the effect of the replay data.
-Combining CEM with other advanced replay-based CL methods warrants further investigation.
+To analyze catastrophic forgetting, we test CuteGPT-13B on Xiezhi, HotpotQA, and GSM8K to evaluate domain knowledge, reasoning, and mathematical skills.
+We use the CEM method to improve performance on Xiezhi and monitor any declines in the other tasks.
+Initially, we sample 2,000 instances from each dataset to establish a baseline.
+To mitigate forgetting, we employ random replay by replacing the General Corpus with 500 randomly replayed instances from HotpotQA and GSM8K in each training round.
+Figure 5 shows that even after three rounds of CEM, performance on Xiezhi improves significantly and maintains an upward trend.
+The combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 70.0% in most cases.
+However, without a replay strategy, initial rounds show better improvement (possibly because the model does not need to adapt to the more complex multitasking required by the random replay strategy), but more rounds lead to greater forgetting, limiting further iteration ,as seen when CEM-R reaches an ER of 54.4% in round three without replay.
+CEM-R, with its larger data capacity, achieves significant performance improvements but shows the greatest decline in other tasks, likely due to the increased volume of CPT data diluting the replay effect.
+Combining CEM with other advanced replay-based methods warrants further investigation.
 6 Discussion
@@
 7 Conclusion
-In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes.
+In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient and cost-effective method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes.
 We believe that the difficulty of collecting CPT data and the limitations of instruction tuning’s superficial alignment have become a bottleneck for further performance improvement.
@@
 The results consistently showing significant improvements in answer accuracy and potential for continual evolution through multiple iterations.
-In the best case, accuracy enhancement reaches up to 17.00%.
-Appendix A Format of CIT Data
+8 Limitations
+This paper focuses on the model’s performance in question-answering tasks. It remains uncertain whether CEM can consistently enhance the model’s capabilities and performance in non-QA tasks.
+Future research will focus on investigating additional rounds of supplemental training to further enhance the LLM and explore the integration of other strategies to mitigate catastrophic forgetting.
+Additionally, it will examine whether repeated supplemental learning on already mastered corpora could have adverse effects.
+Furthermore, we will experiment with diverse data sources, including synthetic corpora, and optimize strategies for collecting supplementary corpora to endow the LLM with advanced continual iterative evolutionary capabilities.
+Appendix A Knowledge Point Extraction
+Table 1 shows the prompt used for extracting knowledge points from questions using Qwen2-7B-Instruct. In the experiment, the variable $x$ in the prompt is set to $5$.
+An Example of Knowledge Point Extraction
+Instruction:
+(task)Please analyze the core knowledge points examined by the following question.
+(question)### Question:
+What is the main component of the cell membrane?
+A. Phospholipids B. Carbohydrates C. Proteins D. Nucleus
+(answer)### Answer
+The answer is A. Phospholipids. Phospholipids are the main component of the cell membrane.
+(requirement)### Requirements:
+- Prioritize identifying directly relevant named entities from the question and options.
+- Knowledge points should be closely related to the question and options, aiding in eliminating incorrect options and selecting the correct one.
+- Knowledge points should be specific, avoiding overly broad, common, or indistinct concepts, e.g., "Prime Minister", "Archaeology".
+- Return the knowledge points in a list format, e.g., ["English Civil War", "Glorious Revolution"].
+- The number of returned knowledge points should not exceed {x}.
+Output:
+(answer)["Cell Membrane Structure", "Phospholipid Bilayer"]
+Table 1: An example illustrating the extraction of knowledge points from questions.
+Appendix B Effectiveness Check of Supplemental Corpus
+Table 5 shows the prompt used for entailment checking of the Supplemental Corpus with Qwen2-72B-Instruct.
+The results are presented in Table 2, where the proportion of useful information from both data sources consistently exceeds 65%.
+Data Source Data Volume Effectiveness
+Bing 8161 0.6718
+Wiki 7694 0.6508
+Table 2: The results of verifying whether the Supplemental Corpus contains information that helps the LLM in correcting errors and answering questions accurately.
+‘Effectiveness’ denotes the proportion of useful information.
+Appendix C Process of Collecting Corpora from Multiple Sources
+The parameter configuration for the experiment is as follows: the value of $k$ is set to $4$ for top-k retrieval.
+Indexing is implemented using FAISS’s IndexFlatIP.
+The threshold for inner product similarity is set at $0.80$.
+If no entries meet this threshold, the entry with the highest similarity exceeding $0.70$ is selected; otherwise, no result is returned.
+Fig.6 and Fig.7 show the process of obtaining supplemental data from Bing and Wikipedia respectively.
+Figure 6: The process of collecting supplemental corpus from Bing.
+Figure 7: The process of collecting supplemental corpus from Wikipedia.
+Appendix D Format of CIT Data
 Table 3 and Table 4 show the IFT data formats of the Normative Corpus and the Extractive Corpus.
 An Example of Normative Corpus
-Question:
+Instruction:
 (task)Please select the correct answer for the following single choice questions:
-(question)In the 2nd century, what became the Roman Empire’s internal sea was _____
+(question)### Question:
+In the 2nd century, what became the Roman Empire’s internal sea was _____
 A. Red Sea B. Caspian Sea C. Mediterranean Sea D. Black Sea
-Answer:
-(answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
+Output:
+(answer)The answer is C. Mediterranean Sea. In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
 Table 3: An example illustrating the format of the Normative Corpus.
 An Example of Extractive Corpus
-Question:
-(task)Please select the correct answer for the following single choice questions:
-(question)In the 2nd century, what became the Roman Empire’s internal sea was _____
+Instruction:
+(task)Please select the correct answer for the following single choice questions based on the supplementary materials provided:
+(question)### Question:
+In the 2nd century, what became the Roman Empire’s internal sea was _____
 A. Red Sea B. Caspian Sea C. Mediterranean Sea D. Black Sea
-(extract prompt)Please answer in the context of the following background material:
-(supplement)The Roman Empire … During the reign of Tulajin (98-117), the Roman Empire reached its greatest extent with the Mediterranean Sea becoming the empire’s inland sea.
-In its heyday, it controlled about 5 million square kilometers of land, making it one of the largest monarchies in the ancient history of the world.
-Answer:
-(answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
+(supplement)### Supplementary Materials:
+The Roman Empire … During the reign of Tulajin (98-117), the Roman Empire reached its greatest extent with the Mediterranean Sea becoming the empire’s inland sea.
+In its heyday, it controlled about 5 million square kilometers of land, making it one of the largest monarchies in the ancient history.
+Output:
+(answer)The answer is C. Mediterranean Sea. In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
 Table 4: An example illustrating the format of the Extractive Corpus. The ellipsis indicates that some supplemental text has been omitted for clarity.
-Appendix B Process of Collecting Corpora from Multiple Sources
-Fig.6 and Fig.7 show the process of obtaining supplemental data from Wikipedia and Bing respectively.
-Figure 6: The process of collecting supplementary corpus from Bing.
-Figure 7: The process of collecting supplementary corpus from Wikipedia.
-Appendix C Experimental Results of LLMs on QA tasks
-Table 3 and Table 4 show the complete experimental results of LLMs using the CEM method on Xiezhi and CMMLU tasks. We use the Plain Strategy to construct the supplemental training dataset.
-Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
+Appendix E Entailment Checking of Extractive Corpus
+Table 5 shows the prompt used for entailment checking of the Extractive Corpus with Qwen2-72B-Instruct.
+An Example of Entailment Checking
+Instruction:
+(task)Can the correct answer to the given multiple-choice question be derived solely from the knowledge in the following material? Please respond with ’Yes’ or ’No’.
+(question)### Question:
+What is the main component of the cell membrane?
+A. Phospholipids B. Carbohydrates C. Proteins D. Nucleus
+(answer)### Answer
+The answer is A. Phospholipids. Phospholipids are the main component of the cell membrane.
+(supplement)### Materials:
+The phospholipid bilayer is a thin polar membrane made of two layers of lipid molecules.
+These membranes are flat sheets that form a continuous barrier around all cells.
+The cell membranes of almost all organisms and many viruses are made of a lipid bilayer, as are the nuclear membrane surrounding the cell nucleus, and membranes of the membrane-bound organelles in the cell…
+Output:
+(answer) Yes.
+Table 5: An example illustrating the entailment checking of the Extractive Corpus. The ellipsis indicates that some supplemental text has been omitted for clarity.
+Appendix F Exploring Fine-Tuning Strategies
+We experiment with fine-tuning strategies on Qwen-7B-Chat, employing full fine-tuning and LoRA fine-tuning for parameter updates.
+For dataset usage, we apply two approaches: single-stage parallel training on combined CPT and CIT data (3000 and 2000 samples, respectively), and multi-stage sequential training, where the model is first fine-tuned on CPT data, followed by further fine-tuning on CIT data.
+Table 6 shows that parallel full fine-tuning yields the best results.
+Therefore, in the experiments, except for ChatGLM-6B, which employs Prompt-tuning Liu et al.
+(2022), all other models use parallel full fine-tuning.
+Pre-Finetuning Parallel Sequential
+Full 0.4162 0.4854 0.3154
+LoRA - 0.4573 0.4271
+Table 6: Results of verifying the impact of CPT and CIT datasets on model performance using different parameter update strategies and training stages.
+Appendix G Data Division
+Table 7 shows the the data division of datasets.
+For Xiezhi and CMMLU, the multiple-choice questions are divided into three parts.
+The Pre-Finetuning part is utilized to train the model to answer questions in a normative format before supplemental training.
+The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
+Dataset Total Pre-Finetuning Testing Supplemental Training Random Replay
+Xiezhi 27,265 10,000 10,000 7,265 -
+CMMLU 11,917 4,000 5,000 2,917 -
+HotpotQA 7,000 2,000 5,000 - 500
+GSM8K 7,000 2,000 5,000 - 500
+Table 7: The data division and usage of Xiezhi, CMMLU, HotpotQA, and GSM8K. For HotpotQA and GSM8K, the data used for Random Replay is included in pre-finetuning.
+Appendix H Experimental Result
+Table 8 presents the comprehensive experimental results for various LLMs employing the CEM method on the Xiezhi and CMMLU tasks, with the supplemental training dataset constructed using the Plain Strategy.
+Table 9 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with supplemental training sets constructed using the Plain, Extractive, and Review strategies, as well as various data sources.
+Table 10 presents the comprehensive experimental results of CuteGPT-13B on the Xiezhi task following two rounds of CEM, with the second round employing the same training data construction strategy as the first.
+Table 11 presents the comprehensive results of catastrophic forgetting experiments for CuteGPT-13B on the Xiezhi, GSM8K, and HotpotQA tasks after one, two, and three rounds of CEM, respectively.
+Task Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
 INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
-Acc 42.93 45.91 46.56 41.62 48.71 47.87 40.44 51.95 53.61
+Xiezhi Acc 42.93 45.91 46.56 41.62 48.71 47.87 40.44 51.95 53.61
 W2R - 22.20 21.91 - 22.17 20.94 - 26.01 25.86
 R2W - 18.57 18.93 - 15.08 14.70 - 14.50 12.69
-Table 3: The table shows the experimental results of LLMs using the CEM method on the Xiezhi task.
-Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
-INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
-Acc 26.44 32.30 35.32 24.72 39.12 40.80 23.00 35.62 35.26
+CMMLU Acc 26.44 32.30 35.32 24.72 39.12 40.80 23.00 35.62 35.26
 W2R - 23.84 25.14 - 25.24 26.94 - 25.96 26.32
 R2W - 17.98 16.26 - 11.04 11.02 - 13.34 14.06
-Table 4: The table shows the experimental results of LLMs using the CEM method on the CMMLU task.
-Appendix D Experimental Results with Various supplemental Training Dataset Construction Strategies
-Table 5 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with supplemental training sets constructed using the Plain, Extractive, and Review strategies, as well as various data sources.
-Data Metrics INIT CEM-N CEM-E CEM-E’ CEM-R0.5 CEM-R1 CEM-R0 CEM-R’
-Wiki Acc 42.93 46.56 52.03 50.69 55.18 54.67 53.40 49.44
-W2R - 22.20 24.12 23.42 24.93 24.61 24.82 22.90
-R2W - 18.57 15.02 15.65 12.68 12.87 14.35 16.38
-Bing Acc 42.93 45.91 48.47 45.11 53.86 54.53 53.90 46.58
-W2R - 21.91 23.15 21.37 24.25 25.18 24.99 22.28
-R2W - 18.93 17.61 19.18 13.32 13.58 14.02 18.62
-Mix Acc 42.93 46.61 53.89 50.78 53.42 54.43 53.37 53.07
-W2R - 21.99 25.12 23.67 24.04 25.06 25.00 24.81
-R2W - 18.31 14.16 15.81 13.55 13.46 14.56 14.67
+Table 8: Experimental results of various LLMs using the CEM method on the Xiezhi and CMMLU tasks
+Data Metrics INIT CEM-N CEM-E CEM-E’ CEM-R0 CEM-R0.5 CEM-R1 CEM-R’
+Wiki Acc 42.93 46.56 52.03 50.69 53.40 55.18 54.67 49.44
+W2R - 22.20 24.12 23.42 24.82 24.93 24.61 22.90
+R2W - 18.57 15.02 15.65 14.35 12.68 12.87 16.38
+Bing Acc 42.93 45.91 48.47 45.11 53.90 53.86 54.53 46.58
+W2R - 21.91 23.15 21.37 24.99 24.25 25.18 22.28
+R2W - 18.93 17.61 19.18 14.02 13.32 13.58 18.62
+Mix Acc 42.93 46.61 53.89 50.78 53.37 53.42 54.43 53.07
+W2R - 21.99 25.12 23.67 25.00 24.04 25.06 24.81
+R2W - 18.31 14.16 15.81 14.56 13.55 13.46 14.67
 Table 9: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplemental training set construction methods.
-Appendix E Data Division
-Table 6 and Table 7 shows the the data division of Xiezhi and CMMLU.
-The multiple-choice questions are divided into three parts.
-The Pre-Finetuning part is utilized to train the model to answer questions in a standard format before supplemental training.
-The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
-Dataset Total Pre-Finetuning Testing Supplemental Training
-Xiezhi 27,265 10,000 10,000 7,265
-CMMLU 11,917 4,000 5,000 2,917
-Table 6: The data division and usage of Xiezhi and CMMLU.
-Dataset Total Pre-Finetuning Testing Random Replay
-HotpotQA 7,000 2,000 5,000 500
-GSM8K 7,000 2,000 5,000 500
-Table 7: The data division and usage of HotpotQA and GSM8K, where the data used for Random Replay is included in pre-finetuning.
-Appendix F Experimental Results of Multiple Rounds of CEM
-We employ the identical training data construction strategy as in the first round to conduct a second round of CEM on the LLM.
-Table 8 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after two rounds of CEM.
 Metrics Wiki Bing
@@
 Round1" indicates W2R and R2W calculated based on the evaluation results after one round of CEM.
-Appendix G Experimental Results of Catastrophic Forgetting Analysis
-Table 9 shows the complete experimental results of CuteGPT-13B on the Xiezhi, GSM8K, and HotpotQA tasks after one, two, and three rounds of CEM, respectively.
-Round Strategy Task Accuracy Relative Rate
-Xiezhi GSM8K HotpotQA ER AFR
+Round Strategy Xiezhi GSM8K HotpotQA ER AFR
 INIT 40.84 13.20 80.60 1.00 1.00

### v6

@@
-Large Language Model Can Continue Evolving From Mistakes
+CEM: A Data-Efficient Method for Large Language Models to Continue Evolving From Mistakes
 Abstract
-As world knowledge evolves and new task schemas emerge, Continual Learning (CL) is crucial for keeping Large Language Models (LLMs) up-to-date and addressing their shortcomings.
-LLMs typically require continual instruction tuning (CIT) and continual pre-training (CPT) to adapt to new tasks and acquire essential knowledge.
-However, collecting sufficient CPT data while addressing knowledge gaps remains challenging, as does optimizing the efficiency of utilizing this data.
+As world knowledge advances and new task schemas emerge, Continual Learning (CL) becomes essential for keeping Large Language Models (LLMs) current and addressing their shortcomings.
+This process typically involves continual instruction tuning (CIT) and continual pre-training (CPT) to enable these models to adapt to novel tasks and acquire critical knowledge.
+However, collecting sufficient CPT data and efficiently bridging knowledge gaps remain significant challenges.
 Inspired by the ‘summarizing mistakes’ strategy, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient approach aiming to collect CPT data and continually improve LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge.
-To enhance data utilization and mitigate forgetting, we introduce a novel training paradigm that combines CIT and CPT data.
-Experiments demonstrate that CEM significantly enhances model performance and continual evolution.
-The code and dataset are available in the GitHub  [^1] https://anonymous.4open.science/r/cem-BB25.
-Large Language Model Can Continue Evolving From Mistakes
+To further optimize data usage and mitigate forgetting, we introduce a novel training paradigm that combines CIT and CPT.
+Experiments show that CEM substantially enhances multiple models’ performance on both in-domain and out-of-domain QA tasks, achieving gains of up to 29.63%.
+Code and datasets are available on GitHub  [^1] https://anonymous.4open.science/r/cem-BB25.
+CEM: A Data-Efficient Method for Large Language Models to Continue Evolving From Mistakes
 Haokun Zhao1, Haixia Han2, Jie Shi1, Chengyu Du 1, Jiaqing Liang3 thanks: Corresponding Author, Yanghua Xiao 1,2 Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University1 Shanghai Institute of Artificial Intelligence for Education, East China Normal University2 Shanghai Key Laboratory of Data Science, School of Data Science, Fudan University3 {hkzhao23,jshi22}@m.fudan.edu.cn, {haixiahan03,cydu2024}@gmail.com, {liangjiaqing,shawyh}@fudan.edu.cn
@@
 Without these, models risk inducing hallucinations due to knowledge boundaries and may cause task misalignment Huang et al.
-(2023).
+(2023a).
 Additionally, addressing and correcting the shortcomings and errors exposed during practical use is crucial.
@@
 (2024) or converting raw CPT corpus into reading comprehension instructions (i.e., CIT data) Cheng et al.
-(2024) to mitigate degradation and enhance the efficiency of knowledge absorption from CPT data.
+(2024b) to mitigate degradation and enhance the efficiency of knowledge absorption from CPT data.
 Nevertheless, reliance on oracle LLMs and the complexity of mining patterns limit their practicality.
@@
 We argue that a model’s mistakes in ‘exams’ reflect its inherent knowledge deficiencies.
-Therefore, we collect corpora from the internet that contain background knowledge related to these mistakes for directly address these deficiencies.
+Therefore, we collect corpora from the internet that contain background knowledge related to these mistakes to directly address these deficiencies.
 To efficiently utilize these CPT data and mitigate forgetting, we design a novel training set construction paradigm that integrates parallel CIT and CPT data.
@@
 (2022), we enable iterative and continual model evolution.
-The CEM method consistently demonstrates significant improvements across multiple models in extensive experiments, achieving up to a 17.00% accuracy enhancement.
+The CEM method consistently demonstrates significant improvements across multiple models in extensive experiments, achieving up to a 29.63% accuracy enhancement.
 2 Related Work
@@
 AdaptLLM Cheng et al.
-(2024) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
+(2024b) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
 3 Methods
-3.1 Task Formulation
-Question Answering (QA) tasks are commonly used to evaluate LLM’s domain knowledge.
-These tasks typically involve a series of single-choice questions with multiple options.
-If an LLM correctly answers these questions, it indicates a strong grasp of the relevant knowledge.
-Conversely, incorrect answers reveal deficiencies in understanding.
-Given a question-answer example $(qi,ai)\in\mathcal{Q}$, a test set for question answering $\mathcal{Q}$, and a reasoning model $\mathcal{M}_{0}$.
-First, the reasoning model $\mathcal{M}_{0}$ is utilized to answer all questions within the test set $\mathcal{Q}$. Subsequently, questions $q_{i}$ that receive incorrect responses are collected,
-$\widetilde{r_{i}}=\mathcal{M}_{0}(q_{i}),\quad Ans(\widetilde{r_{i}})\neq a_{i},$
-where $\widetilde{r_{i}}$ denotes the response that includes incorrect answers to the question $q_{i}$, and $Ans(\cdot)$ is utilized to extract the final answer from the response.
-Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplemental corpus $C_{S_{i}}$ as CPT data,
-$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\ldots\},$
-where $Search(\cdot)$ is used to search for relevant knowledge based on the question and obtain supplemental corpus.
-Subsequently, based on the supplemental corpus $C_{S}$, construct the supplemental training set $D$ to fine-tune model $\mathcal{M}_{0}$,
-$D=Construct(C_{S}),\quad\mathcal{M}_{1}=Train(\mathcal{M}_{0},D),$
-where $Construct(\cdot)$ is utilized to build an supplemental training set, $Train(\cdot)$ denotes fine-tuning.
-$\mathcal{M}_{1}$, represents the model that has undergone one iteration of supplemental training.
-We can continue to employ $\mathcal{M}_{1}$ as the reasoning model on the test set $\mathcal{Q}$, for iterative supplemental training, thereby achieving continual optimization of the model’s performance.
 3.1 CEM Method
@@
 3.2 Supplemental Training Data
-Definition of Corpus Type. The construction of the supplemental training set involves the following five types of corpus.
+Definition of Data Type. The construction of the supplemental training set involves the following five types of data.
 (1) Supplemental Corpus($C_{\underline{S}}$): The supplemental corpus, obtained from multiple data sources, contains knowledge that the LLM lacks and is used as CPT data.
@@
 Details are provided in the Appendix B.
-Below are the three types of CIT data:
-(2) Normative Corpus($C_{\underline{N}}$):
+Below are the four types of CIT data:
+(2) Normative Instruction($I_{\underline{N}}$):
 To enhance the LLM’s understanding of task schemas and formats, task-relevant instruction fine-tuning is commonly applied.
-For QA tasks, we sample questions from the training set and format them as shown in Table 3 to build the Normative Corpus, enabling the LLM to answer multiple-choice questions in a standardized manner.
-Before supplemental training, we fine-tune the LLM using the Normative Corpus (i.e., pre-finetuning).
-A small portion of the Normative Corpus is incorporated during supplemental training to maintain the model’s ability to provide normative answers.
-(3) Extractive Corpus($C_{\underline{E}}$): Inspired by the human learning process, where despite studying from the same textbooks, students who excel at extracting knowledge achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
-To enhance the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample questions from the training set and format them as shown in Table 4 to build the Extractive Corpus.
-We employ Qwen2-72B-Instruct for entailment checking to filter the corpus.
-A minimal amount of Extractive Corpus data yields effective results, ensuring manageable inference costs.
+For QA tasks, we sample questions from the training set and format them as shown in Table 4 to build the Normative Instruction, enabling the LLM to answer multiple-choice questions in a standardized manner.
+Before supplemental training, we fine-tune the LLM using the Normative Instruction (i.e., pre-finetuning).
+A small portion of the Normative Instruction is incorporated during supplemental training to maintain the model’s ability to provide normative answers.
+(3) Extractive Instruction($I_{\underline{E}}$): Inspired by the human learning process, where despite studying from the same textbooks, students who excel at extracting knowledge achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
+To enhance the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we restructured a subset of training set questions into an information-extraction question format, as illustrated in Table 5.
+We employ Qwen2-72B-Instruct for entailment checking to filter the instructions.
+A minimal amount of Extractive Instruction yields effective results, ensuring manageable inference costs.
 Details are provided in the Appendix E.
-(4) Review Corpus($C_{\underline{R}}$):
+(4) Review Instruction($I_{\underline{R}}$):
 The responses of the LLM tend to be diverse under high temperature settings Caccia et al.
@@
 To reinforce the LLM’s memory of the knowledge implied by correctly answered questions, we sample the responses provided correctly by the LLM in the previous round.
-We process some of these responses in the Normative Corpus format and the rest in the Extractive Corpus format to build the Review Corpus.
-The composition of the Review Corpus is described as follows:
-$C_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
-where $\alpha\in[0,1]$, denotes the proportion of responses processed into the Normative Corpus format.
-$r$ denotes the correct responses in the previous round.
-$Normative(\cdot)$ converts responses into the Normative Corpus format, and $Extractive(\cdot)$ converts responses into the Extractive Corpus format.
-(5) General Corpus($C_{\underline{G}}$): We replay previous IFT data to mitigate catastrophic forgetting when training LLMs on new data He et al.
+We process some of these responses in the Normative Instruction format and the rest in the Extractive Instruction format to build the Review Instruction.
+The composition of the Review Instruction is defined as:
+$I_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
+where $\alpha\in[0,1]$ represents the proportion of responses in the Normative Instruction format, and $r$ denotes the correct responses from the previous round.
+$Normative(\cdot)$ and $Extractive(\cdot)$ convert responses into the Normative and Extractive Instruction formats, respectively.
+In subsequent experiments, we examine the effect of varying $\alpha$.
+(5) General Instruction($I_{\underline{G}}$): We replay previous IFT data to mitigate catastrophic forgetting when training LLMs on new data He et al.
 (2021).
@@
 Vicuna_unfiltered
-as the General Corpus to preserve the model’s general capabilities on other tasks and instruction-following ability.
+as the General Instruction to preserve the model’s general capabilities on other tasks and instruction-following ability.
 Supplemental Training Set Construction Strategy.
 We propose a novel CL dataset construction paradigm that parallelizes CIT and CPT data.
-Specifically, we construct three types of supplemental training sets by combining five types of corpora.
-The following sections detail these construction strategies.
-(1) Plain Strategy:
-$D_{\underline{P}}=Concat(C_{S},C_{N},C_{G}),$
-where $C_{S}$ denotes the Supplemental Corpus, $C_{N}$ denotes the Normative Corpus, $C_{G}$ denotes the General Corpus, $Concat{(\cdot)}$ merges these corpora, and $D_{P}$ represents the supplemental training set constructed using the Plain Strategy.
-(2) Extractive Strategy:
-$D_{\underline{E}}=Concat(C_{S},C_{N},C_{G},{\color[rgb]{1,0,0}C_{E}}),$
-where $C_{E}$ denotes the Extractive Corpus and $D_{E}$ represents the supplemental training set constructed using the Extractive Strategy.
-(3) Review Strategy:
-$D_{\underline{R}}=Concat(C_{S},C_{N},C_{G},C_{E},{\color[rgb]{1,0,0}C_{R}}),$
-where $C_{R}$ denotes the Review Corpus and $D_{R}$ represents the supplemental training set constructed using the Review Strategy.
+Specifically, we develop three supplemental training set construction strategies—Plain, Extractive, and Review—by concatenating some or all of five types of data, as shown in Step 5 of Figure 2.
+These strategies are named to reflect the specific data components incorporated.
 4 Experimental Setup
@@
 We train the models on 2 A800 GPUs using ZeRO (Rajbhandari et al., 2020) stage 1 from DeepSpeed (Rasley et al., 2020).
-We adopt AdamW (Kingma and Ba, 2015) as the optimizer and set the batch size to 16 with a maximum sequence length of 2,048 and 1 training epoch.
+We adopt AdamW (Kingma and Ba, 2015) as the optimizer and set the batch size to 32 with a maximum sequence length of 2,048 and 1 training epoch.
 We explore the impact of full fine-tuning, LoRA fine-tuning Hu et al. (2022), and single-stage versus multi-stage training methods. For details, see the Appendix F.
 4.2 Datasets
-We conduct experiments on two question-answering datasets, including Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024a).
+We conduct experiments on two question-answering datasets, Xiezhi Gu et al.
+(2023) and CMMLU Li et al.
+(2024a).
+To evaluate CEM’s out-of-domain generalization, we apply CEM on the in-domain dataset Xiezhi and use CMMLU and C-Eval Huang et al.
+(2023b) as out-of-domain benchmarks.
 To investigate the impact of catastrophic forgetting, we also use the HotpotQA Yang et al.
@@
 (2021) datasets to test the model’s decline in reasoning and mathematical abilities.
-Appendix G shows the details of the data division and usage.
 4.3 Base Language Models
-To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs include CuteGPT-13B [^9] https://huggingface.co/XuYipei/kw-cutegpt-13b-base Liang et al.
-(2023), ChatGLM-6B [^10] https://huggingface.co/THUDM/chatglm3-6b Du et al.
-(2022), and Qwen-7B-Chat [^11] https://huggingface.co/Qwen/Qwen-7B-Chat Bai et al.
+To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, particularly for smaller parameter-scale models, we select open-source LLMs, including Qwen1.5-7B-Chat [^9] https://huggingface.co/Qwen/Qwen1.5-7B-Chat Bai et al.
+(2023), Llama3-8B-Instruct [^10] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct AI@Meta (2024), and CuteGPT-13B-ift [^11] https://huggingface.co/XuYipei/kw-cutegpt-13b-ift Liang et al.
 (2023).
-These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
+Each model has been fine-tuned on the IFT data to ensure it effectively follows instructions.
 4.4 Metrics
@@
 (2024) measures the average relative degradation brought by CEM on tasks, excluding the task using CEM, compared to their initial performance.
-It is calculated by $AFR=\frac{1}{N-1}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\left(\frac{R_{i}^{0}-R_{i}^{*}}{R_{i}^{0}}\right)$, where $R^{0}_{i}$ is the initial accuracy of task $i$ after the initial fine-tuning, and $R_{i}^{*}$ is the accuracy of taks $i$ after applying CEM on task $k$.
+It is calculated by $AFR=\frac{1}{N-1}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\left(\frac{R_{i}^{0}-R_{i}^{*}}{R_{i}^{0}}\right)$, where $R^{0}_{i}$ is the initial accuracy of task $i$ after the initial fine-tuning, and $R_{i}^{*}$ is the accuracy of tasks $i$ after applying CEM on task $k$.
 5 Experiments
 We construct supplemental training sets and control groups for experiments as outlined in Section 3.2.
-To differentiate models trained on various supplemental sets, we employ the following naming conventions:
-INIT represents the model pre-finetuned solely with the Normative Corpus.
-CEM-P: ‘P’ for Plain, indicates the model trained using the supplemental training set constructed by the Plain Strategy.
+To distinguish models trained on different supplemental sets, we use the following naming conventions:
+INIT represents the model pre-finetuned solely with the Normative Instruction, equipping it with the ability to generate standardized responses to multiple-choice questions.
+All subsequent supplemental training is conducted based on this model.
+CEM-P: P for Plain, referring to the model trained using the supplemental set constructed with the Plain Strategy.
 […]
 […]
-Suffixes ‘Wiki’, ‘Bing’, and ‘Mix’ indicate the data sources of the Supplemental Corpus, with ‘Mix’ signifying a combined set from Bing and Wikipedia.
-For instance, ‘EWiki’ denotes the model trained using the Extractive Strategy with a Supplemental Corpus from Wikipedia.
+We also consider several baselines for comparison with our algorithm as follows:
+TempWiki: TemporalWiki uses Wikipedia snapshots as CPT data Jang et al.
+(2022).
+We construct the supplemental training set using corpus from TemporalWiki [^12] https://huggingface.co/datasets/seonghyeonye/TemporalWiki, broadly associated with the evaluation set rather than focusing solely on error cases.
+AdaptLLM: AdaptLLM continues to train the LLMs on the reading comprehension texts constructed based on the raw corpora, mixed with general instructionsCheng et al. (2024b).
+InstructPT: Instruction Pre-Training utilizes an instruction synthesizer [^13] https://huggingface.co/instruction-pretrain to convert raw texts from pre-training corpora into instruction-response pairs for continual tuning Cheng et al.
+(2024a).
 5.1 Main Results
-We use the Plain Strategy to construct supplemental training set, using approximately 4,000 CPT data (based on the initial number of mistakes) and 4,000 CIT data for single-round CEM.
-The main results are shown in Figure 3.
+We apply the CEM method on the Xiezhi task, collecting approximately 60,000 supplemental corpus samples sourced from Wikipedia.
+From this corpus, we downsample 25,000 samples as CPT data, while the remaining four types of CIT data are each set to 2,000 samples.
+To ensure fair comparisons, we add 2,000 Normative Instructions to each of the other baselines, keeping the quantity of raw corpus data consistent.
+Table 1 presents the experimental results, with CMMLU and C-Eval serving as out-of-domain (OOD) benchmarks to evaluate the model’s performance on unseen QA tasks.
 Our observations are summarized as follows:
-Figure 3: This figure shows the accuracy of different models using the CEM-P method on the Xiezhi and CMMLU tasks.
-Firstly, after supplemental training, a marked improvement in accuracy is noted across all models for both tasks.
-This suggests that the CEM method effectively addresses the model’s knowledge deficiencies, despite the limited training set and readily accessible supplemental data.
-For example, the ChatGLM-6B exhibits an increase of up to 16.08% in accuracy on the CMMLU task.
-Secondly, the effectiveness of CEM varies with the innate abilities and parameter scales of LLMs.
-For example, on the Xiezhi task, ChatGLM-6B, with the fewest parameters, improves by up to 13.17%.
-In contrast, CuteGPT-13B, with the highest initial accuracy and most parameters, shows a smaller improvement of up to 3.63%.
-Construction Strategy Analysis.
-To investigate the effectiveness of each component in the supplemental training set, we conducted experiments on CuteGPT-13B with various dataset construction strategies.
-We set the volume of both Extractive Corpus and Review Corpus to 1,000.
-The results are shown in Figure 4.
-Below is a summary of our observations:
+Benchmark INIT Ours Others
+Acc(%) Prompt CEM-P CEM-E CEM-R1 TempWiki AdaptLLM InstructPT
+Qwen1.5-7B-Chat
+Xiezhi 42.41 51.52 52.12 56.87 30.15 35.14 49.05
+CMMLU 36.08 36.30 37.44 40.34 24.22 21.12 38.50
+C-Eval 35.00 37.30 40.34 40.90 27.10 16.50 36.80
+Llama3-8B-Instruct
+Xiezhi 25.09 35.98 40.52 54.72 32.24 41.63 42.46
+CMMLU 23.00 24.92 29.76 33.92 23.40 27.62 31.30
+C-Eval 26.00 26.20 30.20 31.40 22.90 25.30 32.50
+CuteGPT-13B-ift
+Xiezhi 45.70 56.82 57.03 58.67 50.80 45.67 54.04
+CMMLU 32.56 37.96 37.42 38.26 33.44 32.32 35.22
+C-Eval 31.50 35.00 35.20 36.10 31.60 28.50 33.00
+Table 1: The main experimental results of our methods and baseline approaches across various tasks are presented.
+Experiments are conducted using three different LLMs: Qwen1.5-7B-Chat, Llama3-8B-Instruct and CuteGPT-13B-ift.
+The top two performances are highlighted in red bold and black bold, respectively.
+Firstly, CEM leads to significant accuracy improvements across all models and tasks.
+All variations of CEM substantially enhance model performance in both the in-domain Xiezhi task and OOD tasks.
+For instance, the Llama3-8B-Instruct model achieves accuracy gains of up to 29.63% on the Xiezhi task, 10.92% on the CMMLU task, and 5.40% on the C-Eval task when using CEM-R.
+This highlights the effectiveness of the CEM method in addressing the model’s knowledge gaps and its capacity to generalize this knowledge to unseen QA tasks.
+Secondly, CEM demonstrates superior stability compared to other baselines.
+Both TempWiki and AdaptLLM occasionally result in performance degradation for LLMs.
+For instance, AdaptLLM reduces the accuracy of the Qwen1.5-7B-Chat model by 7.27% on the Xiezhi task, 14.96% on the CMMLU task, and 18.50% on the C-Eval task.
+In contrast, all variations of CEM consistently deliver positive performance improvements.
+Thirdly, Extractive Instruction and Review Instruction further enhance performance.
+A clear trend of incremental performance improvement is observed from CEM-P to CEM-E and finally to CEM-R.
+The ablative properties introduced during dataset construction suggest that these two types of CIT data play a crucial role in driving the observed improvements.
+5.2 Data Source Analysis
+We analyze the impact of the Supplementary Corpus from different sources—Wikipedia, Bing, and their concatenation—on CEM’s performance.
+To further validate CEM’s generalizability, CMMLU is included as an in-domain task scenario with data proportions consistent with the main experiment.
+Figure 3: This figure presents the accuracy of models using the CEM-P method with different data sources on Xiezhi and CMMLU tasks.
+The suffixes Wiki, Bing, and Mix indicate the sources of the Supplementary Corpus, with Mix representing using double the samples sourced from the combined dataset of Wikipedia and Bing.
+We find that the CEM method shows robust compatibility across different data sources, consistently improving model performance.
+However, integrating multiple sources does not necessarily outperform using a single source.
+As shown in Figure 3, no significant differences in CEM effectiveness emerge when employing Wikipedia, Bing, or their combination.
+Although the Mix dataset is larger, it often yields lower accuracy than Wiki or Bing, potentially due to data redundancies.
+Moreover, an excessively large supplemental corpus can dilute the influence of CIT data, reducing its capacity to mitigate catastrophic forgetting, support knowledge extraction, and prevent overfitting.
+5.3 Mechanistic Analysis of Extractive and Review Instruction Components
+To clarify the underlying mechanisms by which Extractive Instruction and Review Instruction affect model performance, we conducted experiments on CuteGPT-13B-ift and examined the R2W and W2R metrics.
+Additionally, we explored the effects of varying $\alpha$ (with values $[0,0.5,1]$) on CEM-R to determine whether more diverse and complex Review Instructions yield improved results.
+The corresponding results are illustrated in Figure 4 and Figure 5.
+We summarize the key observations and insights derived from these analyses:
+(1) CEM-E and CEM-R outperform their ablation groups with General Instruction, demonstrating the effectiveness and positive impact of Extractive and Review Instruction on model performance.
+As shown in Figure 5, the Extractive Instruction enhances W2R performance: compared to CEM-P, CEM-E achieves improvements of 1.45%, 3.22%, and 1.26% across the three models, indicating more effective error correction.
+In addition, the Review Instruction significantly reduces the R2W metric, reflecting a decreased tendency for the model to forget previously acquired correct knowledge.
+While the performance benefits of the Review Instruction are readily understandable, the exact mechanism by which the Extractive Instruction improves the model’s ability to integrate crucial knowledge from lengthy passages warrants further discussion.
+We propose the following explanation: The Extractive Instruction provides the model with both the questions and supplementary materials containing relevant information.
+Questions and answer options typically present key points in a concise manner, whereas the supplementary materials often contain a dense network of factual statements Jiang et al.
+(2024).
+Thus, the questions act as a structured induction of the supplementary materials, explicitly linking complex information to the targeted induction (i.e., the questions and options).
+Serving as cues and guides, the questions and options help the model retrieve pertinent information from the corpus while filtering out redundant or irrelevant details.
+This process enhances the model’s ability to efficiently extract and utilize essential information.
+(1) CEM-E and CEM-R outperform their ablation groups with General Instruction, demonstrating the effectiveness and positive impact of Extractive and Review Instruction on model performance.
+Figure 5 shows that Extractive Instruction improves W2R, with CEM-E achieving gains of 1.45%, 3.22%, and 1.26% across the three models compared to CEM-P, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the CPT data.
+This improvement leads to more effective error correction.
+Additionally, the Review Instruction significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
+(2) The proportion factor $\boldsymbol{\alpha}$ has minimal impact on the effectiveness of Review Instruction, as R0, R0.5, and R1 exhibit similar effects on LLM performance.
+This indicates that reviewing knowledge, regardless of format, effectively consolidates it.
 […]
-Firstly, Extractive and Review Corpus significantly enhance LLM performance.
-CEM-E consistently outperforms its control group, showing higher accuracy and W2R metrics.
-Similarly, CEM-R outperforms its control group and significantly reduces the R2W metric.
-The results indicate an improvement in the W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the CPT data.
-This improvement leads to more effective error correction.
-Additionally, the Review Corpus significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
-Secondly, a Supplemental Corpus from multiple sources does not invariably enhance the CEM method more than a single source.
-Results in Figure 4 indicate no substantial difference in CEM effectiveness, whether using Wikipedia, Bing, or their combination.
-Notably, RMix records lower accuracy compared to RWiki and RBing, despite a larger dataset.
-This may arise from data redundancies.
-Moreover, an excessive supplemental corpus can reduce the proportion of CIT data, undermining its role in mitigating catastrophic forgetting, improving knowledge extraction, and increasing the risk of overfitting.
-Thirdly, proportion factor $\boldsymbol{\alpha}$ has limited impact on the effectiveness of Review Corpus.
-Table 9 shows that R0, R0.5, and R1 have similar effects on LLM performance.
-This suggests that reviewing the knowledge in any format consolidates it.
-Therefore, unless specified, R will refer to the R0.5 in the following.
+Figure 5: The figure presents the R2W and W2R metrics of the models after CEM supplemental training.
 5.4 Multiple Iterations Analysis
-The CEM method consistently yields significant improvements across multiple rounds, underscoring its iterative potential.
-We conduct a second round of CEM on CuteGPT-13B for the Xiezhi task, employing the same training strategies and data sources as in the first round.
-Independent CIT data of equivalent volume are used.
-Results in Table 10 indicate that this round significantly enhances the LLM’s accuracy across all dataset construction strategies.
-Notably, RWiki shows a 4.75% increase in accuracy over the previous round and a 17.00% improvement over INIT.
-Further analysis shows that the Review Strategy consistently has significantly lower R2W metrics compared to other dataset construction strategies.
-For example, the R2W metric of RWiki decreases by 7.02% compared to PWiki and by 5.90% compared to EWiki.
-We argue that the Review Strategy significantly reduces the probability of changing a previously correct response to an incorrect one, thereby helping the model avoid severe forgetting during multiple rounds of supplemental training.
-Figure 5: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after multiple rounds of CEM.
+To evaluate the iterability of the CEM method and its potential for catastrophic forgetting during multiple training rounds, we conduct three rounds of CEM training on CuteGPT-13B-ift for the Xiezhi task, using consistent training strategies and non-repetitive CIT data of equivalent size in each round.
+Specifically, the CEM method is applied to the Xiezhi task, while HotpotQA and GSM8K are used to assess reasoning and mathematical skills.
+The ER metric evaluates performance improvements on Xiezhi, and the AFR metric monitors performance degradation in other tasks.
+Initially, 2,000 instances from each dataset are sampled for pre-finetuning to establish a baseline.
+To address forgetting, we implement a random replay strategy, replacing General Instructions with 500 randomly replayed instances from HotpotQA and GSM8K in each training round.
+Figure 6: The figure shows the experimental results of CuteGPT-13B-ift on the Xiezhi task after multiple rounds of CEM.
 The left part represents results without replay, while the right part shows results with a random replay strategy.
 ER indicates the enhancement rate on the Xiezhi task, and AFR represents the average forgetting rate on the HotpotQA and GSM8K tasks.
-Catastrophic Forgetting Analysis
-To analyze catastrophic forgetting, we test CuteGPT-13B on Xiezhi, HotpotQA, and GSM8K to evaluate domain knowledge, reasoning, and mathematical skills.
-We use the CEM method to improve performance on Xiezhi and monitor any declines in the other tasks.
-Initially, we sample 2,000 instances from each dataset to establish a baseline.
-To mitigate forgetting, we employ random replay by replacing the General Corpus with 500 randomly replayed instances from HotpotQA and GSM8K in each training round.
 Figure 6 shows that even after three rounds of CEM, performance on Xiezhi improves significantly and maintains an upward trend.
-The combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 70.0% in most cases.
-However, without a replay strategy, initial rounds show better improvement (possibly because the model does not need to adapt to the more complex multitasking required by the random replay strategy), but more rounds lead to greater forgetting, limiting further iteration ,as seen when CEM-R reaches an ER of 54.4% in round three without replay.
+The combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 75.0% in most cases, highlighting its iterative potential.
+However, without a replay strategy, initial rounds show better improvement (possibly because the model does not need to adapt to the more complex multitasking required by random replay), but more rounds lead to greater forgetting, limiting further iteration ,as seen when CEM-R reaches an ER of 62.4% in round three without replay.
 CEM-R, with its larger data capacity, achieves significant performance improvements but shows the greatest decline in other tasks, likely due to the increased volume of CPT data diluting the replay effect.
 Combining CEM with other advanced replay-based methods warrants further investigation.
-6 Discussion
-Why can the Extractive Corpus be effective?
-LLMs using the Extractive Corpus show higher accuracy and W2R metrics.
-Unlike the Review Corpus, where the role in performance improvement is intuitive, the precise mechanism by which the Extractive Corpus augments the model’s capability to capture pivotal knowledge in lengthy texts remains worthy of discussion.
-Extractive Corpus provides the model with both questions and supplemental material containing relevant knowledge.
-Questions and options typically present straightforwardly and can be condensed into several key points, while the supplemental material tends to intricately interweave numerous factual statements Jiang et al.
-(2024).
-Thus, we suggest that questions act as a structured induction of the supplemental material, where the Extractive Corpus explicitly links complex information to the resulting induction (i.e., the question and options).
-Questions and options can serve as cues and guides for the model to gather relevant information from the corpus, helping it consider how this knowledge is accessed through questions.
-This process enhances the model’s ability to extract useful information from the corpus while filtering out redundant and irrelevant details.
 6 Conclusion
 In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient and cost-effective method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes.
-We believe that the difficulty of collecting CPT data and the limitations of instruction tuning’s superficial alignment have become a bottleneck for further performance improvement.
-To address this, we design a pipeline to collect corpus from mistakes and a novel parallel dataset construction paradigm using both CIT and CPT data to achieve continual and efficient internal knowledge supplementation during fine-tuning stage.
-We conduct experiments on several open-source LLMs to validate the efficacy of CEM.
-The results consistently showing significant improvements in answer accuracy and potential for continual evolution through multiple iterations.
+CEM employs a pipeline that leverages model mistakes to gather extensive CPT data, combined with a novel parallel training set construction paradigm integrating both CIT and CPT data for continuous training.
+Experiments on various open-source LLMs show that CEM significantly improves answer accuracy and demonstrates strong potential for sustained evolution through multiple iterations.
 7 Limitations
@@
 Appendix D Format of CIT Data
-Table 3 and Table 4 show the IFT data formats of the Normative Corpus and the Extractive Corpus.
-An Example of Normative Corpus
+Table 4 and Table 5 show the IFT data formats of the Normative Instruction and the Extractive Instruction.
+An Example of Normative Instruction
 Instruction:
@@
 (answer)The answer is C. Mediterranean Sea. In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
-Table 3: An example illustrating the format of the Normative Corpus.
-An Example of Extractive Corpus
+Table 4: An example illustrating the format of the Normative Instruction.
+An Example of Extractive Instruction
 Instruction:
@@
 (answer)The answer is C. Mediterranean Sea. In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
-Table 4: An example illustrating the format of the Extractive Corpus. The ellipsis indicates that some supplemental text has been omitted for clarity.
-Appendix E Entailment Checking of Extractive Corpus
-Table 5 shows the prompt used for entailment checking of the Extractive Corpus with Qwen2-72B-Instruct.
+Table 5: An example illustrating the format of the Extractive Instruction. The ellipsis indicates that some supplemental text has been omitted for clarity.
+Appendix E Entailment Checking of Extractive Instruction
+Table 6 shows the prompt used for entailment checking of the Extractive Instruction with Qwen2-72B-Instruct.
 An Example of Entailment Checking
@@
 (answer) Yes.
-Table 5: An example illustrating the entailment checking of the Extractive Corpus. The ellipsis indicates that some supplemental text has been omitted for clarity.
+Table 6: An example illustrating the entailment checking of the Extractive Instruction. The ellipsis indicates that some supplemental text has been omitted for clarity.
 Appendix F Exploring Fine-Tuning Strategies
@@
 Table 7: Results of verifying the impact of CPT and CIT datasets on model performance using different parameter update strategies and training stages.
-Appendix G Data Division
-Table 7 shows the the data division of datasets.
-For Xiezhi and CMMLU, the multiple-choice questions are divided into three parts.
-The Pre-Finetuning part is utilized to train the model to answer questions in a normative format before supplemental training.
-The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
-Dataset Total Pre-Finetuning Testing Supplemental Training Random Replay
-Xiezhi 27,265 10,000 10,000 7,265 -
-CMMLU 11,917 4,000 5,000 2,917 -
-HotpotQA 7,000 2,000 5,000 - 500
-GSM8K 7,000 2,000 5,000 - 500
-Table 7: The data division and usage of Xiezhi, CMMLU, HotpotQA, and GSM8K. For HotpotQA and GSM8K, the data used for Random Replay is included in pre-finetuning.
-Appendix H Experimental Result
-Table 8 presents the comprehensive experimental results for various LLMs employing the CEM method on the Xiezhi and CMMLU tasks, with the supplemental training dataset constructed using the Plain Strategy.
-Table 9 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with supplemental training sets constructed using the Plain, Extractive, and Review strategies, as well as various data sources.
-Table 10 presents the comprehensive experimental results of CuteGPT-13B on the Xiezhi task following two rounds of CEM, with the second round employing the same training data construction strategy as the first.
-Table 11 presents the comprehensive results of catastrophic forgetting experiments for CuteGPT-13B on the Xiezhi, GSM8K, and HotpotQA tasks after one, two, and three rounds of CEM, respectively.
-Task Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
-INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
-Xiezhi Acc 42.93 45.91 46.56 41.62 48.71 47.87 40.44 51.95 53.61
-W2R - 22.20 21.91 - 22.17 20.94 - 26.01 25.86
-R2W - 18.57 18.93 - 15.08 14.70 - 14.50 12.69
-CMMLU Acc 26.44 32.30 35.32 24.72 39.12 40.80 23.00 35.62 35.26
-W2R - 23.84 25.14 - 25.24 26.94 - 25.96 26.32
-R2W - 17.98 16.26 - 11.04 11.02 - 13.34 14.06
-Table 8: Experimental results of various LLMs using the CEM method on the Xiezhi and CMMLU tasks
-Data Metrics INIT CEM-N CEM-E CEM-E’ CEM-R0 CEM-R0.5 CEM-R1 CEM-R’
-Wiki Acc 42.93 46.56 52.03 50.69 53.40 55.18 54.67 49.44
-W2R - 22.20 24.12 23.42 24.82 24.93 24.61 22.90
-R2W - 18.57 15.02 15.65 14.35 12.68 12.87 16.38
-Bing Acc 42.93 45.91 48.47 45.11 53.90 53.86 54.53 46.58
-W2R - 21.91 23.15 21.37 24.99 24.25 25.18 22.28
-R2W - 18.93 17.61 19.18 14.02 13.32 13.58 18.62
-Mix Acc 42.93 46.61 53.89 50.78 53.37 53.42 54.43 53.07
-W2R - 21.99 25.12 23.67 25.00 24.04 25.06 24.81
-R2W - 18.31 14.16 15.81 14.56 13.55 13.46 14.67
-Table 9: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplemental training set construction methods.
-Metrics Wiki Bing
-CEM-N CEM-E CEM-R CEM-N CEM-E CEM-R
-Acc INIT 42.93 42.93 42.93 42.93 42.93 42.93
-Round1 46.56 52.02 55.18 45.91 48.46 53.86
-Round2 51.69 55.20 59.93 46.37 48.74 58.33
-W2R vs. INIT 24.46 26.03 28.29 22.07 23.32 27.01
-vs. Round1 20.00 17.56 16.43 18.49 18.33 16.61
-R2W vs. INIT 15.70 13.76 11.29 18.63 17.51 11.61
-vs. Round1 14.87 14.39 11.68 18.03 18.06 12.14
-Table 10: The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task, where "v.s.
-INIT" indicates W2R and R2W calculated based on the initial evaluation results, and "v.s.
-Round1" indicates W2R and R2W calculated based on the evaluation results after one round of CEM.
-Round Strategy Xiezhi GSM8K HotpotQA ER AFR
-INIT 40.84 13.20 80.60 1.00 1.00
-Round 1 CEM-P 42.02 6.64 51.60 1.03 0.730
-CEM-E 42.76 8.36 55.52 1.05 0.681
-CEM-R 51.49 7.54 55.62 1.26 0.673
-Round 2 CEM-P 48.68 9.94 57.84 1.19 0.723
-CEM-E 49.47 9.76 60.24 1.21 0.746
-CEM-R 54.23 6.88 56.68 1.33 0.703
-Round 3 CEM-P 50.58 10.02 60.96 1.24 0.757
-CEM-E 52.69 9.46 62.98 1.29 0.772
-CEM-R 58.91 7.48 59.86 1.44 0.678
-Table 11: The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi, GSM8K, and HotpotQA tasks.

### v7

@@
 As world knowledge advances and new task schemas emerge, Continual Learning (CL) becomes essential for keeping Large Language Models (LLMs) current and addressing their shortcomings.
-This process typically involves continual instruction tuning (CIT) and continual pre-training (CPT) to enable these models to adapt to novel tasks and acquire critical knowledge.
+This process typically involves continual instruction tuning (CIT) and continual pre-training (CPT) to enable these models to adapt to novel tasks and acquire supplementary knowledge.
 However, collecting sufficient CPT data and efficiently bridging knowledge gaps remain significant challenges.
@@
 CEM: A Data-Efficient Method for Large Language Models to Continue Evolving From Mistakes
-Haokun Zhao1, Haixia Han2, Jie Shi1, Chengyu Du 1, Jiaqing Liang3 thanks: Corresponding Author, Yanghua Xiao 1,2 Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University1 Shanghai Institute of Artificial Intelligence for Education, East China Normal University2 Shanghai Key Laboratory of Data Science, School of Data Science, Fudan University3 {hkzhao23,jshi22}@m.fudan.edu.cn, {haixiahan03,cydu2024}@gmail.com, {liangjiaqing,shawyh}@fudan.edu.cn
+Haokun Zhao1, Jinyi Han2, Jie Shi1, Chengyu Du 1, Jiaqing Liang3 thanks: Corresponding Author, Yanghua Xiao 1,2 Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University1 Shanghai Institute of Artificial Intelligence for Education, East China Normal University2 Shanghai Key Laboratory of Data Science, School of Data Science, Fudan University3 {hkzhao23,jshi22}@m.fudan.edu.cn, {haixiahan03,cydu2024}@gmail.com, {liangjiaqing,shawyh}@fudan.edu.cn
 1 Introduction
@@
 However, in real-world applications, the continual emergence of new knowledge and evolving task requirements necessitate ongoing updates and task-specific adaptations for LLMs.
-Without these, models risk inducing hallucinations due to knowledge boundaries and may cause task misalignment Huang et al.
+Without these, models risk hallucinating due to their limited knowledge boundaries and may cause task misalignment Huang et al.
 (2023a).
@@
 (2023).
-Despite CIT’s smaller data size requirement and lower risk of overfitting compared to CPT, it is limited in new knowledge injection, as Zhou et al.
-(2023) highlight that instruction tuning is a superficial alignment focused on interaction styles (i.e., task schema) and does not effectively help the model acquire new knowledge.
-Thus, it cannot address all the issues shown in Figure 1, where LLMs’ poor performance may stem from a lack of requisite knowledge rather than unfamiliarity with the task schema Ren et al.
+Although CIT requires less data and poses a lower overfitting risk than CPT, it is less effective at injecting new knowledge.
+As Zhou et al.
+(2023) highlight, instruction tuning is a superficial alignment focused on interaction styles (i.e., task schema) and is not well-suited for large-scale knowledge supplementation.
+This is because scaling up the necessary task-specific instruction data significantly increases complexity in terms of data selection, careful prompt design, and ensuring coverage across extensive knowledge domains.
+Consequently, CIT cannot address all issues shown in Figure 1, where LLM underperformance may arise from a lack of requisite knowledge rather than unfamiliarity with the task schema Ren et al.
 (2023); Zhang et al.
 (2023b).
-In contrast, CPT is more effective in addressing the lack of knowledge.
+In contrast, CPT better addresses such knowledge deficits.
 However, using CPT to supplement the model with lacking knowledge still presents several challenges.
@@
 (2024b).
-Second, efficiently utilizing CPT data and mitigating degradation are difficult due to weak constraints on parameter updates, leading to larger parameter shifts and a data size gap compared to the initial pre-training.
-To address these issues, the most common data collection methods involve using the latest Wikipedia snapshots or new academic papers in specialized fields Cossu et al.
+Second, efficiently utilizing CPT data and mitigating degradation are difficult.
+As Jiang et al.
+(2024) note, even at minimal perplexity, not all knowledge can be extracted—a ‘perplexity curse’ that limits effective data utilization.
+Moreover, compared to supervised fine-tuning, the larger parameter shifts in CPT increase the risk of performance degradation.
+To address these challenges, common data collection strategies rely on the latest Wikipedia snapshots or new domain academic papers Cossu et al.
 (2022); Jin et al.
@@
 Additionally, recent work has explored using highly relevant QA pairs for CIT before CPT Jiang et al.
-(2024) or converting raw CPT corpus into reading comprehension instructions (i.e., CIT data) Cheng et al.
-(2024b) to mitigate degradation and enhance the efficiency of knowledge absorption from CPT data.
-Nevertheless, reliance on oracle LLMs and the complexity of mining patterns limit their practicality.
-In this paper, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide (1) a practical pipeline for collecting CPT data based on model mistakes and (2) a novel training set construction paradigm to efficiently utilize these data, enabling continual model improvement while mitigating degradation.
-To enhance the targetedness and efficiency of data collection, we draw inspiration from the human learning skill of ‘summarizing mistakes’.
+(2024) or converting raw CPT corpora into CIT instructions (e.g., reading comprehension instructions) Cheng et al.
+(2024b), thereby reducing degradation and enhancing knowledge absorption.
+Nevertheless, their dependence on advanced LLMs and the complexity of mining patterns limit their practical applicability.
+In this paper, we propose the Continue Evolving from Mistakes (CEM) method, which: (1) introduces a practical pipeline to collect CPT data derived from model mistakes, and (2) employs a novel training paradigm that integrates CIT and CPT to efficiently utilize these data.
+Together, these innovations enable continual model improvement while mitigating performance degradation.
+To enhance the focus and efficiency of data collection, we draw inspiration from the human learning skill of ‘summarizing mistakes’.
 We argue that a model’s mistakes in ‘exams’ reflect its inherent knowledge deficiencies.
-Therefore, we collect corpora from the internet that contain background knowledge related to these mistakes to directly address these deficiencies.
-To efficiently utilize these CPT data and mitigate forgetting, we design a novel training set construction paradigm that integrates parallel CIT and CPT data.
-In addition, by integrating CEM with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al.
-(2022), we enable iterative and continual model evolution.
-The CEM method consistently demonstrates significant improvements across multiple models in extensive experiments, achieving up to a 29.63% accuracy enhancement.
+Accordingly, we collect background information from the internet related to these identified mistakes, directly addressing the specific knowledge deficiencies they reveal.
+To further optimize CPT data utilization and mitigate forgetting, we construct training sets integrating CIT and CPT data in parallel.
+Coupled with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al.
+(2022), CEM supports iterative, continual model evolution.
+Across extensive experiments, CEM yields substantial gains for multiple models, increasing accuracy by up to 29.63%.
 2 Related Work
@@
 In domain-incremental pre-training, Cossu et al.
-(2022) continually pre-trained LLMs on new multimodal data streams, preparing them for various downstream tasks.
+(2022) continually pre-trained LLMs on new multimodal data streams, preparing them for various asks.
 Ke et al.
@@
 (2024).
-Task-incremental Continual Instruction Tuning (CIT) aims to continually fine-tune LLMs on new task-specific instructions.
+Task-incremental CIT aims to continually fine-tune LLMs on new task-specific instructions.
 While a straightforward approach involves generating new instruction-tuning data and fine-tuning LLMs directly Wang et al.
@@
 We search for extracted knowledge points using Bing and select the top-k results based on relevance, scraping the main textual content as a supplemental corpus.
-To prevent LLMs from simply memorizing answers, we filter out results that contain both the original question and its answer.
+To prevent LLMs from simply memorizing answers, we employ a 13-gram matching to filter out any results that contain both the original question and its answer.
 Moreover, we use newspaper3k [^3] https://github.com/codelucas/newspaper and readability [^4] https://github.com/buriy/python-readability to exclude irrelevant and spam content.
@@
 This can serve as the first step for the next round of CEM.
-It should be noted that supplemental training should not be endless, as it may lead to catastrophic forgetting and diminishing returns in accuracy improvement.
+It should be noted that supplemental training should not be endless, as it may lead to catastrophic forgetting and diminishing returns in accuracy improvement (i.e., ‘perplexity curse’).
 Subsequent experiments explore the impact of the number of training rounds on model performance.
@@
 (2024a).
-To evaluate CEM’s out-of-domain generalization, we apply CEM on the in-domain dataset Xiezhi and use CMMLU and C-Eval Huang et al.
+To evaluate CEM’s out-of-domain generalization, we apply CEM on Xiezhi as the in-domain dataset and use CMMLU and C-Eval Huang et al.
 (2023b) as out-of-domain benchmarks.
@@
 ER: The Enhancement Rate measures the relative improvement brought by CEM on a specific task compared to the initial performance.
-It is calculated by $ER=\frac{R_{k}^{*}-R_{k}^{0}}{R_{k}^{0}},$ where $k$ denotes the task using CEM, $R^{0}_{k}$ is the initial accuracy, and $R^{*}_{k}$ is the accuracy after CEM.
+It is calculated by $ER=\frac{A_{k}^{*}-A_{k}^{0}}{A_{k}^{0}},$ where $k$ denotes the task using CEM, $A^{0}_{k}$ is the initial accuracy, and $A^{*}_{k}$ is the accuracy after CEM.
 AFR: The Average Forgetting Rate Wang et al.
 (2024) measures the average relative degradation brought by CEM on tasks, excluding the task using CEM, compared to their initial performance.
-It is calculated by $AFR=\frac{1}{N-1}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\left(\frac{R_{i}^{0}-R_{i}^{*}}{R_{i}^{0}}\right)$, where $R^{0}_{i}$ is the initial accuracy of task $i$ after the initial fine-tuning, and $R_{i}^{*}$ is the accuracy of tasks $i$ after applying CEM on task $k$.
+It is calculated by $AFR=\frac{1}{N-1}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\left(\frac{A_{i}^{0}-A_{i}^{*}}{A_{i}^{0}}\right)$, where $A^{0}_{i}$ is the initial accuracy of task $i$ after the initial fine-tuning, and $A_{i}^{*}$ is the accuracy of tasks $i$ after applying CEM on task $k$.
 5 Experiments
@@
 Firstly, CEM leads to significant accuracy improvements across all models and tasks.
-All variations of CEM substantially enhance model performance in both the in-domain Xiezhi task and OOD tasks.
+All variations of CEM substantially enhance model performance in both the in-domain (ID) Xiezhi task and OOD tasks.
 For instance, the Llama3-8B-Instruct model achieves accuracy gains of up to 29.63% on the Xiezhi task, 10.92% on the CMMLU task, and 5.40% on the C-Eval task when using CEM-R.
@@
 We analyze the impact of the Supplementary Corpus from different sources—Wikipedia, Bing, and their concatenation—on CEM’s performance.
-To further validate CEM’s generalizability, CMMLU is included as an in-domain task scenario with data proportions consistent with the main experiment.
+To further validate CEM’s applicability, CMMLU is included as an ID task scenario with data proportions consistent with the main experiment.
 Figure 3: This figure presents the accuracy of models using the CEM-P method with different data sources on Xiezhi and CMMLU tasks.
@@
 Although the Mix dataset is larger, it often yields lower accuracy than Wiki or Bing, potentially due to data redundancies.
-Moreover, an excessively large supplemental corpus can dilute the influence of CIT data, reducing its capacity to mitigate catastrophic forgetting, support knowledge extraction, and prevent overfitting.
+Moreover, an excessively large supplemental corpus can dilute the influence of CIT data, reducing its capacity to mitigate catastrophic forgetting, enhance knowledge extraction, and prevent overfitting.
 5.3 Mechanistic Analysis of Extractive and Review Instruction Components
-To clarify the underlying mechanisms by which Extractive Instruction and Review Instruction affect model performance, we conducted experiments on CuteGPT-13B-ift and examined the R2W and W2R metrics.
-Additionally, we explored the effects of varying $\alpha$ (with values $[0,0.5,1]$) on CEM-R to determine whether more diverse and complex Review Instructions yield improved results.
+To clarify the underlying mechanisms by which Extractive Instruction and Review Instruction affect model performance, we conduct experiments on CuteGPT-13B-ift and examine the R2W and W2R metrics.
+Additionally, we explore the effects of varying $\alpha$ (with values $[0,0.5,1]$) on CEM-R to determine whether more diverse and complex Review Instructions yield improved results.
 The corresponding results are illustrated in Figure 4 and Figure 5.
-We summarize the key observations and insights derived from these analyses:
+[…]
+[…]
+We summarize the key observations and insights derived from these experiments:
 (1) CEM-E and CEM-R outperform their ablation groups with General Instruction, demonstrating the effectiveness and positive impact of Extractive and Review Instruction on model performance.
 As shown in Figure 5, the Extractive Instruction enhances W2R performance: compared to CEM-P, CEM-E achieves improvements of 1.45%, 3.22%, and 1.26% across the three models, indicating more effective error correction.
-In addition, the Review Instruction significantly reduces the R2W metric, reflecting a decreased tendency for the model to forget previously acquired correct knowledge.
+In addition, the Review Instruction leads to a significant reduction in the R2W metric, indicating a reduced tendency for the model to forget previously acquired correct knowledge.
+For example, using CEM-R1, Llama3-8B-Instruct’s R2W performance decreases by 5.61% and 4.29%, respectively compared to CEM-P and CEM-E.
 While the performance benefits of the Review Instruction are readily understandable, the exact mechanism by which the Extractive Instruction improves the model’s ability to integrate crucial knowledge from lengthy passages warrants further discussion.
@@
 This process enhances the model’s ability to efficiently extract and utilize essential information.
-(1) CEM-E and CEM-R outperform their ablation groups with General Instruction, demonstrating the effectiveness and positive impact of Extractive and Review Instruction on model performance.
-Figure 5 shows that Extractive Instruction improves W2R, with CEM-E achieving gains of 1.45%, 3.22%, and 1.26% across the three models compared to CEM-P, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the CPT data.
-This improvement leads to more effective error correction.
-Additionally, the Review Instruction significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
 (2) The proportion factor $\boldsymbol{\alpha}$ has minimal impact on the effectiveness of Review Instruction, as R0, R0.5, and R1 exhibit similar effects on LLM performance.
 This indicates that reviewing knowledge, regardless of format, effectively consolidates it.
-[…]
-Figure 5: The figure presents the R2W and W2R metrics of the models after CEM supplemental training.
 5.4 Multiple Iterations Analysis
@@
 Initially, 2,000 instances from each dataset are sampled for pre-finetuning to establish a baseline.
-To address forgetting, we implement a random replay strategy, replacing General Instructions with 500 randomly replayed instances from HotpotQA and GSM8K in each training round.
+To address forgetting, we implement a random replay strategy, replacing General Instructions with 400 randomly replayed instances from HotpotQA and GSM8K in each training round.
 Figure 6: The figure shows the experimental results of CuteGPT-13B-ift on the Xiezhi task after multiple rounds of CEM.
@@
 The combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 75.0% in most cases, highlighting its iterative potential.
-However, without a replay strategy, initial rounds show better improvement (possibly because the model does not need to adapt to the more complex multitasking required by random replay), but more rounds lead to greater forgetting, limiting further iteration ,as seen when CEM-R reaches an ER of 62.4% in round three without replay.
+However, without a replay strategy, initial rounds show better improvement (possibly because the model does not need to adapt to the more complex multitasking instructions introduced by random replay), but more rounds lead to greater forgetting, limiting further iteration, as seen when CEM-R reaches an ER of 62.4% in round three without replay.
 CEM-R, with its larger data capacity, achieves significant performance improvements but shows the greatest decline in other tasks, likely due to the increased volume of CPT data diluting the replay effect.
-Combining CEM with other advanced replay-based methods warrants further investigation.
+Combining CEM with other advanced replay-based methods and data balancing warrants further investigation.
 6 Conclusion
-In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient and cost-effective method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes.
-CEM employs a pipeline that leverages model mistakes to gather extensive CPT data, combined with a novel parallel training set construction paradigm integrating both CIT and CPT data for continuous training.
-Experiments on various open-source LLMs show that CEM significantly improves answer accuracy and demonstrates strong potential for sustained evolution through multiple iterations.
+In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient and cost-effective method for collecting CPT data and continually evolving LLMs on specific tasks by identifying and addressing internal knowledge deficiencies revealed by models’ mistakes.
+CEM employs a pipeline that leverages model mistakes to gather extensive and targeted CPT data, combined with a novel parallel training set construction paradigm integrating both CIT and CPT data for continuous training.
+Experiments on various open-source LLMs show that CEM significantly enhances models’ answer accuracy and supplements their internal knowledge deficiencies.
+These improvements effectively generalize to out-of-domain tasks.
+Furthermore, when integrated with replay-based methods, CEM demonstrates substantial potential for sustained evolution across multiple iterations while mitigating performance degradation.
 7 Limitations